### v1

# Efficient simulation of the SABR model

Abstract
We propose an efficient and reliable simulation scheme for the stochastic-alpha-beta-rho (SABR) model. The two challenges of the SABR simulation lie in sampling (i) the integrated variance conditional on terminal volatility and (ii) the terminal price conditional on terminal volatility and integrated variance. For the first sampling procedure, we analytically derive the first four moments of the conditional average variance, and sample it from the moment-matched shifted lognormal approximation. For the second sampling procedure, we approximate the conditional terminal price as a constant-elasticity-of-variance (CEV) distribution. Our CEV approximation preserves the martingale condition and precludes arbitrage, which is a key advantage over Islah’s approximation used in most SABR simulation schemes in the literature. Then, we adopt the exact sampling method of the CEV distribution based on the shifted-Poisson-mixture Gamma random variable. Our enhanced procedures avoid the tedious Laplace inversion algorithm for sampling integrated variance and non-efficient inverse transform sampling of the forward price in some of the earlier simulation schemes. Numerical results demonstrate our simulation scheme to be highly efficient, accurate, and reliable.
keywords:
SABR model, Monte Carlo simulation, constant elasticity of variance process, shifted-Poisson-mixture gamma distribution
journal: arXiv.org
1 Introduction
The stochastic-alpha-beta-rho (SABR) model proposed by Hagan et al. (2002) has been widely adopted in option pricing due to its ability to capture volatility smile or skew using a few parameters, and exhibiting consistency in revealing the dynamic behavior between price and smile. On one hand, it is common in equity options that the implied volatility of out-of-money options are generally higher than that of the in-the-money counterparts, known as the volatility skew. On the other hand, the relatively symmetric volatility smile is common in the foreign exchange options. The SABR model can capture the volatility skew and volatility smile in the respective markets. In addition, the SABR model exhibits consistency in the dynamic behavior of volatility smile or skew, where an increase of the price of the underlying asset leads to shifting of the volatility smile or skew to a higher price (Hagan et al., 2002). The SABR model can generate such co-movements correctly while the local volatility model fails to do so (Derman and Kani, 1998; Dupire, 1997).
In their original paper, Hagan et al. (2002) derived an analytic implied volatility formula under the SABR model via an asymptotic expansion in small time-to-maturity. It has been a standard practice for practitioners to obtain the option price from the implied volatility using the Black–Scholes formula. However, Hagan et al. (2002)’s approximation becomes unreliable under large time-to-maturity or when the option is deep-out-of-the-money. As a result, one important stream of the SABR model research has been improving the implied volatility approximation (Obłój, 2007; Paulot, 2015; Lorig et al., 2017; Gulisashvili et al., 2018; Yang et al., 2017; Choi and Wu, 2021a). Notably, Antonov and Spector (2012) obtained a more accurate approximation by mapping the implied volatilities under the correlated cases to those of the uncorrelated cases, under which more analytic properties of the SABR model become available.
Another stream of research, to which this paper belongs, is developing efficient simulation algorithms of the SABR model. With growing popularity of the SABR model, there has been demand for pricing path-dependent derivatives using the model. Since analytic expansion approaches are limited to pricing the European options, the Monte Carlo method becomes a natural choice. In particular, the research on the SABR simulation has been stimulated by less than perfect performance of the Euler and Milstein time-discretization schemes. Since the SABR model imposes an absorbing boundary condition at the origin, the time discretization scheme should be more carefully implemented; naive truncation at the origin leads to large deviation from the original SABR model. Typically, it is necessary to use a small time discretization step to decrease the discretization errors due to its low order of convergence, thus suffering from heavy computation load.
Several simulation algorithms (Chen et al., 2012; Cai et al., 2017; Leitao et al., 2017a, b; Cui et al., 2018; Grzelak et al., 2019; Kyriakou et al., 2023) have been proposed for efficient simulation of the SABR model. Simulating the SABR process over a relatively large time step requires sequentially sampling terminal volatility, average variance, and terminal price at the next simulation time point. Since the last two sampling procedures are challenging, existing methods are differentiated with reference to different choices of the algorithms for these two steps. These earlier simulation algorithms have achieved some success in that they overcome the issue of the time-discretization method. However, there are still many gaps to be filled in order to make SABR simulation more efficient.
Regarding the average variance conditional on the terminal volatility, Chen et al. (2012) sampled the quantity through the mean-and-variance matched lognormal (LN) variable. While the sampling procedures are fast, the time step cannot be large since the derived mean and variance are valid in small-time limit only. For more accurate sampling, researchers used various version of the transform of the average variance. Leitao et al. (2017a, b) derived the approximate Fourier transform of the unconditional average variance recursively, while Cai et al. (2017) adopted the analytically available Laplace transform of the reciprocal of the conditional average variance (Matsumoto and Yor, 2005). However, these approaches are cumbersome and time-consuming due to the tedious numerical inverse transform for constructing the cumulative distribution function (CDF) and iterative root-finding for inverse transform sampling. Cui et al. (2021) approximated the volatility dynamics with the continuous-time Markov chain (CTMC) over volatility grids. Their approach is also not immune from heavy computation and complicated implementation. Kyriakou et al. (2023) used moment-matched Pearson family of distributions for efficient sampling. For such purpose, they rely on numerical evaluation of the four moments from the conditional Laplace transform.
Sampling the terminal price conditional on the terminal variance and average variance is perhaps more challenging. Unlike the other stochastic volatility models that are based on the geometric Brownian motion (BM), the constant-elasticity-of-variance (CEV) feature of the SABR model does not admit an analytically tractable form of the conditional price distribution. As a result, the distribution has to be approximated by an analytic form first even before the corresponding sampling algorithm is considered. For the approximation of the conditional price distribution, almost all existing algorithms (Chen et al., 2012; Cai et al., 2017; Leitao et al., 2017a, b; Cui et al., 2018; Grzelak et al., 2019; Kyriakou et al., 2023) adopted Islah (2009)’s noncentral chi-squared (NCX2) distribution approximation unquestionably. However, the failure of the martingale property in this approximation has not been critically examined among these papers. Leitao et al. (2017a, b) applied an ad-hoc correction on the simulation prices to enforce the martingale property. Therefore, the possibility of better approximation of the conditional price distribution that preserves the martingale should be explored.
The lack of fast sampling algorithm for the approximated distribution is the last hurdle in constructing an efficient SABR simulation algorithm. The inverse transform sampling with numerical root-finding is accurate but tediously slow. The mean-and-variance-match quadratic Gaussian sampling has been adopted in Chen et al. (2012), Leitao et al. (2017a, b), and Cui et al. (2021) to speed up sampling. However, it is limited to small time step due to the nature of approximation. Also, it cannot be used efficiently when the price is close to zero due to the absorbing boundary condition.
This paper proposes an accurate, efficient, and reliable SABR simulation algorithm, filling the gaps in all steps of the simulation procedure. The contributions of this paper are three-fold. Firstly, we derive the first four conditional moments of the conditional average variance analytically. We then use the shifted lognormal (SLN) distribution matching the moments for fast sampling of the average variance. As we use the moments that are higher in order and exact for any arbitrary time step, the validity of our algorithm can be extended to larger time step. Secondly, we present an alternative approximation of the conditional price distribution based on the CEV process. Our approximation is superior to the widely adopted Islah (2009)’s approximation because it preserves the martingale property by construction. Lastly, we employ the algorithms of Makarov and Glew (2010) and Kang (2014) that are capable of sampling the CEV price process for an arbitrary time step. Their algorithms are fast since the CEV price can be expressed by a compound random variable composed of three elementary (i.e., gamma-Poison-gamma) random variables.
This paper is organized as follows. Section 2 presents the formulation of the SABR model and reviews some of its analytic properties. Section 3 describes our SLN approximation for sampling the average variance. In Section 4, we present the CEV distribution approximation of the conditional forward price, and an exact sampling algorithm for the CEV distribution via the shifted Poisson-mixture Gamma distribution. In Sections 3 and 4, our simulation algorithm is compared to those in the existing literature. In particular, we argue that the CEV distribution is a better alternative to the widely adopted Islah (2009)’s approximation as it preserves the martingale condition. Section 5 presents our comprehensive numerical experiments that serve to assess accuracy, efficiency and reliability of different simulation schemes and analytic approximation methods, focusing on the more challenging choices of parameters, such as large time-to-maturity. Conclusive remarks are summarized in Section 6.
2 Formulation of the SABR model and its special cases
In this section, we introduce the SABR model and review some of its analytic properties. The two steps of the SABR simulation are outlined. Specifically, we discuss some special cases where the SABR simulation becomes straightforward.
2.1 SABR model
The governing stochastic differential equations (SDE) for the SABR volatility model (Hagan et al., 2002) are given by
$\frac{\mathrm{d}F_{t}}{F_{t}^{\beta}}=\sigma_{t}\,\mathrm{d}W_{t}\quad\text{and}\quad\frac{\mathrm{d}\sigma_{t}}{\sigma_{t}}=\nu\,\mathrm{d}Z_{t},\quad(0\leq\beta\leq 1),$ (1)
where $F_{t}$ and $\sigma_{t}$ are the stochastic processes for the forward price and volatility, respectively, $\nu$ is the vol-of-vol, $\beta$ is the elasticity of variance parameter, and $W_{t}$ and $Z_{t}$ are the standard BMs that are correlated with correlation coefficient $\rho$. We also define two variables for later use:
$\rho_{\ast}:=\sqrt{1-\rho^{2}}\quad\text{and}\quad\beta_{\ast}:=1-\beta.$
2.2 Average variance, conditional expectation, and two-step simulation procedures
Regarding the simulation step, without loss of generality, we sample the volatility and forward price at time $t+h$ (i.e., $\sigma_{t+h}$ and $F_{t+h}$) for some time step $h$ given time $t$. Accordingly, the filtration up to time $t$ (e.g., $\sigma_{t}$ and $F_{t}$) is implicitly assumed.
Since the volatility process in Eq. (1) follows a geometric BM, $\sigma_{t}=\sigma_{0}\exp\left(\nu Z_{t}-\nu^{2}t/2\right)$, simulating $\sigma_{t}$ from time $t$ to $t+h$ can be easily done by observing
$\sigma_{t+h}=\sigma_{t}\exp\left(\nu Z_{h}-\frac{\nu^{2}h}{2}\right)=\sigma_{t}\exp\left(\hat{\nu}\hat{Z}\right)\quad\quad\text{where}\quad\hat{Z}\sim\mathcal{N}\left(-\frac{\hat{\nu}}{2},1\right).$ (2)
Here, $\hat{\nu}:=\nu\sqrt{h}$ is the standard deviation of $\ln\sigma_{t}$ over the time step $h$. This quantity is used as a measure of small-time limit.
Like the simulation schemes of other SV models, integrated variance conditional on initial and terminal volatility values plays an important role in the SABR model simulation. Given the initial and terminal volatility (i.e., $\sigma_{t}$ and $\sigma_{t+h}$), we define the conditional average variance between $t$ and $t+h$ as
$I_{t}^{h}(\sigma_{t+h}):=\frac{1}{\sigma_{t}^{2}h}\int_{t}^{t+h}\sigma_{s}^{2}\,\mathrm{d}s\;\Big{|}_{\sigma_{t+h}}.$ (3)
Here, we normalize the integral by $\sigma_{t}^{2}h$ in order that $I_{t}^{h}$ becomes a dimensionless quantity in the order of one. Specifically, $I_{t}^{h}$ converges to one in the limit of $\nu\downarrow 0$. From the simulation step in Eq. (2), $I_{t}^{h}$ is equivalently conditioned by $\hat{Z}$ as
$I_{t}^{h}(\hat{Z})\;\sim\;\int_{0}^{1}e^{2\hat{\nu}Z_{s}}\,\mathrm{d}s\;\Big{|}_{Z_{1}=\hat{Z}}\;.$
For notational simplicity, we omit the dependence on $\sigma_{t+h}$ or $\hat{Z}$ and simply write $I_{t}^{h}$.
Almost all SABR simulation algorithms are required to perform the following two steps, given the filtration up to $t$, to sufficient accuracy:
Step 1
: simulation of the average variance $I_{t}^{h}$ conditional on $\sigma_{t+h}$ (or $\hat{Z}$),
Step 2
: simulation of the forward price $F_{t+h}$ conditional on $\sigma_{t+h}$ and $I_{t}^{h}$.
As both simulation steps pose challenges, the previous studies on the topic have proposed numerous algorithms for the two steps. This paper also aims to innovate the two steps, details are presented in Sections 3 and 4, respectively.
Regarding Step 2, let us also define the conditional expectation of $F_{t+h}$:
$\bar{F}_{t}^{h}(\sigma_{t+h},I_{t}^{h}):=E\left(F_{t+h}\,|\,\sigma_{t+h},I_{t}^{h}\right),$ (4)
which is essential for the discussion of the simulation algorithms of Step 2. We also omit the dependence on $\sigma_{t+h}$ and $I_{t}^{h}$, and simply write $\bar{F}_{t}^{h}$.
2.3 Special cases
Before introducing our simulation algorithm, we review several special cases of the SABR model. In these cases, Step 2 is relatively easier while Step 1 remains challenging. Reviewing these cases also helps building intuitions for our new algorithm.
Normal ($\beta=0$) SABR model: When $\beta=0$, conditional on $\sigma_{T}$ and $I_{0}^{T}$, integrating $F_{t}$ from $t=0$ to $T$ reveals that the terminal forward price $F_{T}$ follows the normal distribution with mean $\bar{F}_{0}^{T}$ and variance $\rho_{\ast}^{2}\sigma_{0}^{2}TI_{0}^{T}$:
$F_{T}\,|\,\sigma_{T},I_{0}^{T}\;\sim\;\mathcal{N}\left(\bar{F}_{0}^{T},\rho_{\ast}^{2}\sigma_{0}^{2}TI_{0}^{T}\right)$
where the conditional expectation $\bar{F}_{0}^{T}$ is given by
$\bar{F}_{0}^{T}=F_{0}+\frac{\rho}{\nu}\big{(}\sigma_{T}-\sigma_{0}\big{)}.$ (5)
Therefore, Step 2 becomes trivial as long as $I_{t}^{h}$ can be sampled accurately in Step 1. Choi et al. (2019) have also found an exact closed-form expression for sampling $F_{T}$ without separating Steps 1 and 2. Moreover, note that $F_{T}$, as well as $\bar{F}_{0}^{T}$ can be negative in this special case as the origin is not a boundary. The normal SABR model is different from the $\beta\downarrow 0$ limit of the SABR model, where the origin is an absorbing boundary. In the view of the above, the SABR model under $\beta=0$ is not a concern of our study.
Lognormal ($\beta=1$) SABR model: When $\beta=1$, conditional on $\sigma_{T}$ and $I_{0}^{T}$, integrating $\ln F_{t}$ from $t=0$ to $T$ yields that the terminal forward price $F_{T}$ follows the LN distribution with mean $\bar{F}_{0}^{T}$ and log variance $\rho_{\ast}^{2}\sigma_{0}^{2}TI_{0}^{T}$:
$F_{T}\,|\,\sigma_{T},I_{0}^{T}\;\sim\;\bar{F}_{0}^{T}\,\exp\left(\,\rho_{\ast}\sigma_{0}\sqrt{TI_{0}^{T}}\,X-\frac{\rho_{\ast}^{2}\sigma_{0}^{2}TI_{0}^{T}}{2}\right),$ (6)
where $X$ is a standard normal variate independent of $\sigma_{T}$ and $I_{0}^{T}$. The conditional expectation $\bar{F}_{0}^{T}$ is given by
$\bar{F}_{0}^{T}=F_{0}\exp\left(\frac{\rho}{\nu}(\sigma_{T}-\sigma_{0})-\frac{\rho^{2}\sigma_{0}^{2}TI_{0}^{T}}{2}\right).$ (7)
Therefore, Step 2 becomes trivial as well when $\beta=1$.
This special case provides an important observation, which derives an intuition for our CEV approximation when $10<\beta<$ 1 (to be discussed Section 4).
Remark 1.
The joint distribution of $\sigma_{T}$ and $I_{0}^{T}$ satisfies
$E\left[\exp\left(\frac{\rho}{\nu}(\sigma_{T}-\sigma_{0})-\frac{\rho^{2}\sigma_{0}^{2}TI_{0}^{T}}{2}\right)\right]=1,$
for any initial volatility $\sigma_{0}>0$.
Proof.
It follows from the law of total expectation:
$E(\bar{F}_{0}^{T})=E\left(E\left(F_{T}|\sigma_{T},I_{0}^{T}\right)\right)=E(F_{T})=F_{0}.\qed$
Zero vol-of-vol ($\nu=0$) SABR model: When $10<\beta<$ 1, the SABR model is visualized as the stochastic volatility version of the CEV model:
$\frac{\mathrm{d}F_{t}}{F_{t}^{\beta}}=\sigma_{0}\,\mathrm{d}W_{t},$ (8)
with the absorbing boundary condition at the origin.
Definition 1.
We define the distribution of $F_{T}$ resulting from the CEV dynamics in Eq. (8) as the CEV distribution,
$F_{T}\sim\mathcal{CEV}_{\beta}(F_{0},\sigma_{0}^{2}T),$
parameterized by the elasticity-of-variance parameter $\beta$, mean $F_{0}$ [^1] If $F_{T}\sim\mathcal{CEV}_{\beta}(F_{0},\sigma_{0}^{2}T)$, then $E(F_{T})=F_{0}$ regardless of the values of $\beta$ and $\sigma_{0}^{2}T$., and variance $\sigma_{0}^{2}T$. The probability distribution function (PDF) and CDF of the CEV distribution will be shown later in Proposition 5.
Therefore, in the zero vol-of-vol ($\nu\downarrow 0$) limit, $F_{T}$ from the SABR model (unconditionally) follows $\mathcal{CEV}_{\beta}(F_{0},\sigma_{0}^{2}T)$. Step 2 is reduced to sampling $F_{T}$ from the CEV distribution. However, this is not a trivial task as we discuss later in Section 4.
Uncorrelated ($\rho=0$) SABR model: The terminal forward price $F_{T}$ also exhibits a CEV distribution under non-zero vol-of-vol ($\nu>0$) and zero correlation ($\rho=0$) case, albeit conditionally. Conditional on $I_{0}^{T}$, the terminal price $F_{T}$ follows the CEV distribution:
$F_{T}\,|\,I_{0}^{T}\sim\mathcal{CEV}_{\beta}(F_{0},\sigma_{0}^{2}TI_{0}^{T}).$
Basically, $TI_{0}^{T}$ plays the role of stochastic time clock of the CEV model. As in the $\nu=0$ case, Step 2 is reduced to sampling $F_{T}$ from the CEV distribution in this case. Unlike the $\beta=1$ case, $F_{T}$ is conditioned by $I_{0}^{T}$ only since conditional expectation is always $F_{0}$ without depending on $\sigma_{T}$. Thus, $I_{0}^{T}$ should be understood as the unconditional average variance in this context. Therefore, the European option price can be simplified to one dimensional integral of the CEV option price over the (unconditional) distribution of $I_{0}^{T}$. Choi and Wu (2021b) have approximated the integral using the Gaussian quadrature together with the LN distribution fitted to the (unconditional) mean and variance of $I_{0}^{T}$.
Completing the discussion on the four special cases, we are left with the most general case with $10<\beta<$ 1 and $\rho\neq 0$, under which the conditional distribution of $F_{T}$ is not analytically tractable. One has to resort to an approximation. The approximation should be reduced to the special cases discussed above under various continuous limits.
3 First simulation step: Sampling average variance $I_{t}^{h}$
This section presents our simulation method for sampling $I_{t}^{h}$. We also compare with other most popular simulation methods, highlighting the computational advantages of our simulation scheme.
3.1 Sampling $I_{t}^{h}$ from the moment-matched SLN approximation
We sample $I_{t}^{h}$ from the SLN distribution matched to the first three moments of $I_{t}^{h}$.
Definition 2 (SLN variable).
The SLN random variable, $Y\sim\mathcal{SLN}(\mu,\sigma^{2},\lambda)$, is given by
$Y\sim\mu\left[(1-\lambda)+\lambda\exp\left(\sigma X-\frac{\sigma^{2}}{2}\right)\right]\quad\text{with}\quad\mu>0,\;\sigma>0,\;0<\lambda\leq 1,$
where $X$ is a standard normal variate. Here, $\mu$ is the mean of $Y$, $\lambda$ and $\sigma^{2}$ are the weight and variance of the LN component, respectively. When $\lambda=1$, $Y$ is reduced to the LN variable with mean $\mu$ and log variance $\sigma^{2}$.
The mean, coefficient of variation, skewness, and ex-kurtosis of $Y$ are given by
$E(Y)=\mu,\quad\text{Cv}(Y)=\lambda\sqrt{w},\quad\text{Sk}(Y)=\sqrt{w}(w+3),\quad\text{and}\quad\text{Ek}(Y)=w(w^{3}+6w^{2}+15w+16),$
where $w=\exp\left(\sigma^{2}\right)-1\;(w>0)$.
Remark 2.
While the skewness and ex-kurtosis of distribution $Y$ follow the conventional definitions, the coefficient of variation is defined as the ratio of standard deviation to mean:
$\mathrm{Cv}(Y)=\frac{\sqrt{\mathrm{Var}(Y)}}{E(Y)}.$
The coefficient of variation is an important dimensionless quantity that characterizes non-negative distributions such as the LN or SLN distributions.
Proposition 1 (SLN distribution matched to the first three moments).
Given mean $\mu>0$, coefficient of variance $v>0$, and skewness $s>0$, the parameters of the SLN distribution matching $\mu$, $v$, and $s$ are analytically given by
$\sigma=\sqrt{\ln(1+w)}\quad\text{for}\quad w=4\sinh^{2}\left(\frac{1}{6}\operatorname{arcosh}\left(1+\frac{s^{2}}{2}\right)\right)$
$\text{and}\quad\lambda=\frac{v}{2\sinh\left(\frac{1}{6}\operatorname{arcosh}\left(1+\frac{s^{2}}{2}\right)\right)}.$
Proof.
We first fit $\sigma$ to the given skewness $s$:
$s^{2}=w(w+3)^{2}\quad\text{for}\quad w=\exp\left(\sigma^{2}\right)-1.$
Substituting $x=w+2$ transforms the equation to a special case of the cubic equation:
$x^{3}-3x-(2+s^{2})=0,$
where the unique positive root $x$ is analytically given by  [^2] See https://en.wikipedia.org/wiki/Cubic_equation#Hyperbolic_solution_for_one_real_root.
$x=2\cosh\left(\frac{1}{3}\operatorname{arcosh}\left(1+\frac{s^{2}}{2}\right)\right).$
The expression for $w$ follows from $\cosh\theta-1=2\sinh^{2}(\theta/2)$. Next, we fit $\lambda$ to the given coefficient of variance $v$:
$\lambda=\frac{v}{\sqrt{w}}=\frac{v}{2\sinh\left(\frac{1}{6}\operatorname{arcosh}\left(1+\frac{s^{2}}{2}\right)\right)}.\qed$
Remark 3.
For an LN random variable (i.e., $\lambda=1$), $\sigma$ is determined by matching the coefficient of variation $v$:
$\sigma=\sqrt{\ln\left(1\,+\,v^{2}\right)}.$
Chen et al. (2012) has sampled $I_{t}^{h}$ through the LN random variable fitted to the mean and variance of $I_{t}^{h}$ in the small-time limit.
Next, we derive the first four conditional moments of $I_{t}^{h}$, whose analytic formulas are presented in the next proposition.
Proposition 2 (Conditional moments of average variance).
Given $\sigma_{t}$ and $\sigma_{t+h}$, the first four raw moments of $I_{t}^{h}$ (i.e., $\mu=E(I_{t}^{h})$ and $\mu^{\prime}_{k}=E\left((I_{t}^{h})^{k}\right)$) are obtained as below.
$\mu$ $=\left(\frac{\sigma_{t+h}}{\sigma_{t}}\right)m_{1},$
$\mu^{\prime}_{2}$ $=\left(\frac{\sigma_{t+h}}{\sigma_{t}}\right)^{2}\frac{1}{\hat{\nu}^{2}}\left[m_{2}-cm_{1}\right],$
$\mu^{\prime}_{3}$ $=\left(\frac{\sigma_{t+h}}{\sigma_{t}}\right)^{3}\frac{1}{8\hat{\nu}^{4}}\left[3m_{3}-8cm_{2}+\left(4c^{2}+1\right)m_{1}\right],$
$\mu^{\prime}_{4}$ $=\left(\frac{\sigma_{t+h}}{\sigma_{t}}\right)^{4}\frac{1}{24\,\hat{\nu}^{6}}\left[2m_{4}-9cm_{3}+\left(12c^{2}+2\right)m_{2}-c\left(4c^{2}+3\right)m_{1}\right],$
where the conditioning variable, $\sigma_{t+h}$ and $\hat{Z}$, are exchangeable by $\hat{Z}=\frac{1}{\hat{\nu}}\ln(\sigma_{t+h}/\sigma_{t})$, and the coefficients, $c$ and $m_{k}\,(k=1,2,\dots,4)$, are the functions of them:
$c(\hat{Z}):=\cosh(\hat{\nu}\hat{Z})=\frac{1}{2}\left(\frac{\sigma_{t+h}}{\sigma_{t}}+\frac{\sigma_{t}}{\sigma_{t+h}}\right)\quad\text{and}\quad m_{k}(\hat{Z})=\frac{N(\hat{Z}+k\hat{\nu})-N(\hat{Z}-k\hat{\nu})}{2k\hat{\nu}\;n\left(\sqrt{\hat{Z}^{2}+(k\hat{\nu})^{2}}\right)},$
and $n(z)$ and $N(z)$ are the PDF and CDF of the standard normal distribution, respectively.
Proof.
The analytic derivation is based on the conditional moments of the exponential functional of BM (Matsumoto and Yor, 2005, (5.4)). See A for the derivation. ∎
Remark 4.
Kennedy et al. (2012, Eqs. (7) and (8)) have derived $\mu$ and $\mu^{\prime}_{2}$ independently without resorting to Matsumoto and Yor (2005).
Remark 5.
From the first four raw moments, the coefficient of variation $v$, skewness $s$, and ex-kurtosis $\kappa$ of $I_{t}^{h}$ are obtained as
$v=\frac{\sqrt{\mu_{2}}}{\mu},\quad s=\frac{\mu^{\prime}_{3}-3\mu\mu^{\prime}_{2}+2\mu^{3}}{\mu_{2}\sqrt{\mu_{2}}},\quad\text{and}\quad\kappa=\frac{\mu^{\prime}_{4}-4\mu\mu^{\prime}_{3}+6\mu^{2}\mu^{\prime}_{2}-3\mu^{4}}{\mu_{2}^{2}}-3,$
where $\mu_{2}=\mu^{\prime}_{2}-\mu^{2}$ is the variance.
From Propositions 1 and 2, it is possible to sample $I_{t}^{h}$ from the SLN variable whose parameters are fitted to the first three moments of $I_{t}^{h}$ exactly. However, we take a simpler computational approach that uses a fixed $\lambda$ obtained from the small-time limit. It is shown to be effective as well.
Proposition 3 (SLN in the small-time limit).
In the small-time limit ($\hat{\nu}\downarrow 0$), $I_{t}^{h}$ is approximately sampled via an SLN variable with $\lambda=5/6$:
$I_{t}^{h}\sim\frac{\mu}{6}\left[1+5\exp\left(\sigma X-\frac{\sigma^{2}}{2}\right)\right]\quad\text{with}\quad\sigma=\sqrt{\ln\left(1+\frac{36}{25}v^{2}\right)},$ (9)
where $\mu$ and $v$ are the mean and coefficient of variance of $I_{t}^{h}$ given by Proposition 2 and Remark 5.
Proof.
In A, the coefficient of variation and skewness of $I_{t}^{h}$ are expanded around $\hat{\nu}=0$ as
$v=\frac{\sqrt{\mu_{2}}}{\mu}=\frac{\hat{\nu}}{\sqrt{3}}+O(\hat{\nu}^{3})\quad\text{and}\quad s=\frac{\mu_{3}}{\mu_{2}\sqrt{\mu_{2}}}=\frac{6\sqrt{3}}{5}\hat{\nu}+O(\hat{\nu}^{3}).$
Using Taylor’s expansion, we have
$2\sinh\left(\frac{1}{6}\operatorname{arcosh}\left(1+\frac{s^{2}}{2}\right)\right)=\frac{s}{3}-\frac{s^{3}}{81}+O(s^{5}),$
so that the shift parameter $\lambda$ converges to
$\lambda=\frac{v}{2\sinh\left(\frac{1}{6}\operatorname{arcosh}\left(1+\frac{s^{2}}{2}\right)\right)}\to\frac{5}{6}\quad\text{as}\quad\hat{\nu}\downarrow 0.$
Once $\lambda$ is determined, we fit $\sigma$ to match the coefficient of variation:
$\sigma=\sqrt{\ln\left(1+\frac{v^{2}}{\lambda^{2}}\right)}=\sqrt{\ln\left(1+\frac{36}{25}v^{2}\right)}.$
As we demonstrate in the numerical results shown in Figure 1, this approximation is highly reliable not only for small $\hat{\nu}$ but also for $O(\hat{\nu})=1$. ∎
Figure 1: Variance (left), skewness (middle), and Ex-kurtosis (right) of $I_{t}^{h}$ as functions of $\hat{Z}$ for $\hat{\nu}=0.4$. We compare our SLN approximation with $\lambda=5/6$ (yellow), the SLN approximation fitted up to skewness (green), and the LN approximation fitted up to variance (blue) to the true values (dashed).
3.2 Comparison to other simulation methods in literature
Our sampling method for $I_{t}^{h}$ is much faster and easier to implement compared to other methods in the literature, such as Chen et al. (2012), Cai et al. (2017), Leitao et al. (2017a, b) and Cui et al. (2021).
Our method is considered as an extension of Chen et al. (2012) as they use an LN approximation matched to mean and variance of $I_{t}^{h}$. However, they derive the mean and variance only in the small-time limit, so the time step should be small accordingly. We enhance the method by deriving the moments analytically and matching to higher moments with the SLN approximation. Figure 1 displays the accuracy of variance, skewness, and ex-kurtosis of $I_{t}^{h}$ obtained from our SLN approximation. The left figure shows that Chen et al. (2012)’s small-time approximation of the variance is not accurate across the dependency $Z$ while our approximation is exact. The middle figure shows that the SLN approximation gives the correct skewness value as the parameters are fitted up to skewness. Moreover, the skewness under the simple SLN scheme with $\lambda=5/6$ also agrees well for $\nu=0.2$. Even though the two SLN schemes do not use the ex-kurtosis in the parameter calibration, the ex-kurtosis is close to the true value, justifying our choice of the SLN random variable.
Cai et al. (2017) computed the analytic form of the Laplace transform of the CDF of the reciprocal of the average variance 1/$I_{t}^{h}$, then used the Euler inversion formula to obtain the CDF of 1/$I_{t}^{h}$. After obtaining the CDF, they used the inverse transform method, which requires some tedious root-finding procedures such as Newton’s method or bisection method to sample $I_{t}^{h}$.
Leitao et al. (2017a, b) used recursion schemes to find the characteristic function of the logarithm of $I_{t}^{h}$ and then employed the inverse Fourier technique COS method (Fang and Oosterlee, 2008) to recover its PDF. They then computed the Pearson correlation coefficient of $\ln I_{t}^{h}$ and $\ln\sigma_{t+h}$, then the bivariate copula distribution based on the CDF of $\ln\sigma_{T}$ and approximate CDF of $\ln I_{t}^{h}$. Finally, they adopted the direct inverse transform method based on linear interpolation to sample $I_{t}^{h}$.
Cui et al. (2021) used the continuous-time Markov chain (CTMC) to approximate the space of the volatility process with finite number of states. In their procedure, they first fixed the volatility grids, computed the transition probability density at these grids based on the exponential of the generator matrix and used the inverse transform method to sample the volatility. After obtaining the volatility values, they computed the conditional characteristic function of the integrated variance, again based on the exponential of the generator matrix. They then adopted the Fourier sampler to get the CDF of the integrated variance, which first computed the PDF of the integrated variance via the orthogonal projection. Finally, they simulated the integrated variance using the closed-form inverse function.
Putting aside the issue of accuracy of the simulation methods of $I_{t}^{h}$, these simulation methods invariably include some tedious and time-consuming procedures such as the adoption of the inverse Laplace/Fourier transform calculation, root-finding steps in the inverse transform methods or computation of the exponential of matrices. However, our SLN simulation methods for $I_{t}^{h}$ exhibits easy and fast implementation. As shown in the numerical experiments reported in Section 5, our simulation methods for $I_{t}^{h}$ demonstrate significant savings in computational time and high level of accuracy.
4 Second simulation step: Sampling forward price $F_{t+h}$
This section presents our CEV approximation for $F_{t+h}$ and its sampling method.
4.1 CEV approximation of the conditional price
Proposition 4 proposes our approximation of $F_{t+h}$ for the general case. Our choice of the approximation is a very natural and intuitive extension of the special cases discussed in Section. 2.3. The only technical component is the choice of the conditional terminal forward price that we adopt from Remark 1.
Proposition 4 (CEV approximation of the conditional terminal forward price).
For $10<\beta<$ 1 and $-1\leq\rho\leq 1$, conditional on $\sigma_{t+h}$ and $I_{t}^{h}$ (and the filtration up to time $t$), we approximate $F_{t+h}$ as a CEV distribution
$F_{t+h}\,|\,\sigma_{t+h},I_{t}^{h}\;\sim\;\mathcal{CEV}_{\beta}(\bar{F}_{t}^{h},\rho_{\ast}^{2}\sigma_{t}^{2}hI_{t}^{h}),$ (10)
where the conditional expectation $\bar{F}_{t}^{h}$ is approximated by
$\bar{F}_{t}^{h}\approx F_{t}\exp\left(\frac{\rho(\sigma_{t+h}-\sigma_{t})}{\nu F_{t}^{\beta_{\ast}}}-\frac{\rho^{2}\sigma_{t}^{2}hI_{t}^{h}}{2F_{t}^{2\beta_{\ast}}}\right).$ (11)
Proof.
We provide the rationales for each component of the CEV approximation. First, with regard to distribution, the choice of $\mathcal{CEV}_{\beta}$ is a natural extension of the $\nu=0$ and $\rho=0$ cases, where the forward price is distributed by $\mathcal{CEV}_{\beta}$ unconditionally and conditionally, respectively. Given that we choose the CEV distribution as the approximation, we show how to select the mean and variance. The variance, $\rho_{\ast}^{2}\sigma_{t}^{2}hI_{t}^{h}$, is also a natural choice, since it is consistent with all special cases in Section 2.3. The variance in the $\rho=0$ case is merely the special cases of $\rho_{\ast}^{2}\sigma_{t}^{2}hI_{t}^{h}$ when $\rho_{\ast}=1$.
Lastly, we choose the conditional expectation $\bar{F}_{t}^{h}$ as the mean of the CEV distribution. Note that it is analytically intractable to obtain the exact conditional expectation under the original SABR dynamics when $\rho\neq 0$. Instead, we approximate the SABR dynamics in the interval $[\,t,t+h]$ by the lognormal SABR model:
$\frac{\mathrm{d}F_{s}}{F_{s}}=\frac{\sigma_{s}}{F_{s}^{\beta_{\ast}}}\,\mathrm{d}W_{s}\approx\frac{\sigma_{s}}{F_{t}^{\beta_{\ast}}}\,\mathrm{d}W_{s}\quad\text{for}\quad t\leq s\leq t+h,$ (12)
where the stochastic volatility $\sigma_{s}$ is replaced by $\sigma_{s}/F_{t}^{\beta_{\ast}}$. Under this approximation, the conditional expectation in Eq. (11) follows from $\beta=1$ case in Eq. (7). Note that the lognormal SABR approximation is too loose to be used for sampling $F_{t+h}$. We only use it for approximating $\bar{F}_{t}^{h}$. Most importantly, our choice of $\bar{F}_{t}^{h}$ ensures the martingale property of $F_{t}$. Based on Remark 1, we observe
$E(F_{t+h})=E\left(E(F_{t+h}|\sigma_{t+h},I_{t}^{h})\right)=E(\bar{F}_{t}^{h})=F_{t},$ (13)
when $F_{t+h}$ is sampled according to this proposition. The martingale condition significantly improves the accuracy of our simulation algorithm (see numerical results illustrated in Figure 3). ∎
4.2 Sampling conditional price from a CEV distribution
The availability of analytic distribution function does not guarantee an efficient sampling algorithm of the random variable. The slow root-finding step is often used to invert the distribution function, and it has been the case with some existing SABR simulation methods (e.g., Cai et al. (2017)). Our SABR simulation algorithm would be completed when our CEV approximation in Proposition 4 is paired with an efficient sampling algorithm of the CEV distribution. For the purpose, we adopt the algorithms of Makarov and Glew (2010) and Kang (2014) for sampling the CEV distribution exactly. Despite its simplicity and efficiency, the algorithm is not widely known in the literature.
Before further discussion, let us present probability distributions that are related to the CEV distribution, and review several related analytic properties.
Definition 3 (Noncentral chi-squared random variable).
Let ${\chi^{2}}(\delta,r)$ be the NCX2 random variable with degree of freedom $\delta$ and noncentrality $r$. The PDF and CDF of ${\chi^{2}}(\delta,r)$ are respectively expressed by
$f_{\chi^{2}}(x\,;\,\delta,r)=\frac{1}{2}\left(\frac{x}{r}\right)^{\alpha/2}I_{\alpha}(\sqrt{x\,r})\;e^{-(x+r)/2}\quad\text{and}\quad P_{\chi^{2}}(x;\delta,r)=\int_{0}^{x}f_{\chi^{2}}(t\,;\,\delta,r)\,\mathrm{d}t,$ (14)
where $\alpha=\delta/2-1$ and $I_{\alpha}(z)$ is the modified Bessel function of the first kind,
$I_{\alpha}(z)=\sum_{k=0}^{\infty}\frac{(z/2)^{\alpha+2k}}{k!\,\Gamma(k+\alpha+1)}.$
Remark 6 (Relation to the gamma distribution).
The central chi-squared (i.e., $r=0$) distribution is a gamma distribution. Let $\mathcal{G}(\alpha)$ be a gamma random variable with shape parameter $\alpha$ and unit scale parameter, the two random variables are related by
${\chi^{2}}(\delta,0)\sim 2\,\mathcal{G}(\delta/2).$
The PDF and CDF of $\mathcal{G}(\alpha)$ are respectively expressed by  [^3] The CDF, $P_{\mathcal{G}}(x;\alpha)$, is the scaled lower incomplete gamma function $\gamma(\alpha,x)$: $P_{\mathcal{G}}(x;\alpha)=\frac{\gamma(\alpha,x)}{\Gamma(\alpha)}\quad\text{where}\quad\gamma(\alpha,x)=\int_{0}^{x}t^{\alpha-1}e^{-t}\,\mathrm{d}t.$
$f_{\mathcal{G}}(x;\alpha)=\frac{1}{\Gamma(\alpha)}x^{\alpha-1}e^{-x}\quad\text{and}\quad P_{\mathcal{G}}(x;\alpha)=\frac{1}{\Gamma(\alpha)}\int_{0}^{x}t^{\alpha-1}e^{-t}\,\mathrm{d}t.$ (15)
Definition 4 (Shifted Poisson random variable).
The shifted Poisson (SP) variate with intensity $\lambda$ and shift $\alpha$, $N\sim\mathcal{SP}(\lambda,\alpha)$, takes non-negative integer values with mass probability function:
$P_{\mathcal{SP}}(n;\lambda,\alpha)=\mathrm{Prob}(N=n)=\frac{1}{P_{\mathcal{G}}(\lambda;\alpha)}\frac{\lambda^{\alpha+n}\,e^{-\lambda}}{\Gamma(n+\alpha+1)}.$ (16)
Remark 7 (Relation to the Poisson distribution).
When the shift parameter $\alpha$ is zero, the SP distribution is reduced to the Poisson distribution with intensity $\lambda$, $\mathcal{POIS}(\lambda)$, with mass probability function:
$P_{\mathcal{POIS}}(n;\lambda)=\frac{\lambda^{n}\,e^{-\lambda}}{n!}.$
Remark 8.
The SP variate arises from the series expansion of $P_{\mathcal{G}}(x;\alpha)$ (Abramowitz and Stegun, 1972, 6.5.29):
$1=\frac{1}{P_{\mathcal{G}}(\lambda;\alpha)}\frac{\gamma(\alpha,\lambda)}{\Gamma(\alpha)}=\frac{1}{P_{\mathcal{G}}(\lambda;\alpha)}\sum_{n=0}^{\infty}\frac{\lambda^{\alpha+n}\,e^{-\lambda}}{\Gamma(n+\alpha+1)}=\sum_{n=0}^{\infty}P_{\mathcal{SP}}(n;\lambda,\alpha).$
Next, we discuss the connection of the above defined distributions to the CEV distribution. It is well known that the transition density of $F_{T}$ is closely related to the NCX2 distribution (Schroder, 1989; Cox, 1996).
Proposition 5.
Define $z_{t}$ be a transformation of the CEV process, $F_{t}$, in Eq. (8) by the function $z(\cdot)$:
$z_{t}:=z(F_{t})\quad(0\leq t\leq T)\quad\text{and}\quad z(y):=\frac{y^{2\beta_{\ast}}}{\beta_{\ast}^{2}\sigma_{0}^{2}T}.$ (17)
The PDF of $z_{T}>0$ is expressed in terms of the NCX2 distribution:
$f_{z_{T}}(z;z_{0})\,\mathrm{d}z=\mathrm{Prob}(z_{T}\in[z,z+\mathrm{d}z])=f_{\chi^{2}}\left(z_{0};\frac{1}{\beta_{\ast}}+2,z\right)\mathrm{d}z\quad(z>0).$
Since the location variable $z$ plays the role of the noncentrality parameter, the distribution of $z_{T}$ is not a NCX2 distribution. The complementary CDF of $F_{T}>0$ and the probability of absorption (i.e., mass at zero) are respectively given by
$\mathrm{Prob}(F_{T}>y)$ $=\mathrm{Prob}(z_{T}>z(y))=P_{\chi^{2}}\left(z_{0};\frac{1}{\beta_{\ast}},z(y)\right),$ (18)
$\mathrm{Prob}(F_{T}=0)$ $=\mathrm{Prob}(z_{T}=0)=1-P_{\chi^{2}}\left(z_{0};\,\frac{1}{\beta_{\ast}},0\right)=1-P_{\mathcal{G}}\left(\frac{z_{0}}{2};\,\frac{1}{2\beta_{\ast}}\right).$ (19)
The last equality in the second line comes from Remark 6.
Proposition 6 (Mixture gamma representation of the CEV distribution (Makarov and Glew, 2010, § 3.3)).
Under the CEV model with absorbing boundary condition, given $z_{T}>0$ ($F_{t}$ is not absorbed at the origin until time $T$), the transition from $z_{0}$ to $z_{T}$ can be sampled by an SP-mixture gamma distribution:
$z_{T}\sim 2\,\mathcal{G}(N+1)\quad\text{where}\quad N\sim\mathcal{SP}\left(\frac{z_{0}}{2},\frac{1}{2\beta_{\ast}}\right)$
Proof.
We employ the infinite series expansion of $I_{\alpha}(\sqrt{z_{0}\,z_{T}})$ to obtain
$f_{\chi^{2}}\left(z_{0};\frac{1}{\beta_{\ast}}+2,z_{T}\right)$ $=\frac{1}{2}\left(\frac{z_{0}}{z_{T}}\right)^{\alpha/2}I_{\alpha}(\sqrt{z_{0}\,z_{T}})\;e^{-(z_{T}+z_{0})/2}\quad\left(\alpha=\frac{1}{2\beta_{\ast}}\right)$
$=\frac{1}{2}\left(\frac{z_{0}}{z_{T}}\right)^{\alpha/2}\sum_{k=0}^{\infty}\frac{(\sqrt{z_{0}\,z_{T}}/2)^{\alpha+2k}}{k!\,\Gamma(k+\alpha+1)}\;e^{-(z_{T}+z_{0})/2}$
$=\frac{1}{2}\sum_{k=0}^{\infty}\frac{(z_{0}/2)^{k+\alpha}e^{-z_{0}/2}}{\Gamma(k+\alpha+1)}\frac{(z_{T}/2)^{k}}{k!}\;e^{-z_{T}/2}$
$=P_{\mathcal{G}}\left(\frac{z_{0}}{2};\,\alpha\right)\sum_{k=0}^{\infty}P_{\mathcal{SP}}\left(k;\frac{z_{0}}{2},\alpha\right)\cdot\frac{1}{2}f_{\mathcal{G}}\left(\frac{z_{T}}{2};k+1\right).$
so that the transition density of $z_{T}$ conditional on $z_{T}>0$ is given by
$\frac{f_{z_{T}}(z_{T};z_{0})}{\mathrm{Prob}(z_{T}>0)}=\sum_{k=0}^{\infty}P_{\mathcal{SP}}\left(k;\frac{z_{0}}{2},\alpha\right)\cdot\frac{1}{2}f_{\mathcal{G}}\left(\frac{z_{T}}{2};k+1\right).$
The right hand side is the composite probability density of the gamma variable, $2\,\mathcal{G}(k+1)$, where $k$ is sampled from the SP variable, $\mathcal{SP}\left(z_{0}/2,1/2\beta_{\ast}\right)$. Therefore, $z_{T}>0$ can be sampled as $2\,\mathcal{G}(N+1)$. ∎
Proposition 7 (Sampling an SP variable (Kang, 2014, Algorithm 3)).
The SP variable, $N\sim\mathcal{SP}(\lambda,\alpha)$, can be sampled by a gamma-mixture Poisson variable:
$N\sim\mathcal{POIS}(\lambda-X)\quad\text{where}\quad X\sim\mathcal{G}(\alpha)\quad\text{conditional on}\quad X<\lambda.$
Proof.
For proof, see Kang (2014). ∎
Combining Propositions 6 and 7 leads to an exact simulation algorithm for sampling the CEV distribution. Moreover, Proposition 7 can be further optimized in the context of the CEV simulation. This is the reason we select ‘Algorithm 3’ among the three algorithms in Kang (2014). As this advantage is not mentioned in Kang (2014), we state explicitly it as a remark:
Remark 9 (Kang (2014)’s ‘Algorithm 3’ for the CEV distribution).
If simulating $N\sim\mathcal{SP}(\lambda,\alpha)$ is the final goal, $X<\lambda$ can be obtained by repeatedly sampling $X$ until $X<\lambda$. If $N$ is an intermediate variable for the CEV simulation in Proposition 6, however, the repeated sampling of $X$ is unnecessary since
$\mathrm{Prob}(X\geq\lambda)=1-P_{\mathcal{G}}\left(\frac{z_{0}}{2};\,\frac{1}{2\beta_{\ast}}\right)=\mathrm{Prob}(F_{T}=0).$
Therefore, we put $F_{T}=0$ if $X\geq\lambda$.
In summary, an exact simulation of the CEV distribution consists of successive generation of elementary random variables, for which efficient numerical algorithms are available in standard numerical library. Finally, the simulation algorithm for $F_{T}\sim\mathcal{CEV}_{\beta}(F_{0},\sigma_{0}^{2}T)$ is summarized in Algorithm 1:
Input: $F_{0}$, $\sigma_{0}$, $\beta$, $T$
Output: $F_{T}$
$\alpha\leftarrow\frac{1}{2\beta_{\ast}}$ and $z_{0}\leftarrow\frac{F_{0}^{2\beta_{\ast}}}{\beta_{\ast}^{2}\sigma_{0}^{2}T}$;
$X\leftarrow\mathcal{G}\left(\alpha\right)$;
if $X\geq z_{0}/2$ then
       return $F_{T}\leftarrow 0$;
else
       $z_{T}\leftarrow 2\,\mathcal{G}(\mathcal{POIS}(z_{0}/2-X)+1)$;
       return $F_{T}\leftarrow\left(\beta_{\ast}^{2}\sigma_{0}^{2}Tz_{T}\right)^{1/2\beta_{\ast}}$;
end if
Algorithm 1 Exact simulation of the CEV distribution $F_{T}$ in Eq. (8)
The CEV sampling algorithm can be applied to the conditional distribution of $F_{t+h}$ in Proposition 4. We need to sample $z_{t+h}=z(F_{t+h})$ from $z_{t}=z(\bar{F}_{t}^{h})$, where the transformation $z(\cdot)$ is modified in the context of Proposition 4:
$z(y)=\frac{y^{2\beta_{\ast}}}{\beta_{\ast}^{2}\rho_{\ast}^{2}\sigma_{t}^{2}hI_{t}^{h}}.$
We summarize the simulation of $F_{t}$ in the context of the SABR model in Algorithm 2:
Input: $F_{0}$, $\sigma_{0}$, $\nu$, $\beta$, $\rho$, $T$, $h$
Output: $F_{T}$
$\alpha\leftarrow\frac{1}{2\beta_{\ast}}$, $t\leftarrow 0$, $F_{t}\leftarrow F_{0}$, and $\sigma_{t}\leftarrow\sigma_{0}$;
repeat
       $\sigma_{t+h}\leftarrow$ Eq. (2) ;
       $I_{t}^{h}\leftarrow$ Proposition 3 ;
       $\bar{F}_{t}^{h}\leftarrow$ Eq. (11) ;
       $z_{t}\leftarrow\frac{\left(\bar{F}_{t}^{h}\right)^{2\beta_{\ast}}}{\beta_{\ast}^{2}\rho^{2}\sigma_{t}^{2}hI_{t}^{h}}$;
       $X\leftarrow\mathcal{G}\left(\alpha\right)$ ;
       if $X\geq z_{t}/2$ then
             $F_{T}\leftarrow 0$ ;
             break ;
      else
             $z_{t+h}\leftarrow 2\,\mathcal{G}(\mathcal{POIS}(z_{t}/2-X)+1)$;
             $F_{t+h}\leftarrow\left(\beta_{\ast}^{2}\rho^{2}\sigma_{t}^{2}hI_{t}^{h}z_{t+h}\right)^{1/2\beta_{\ast}}$;
       end if
      $t\leftarrow t+h$ ;
until t = T;
return $F_{T}$ ;
Algorithm 2 Simulation of $F_{t}$ under the SABR model
4.3 Comparison to Islah (2009)’s approximation used in other simulation schemes
While earlier SABR simulation methods show various approach for sampling $I_{t}^{h}$ (Step 1), almost all of them (Chen et al., 2012; Cai et al., 2017; Leitao et al., 2017a, b; Grzelak et al., 2019; Cui et al., 2021; Kyriakou et al., 2023) are based on Islah (2009)’s approximation for sampling $F_{t+h}$ (Step 2). We compare our approach to Islah (2009)’s in this section.
Proposition 8 (Islah (2009)’s approximation).
For $10<\beta<$ 1, conditional on $\sigma_{t+h}$ and $I_{t}^{h}$ (and the filtration up to time $t$), the distribution of $F_{t+h}$ is approximated by
$\mathrm{Prob}(F_{t+h}\geq y)\approx P_{\chi^{2}}\left(z^{\prime}_{t};\frac{1-\beta_{\ast}\rho^{2}}{\beta_{\ast}\rho_{\ast}^{2}},z(y)\right),$ (20)
where
$z^{\prime}_{t}=\frac{1}{\beta_{\ast}^{2}\rho_{\ast}^{2}\sigma_{t}^{2}hI_{t}^{h}}\left[F_{t}+\frac{\beta_{\ast}\rho}{\nu}(\sigma_{t+h}-\sigma_{t})\right]^{2}\quad\text{and}\quad z(y)=\frac{y^{2\beta_{\ast}}}{\beta_{\ast}^{2}\rho_{\ast}^{2}\sigma_{t}^{2}hI_{t}^{h}}.$
Proof.
For the ideas behind this approximation, see Islah (2009, § 2.2), Chen et al. (2012, Result 2.4), or Cui et al. (2021, Appendix B.1)
Islah’s approximation is compared with the probability under our CEV approximation in Proposition 4:
$\mathrm{Prob}(F_{t+h}\geq y)\approx P_{\chi^{2}}\left(z_{t};\frac{1}{\beta_{\ast}},z(y)\right)\quad\text{with}\quad z_{t}=z\left(\bar{F}_{t}^{h}\right).$
The key difference between the two is that Islah’s approximation uses the modified degree of freedom, $(1-\beta_{\ast}\rho^{2})/(\beta_{\ast}\rho_{\ast}^{2})$, in Eq. (20), and it is not equal to $1/\beta_{\ast}$ (except $\rho=0$). However, the transformation $z(\cdot)$ uses $\beta_{\ast}$ inconsistently. On the other hands, our CEV approximation uses $1/\beta_{\ast}$, which is the same degree of freedom for the CEV distribution. This means that $F_{t+h}$ under Islah’s approximation is not a CEV distribution, which leads to two drawbacks discussed in the next two remarks.
Remark 10.
As $\hat{\nu}\downarrow 0$, $F_{t+h}$ under Islah’s approximation does not converge to the CEV distribution with $\beta$ if $\rho\neq 0$. This violates the limiting property that the SABR model should converge to the CEV model with $\beta$ as $\nu\downarrow 0$.
Remark 11.
The martingale property of $F_{t}$ is not guaranteed under Islah’s approximation: $F_{t}\neq E(F_{t+h}).$ In particular, when $\beta\downarrow 0$ ($\beta_{\ast}\uparrow 1$), we can show that $F_{t}$ is a supermartingale: $F_{t}\leq E(F_{t+h})$. In the limit, $(1-\beta_{\ast}\rho^{2})/(\beta_{\ast}\rho_{\ast}^{2})$, as well as $1/\beta_{\ast}$, approaches to 1, and the conditional expectation is given by
$E\left(F_{t+h}\,|\,\sigma_{t+h},I_{t}^{h}\right)=z^{-1}\left(z_{t}^{\prime}\right)=\left|F_{t}+\frac{\rho}{\nu}(\sigma_{t+h}-\sigma_{t})\right|.$
Taking expectation on both sides over the joint distribution of $(\sigma_{t+h},I_{t}^{h})$, we arrive at
$F_{t}=E\left(F_{t}+\frac{\rho}{\nu}(\sigma_{t+h}-\sigma_{t})\right)\leq E\left(\left|F_{t}+\frac{\rho}{\nu}(\sigma_{t+h}-\sigma_{t})\right|\right)=E\left(E\left(F_{t+h}\,|\,\sigma_{t+h},I_{t}^{h}\right)\right)=E(F_{t+h}).$
In order to preserve the martingale property, Leitao et al. (2017a, b) artificially add an additive adjustment, commonly called the martingale correction, to the simulated samples of $F_{t+h}$. This ad-hoc correction is inferior to our approach where the martingale condition is naturally observed.
Apart from the failure of martingale property in Islah’s approximation, a fast and exact algorithm for sampling $F_{t+h}$ according to Eq. (20) has not been available in the literature. The methods presented so far are either slow (but accurate) or inaccurate (but fast). Chen et al. (2012) (partially) and Cai et al. (2017) adopted numerical root finding of $P_{\chi^{2}}(\,\cdot\,,\,\cdot\,,z_{K})=U$ for a uniform random number $U$, which belongs to the former type. This approach is slow due to the cumbersome evaluations of $P_{\chi^{2}}$ despite using some good initial guess and efficient calculation algorithm. Grzelak et al. (2019) speeded up sampling by using the stochastic collocation method, which is an optimal caching of the distribution function. Chen et al. (2012) (partially) and Cui et al. (2021) adopted the moment-matched quadratic Gaussian approximation to Eq. (20), which is proven to be effective in the Heston simulation (Andersen, 2008). While fast, it becomes inaccurate for a larger time step as it only uses the first two moments. This approach belongs to the latter type.
It is interesting to note that the exact CEV sampling algorithm in Section 4.2 can also be modified to sample $F_{t+h}$ under Islah’s approximation though it is not recommended due to the inherent drawbacks of Islah’s approximation. We outline the procedure in B. With the use of the exact CEV sampling procedure, we numerically illustrate the supermartingale property of $F_{t}$ in Section 5.
5 Numerical Results
We demonstrate the accuracy and effectiveness of our simulation algorithm with an extensive set of numerical examples. Table 1 shows the five parameter sets used in our numerical tests. For easier comparison to other existing methods, we select the sets from previous literature: the first two from Antonov and Spector (2012) and the rest from Cai et al. (2017). Following the literature (Chen et al., 2012; Cai et al., 2017), we price European options. Although the versatility of our simulation scheme goes beyond prcing European options, it offers a good testing case. Pricing European options also benefits from the availability of highly accurate benchmark option prices obtained from using the finite difference method (FDM) (Cai et al., 2017).
Cases $F_{0}$ $\sigma_{0}$ $\nu$ $\rho$ $\beta$ $T$ $K$
Case I 1 0.25 0.3 $-0.8$ 0.3 10 [0.2, 2.0]
Case II 1 0.25 0.3 $-0.5$ 0.6 10 [0.2, 2.0]
Case III 0.05 0.4 0.6 0.0 0.3 1 [0.02, 0.10]
Case IV 1.1 0.4 0.8 $-0.3$ 0.3 4 1.1
Case V 1.1 0.3 0.5 $-0.8$ 0.4 [1, 10] 1.1
Table 1: Parameter sets used in our numerical experiments.
In our numerical tests, unless otherwise specified, we used $N=10^{5}$ paths for each simulation, and repeated $m=50$ times. Out of $m$ prices, we computed the average bias (used the FDM price as the benchmark) and standard deviation. Our algorithm is implemented in MATLAB R2023b on a laptop with a 13th Gen Intel Core i7-1360P 2.20 GHz processor. In Section 5.2, to demonstrate the speed of algorithm, we compare the CPU time of our method with those reported in Cai et al. (2017). Cai et al. (2017) used Matlab 7 on an Intel Core2 Q9400 2.66GHZ processor. One may argue that comparing CPU time in this manner may not be fair as we used a newer system. However, the two CPUs are not significantly different to overturn the conclusion. As shown, our algorithm is faster than the existing algorithms by one or two orders of magnitude.
5.1 Accuracy and comparison to analytic approximations
In this subsection, we test the accuracy of our algorithm (Algorithm 2) using Cases I and II in Table 1. Tables 2 and 3 show the numerical results for the two sets. The two examples are previously tested in Antonov and Spector (2012) to demonstrate the performance of the cutting-edge analytic approximations for European vanilla options, map to the zero correlation (ZC Map, and hybrid map to the zero correlation (Hyb ZC Map), against the widely used Hagan et al. (2002)’s implied volatility formula. [^4] Since these approximation formulas are all expressed in terms of implied volatilities, we use the Black-Scholes formula to obtain the call option price. Antonov and Spector (2012) use MC method for the benchmark option price. Instead, we use FDM price for the benchmark, which is more precise. Therefore, we also display those option prices for comparison.
Tables 2 and 3 demonstrate that our simulation scheme is highly accurate. As we decrease the time step $h$ from 1 to $1/16$, the option price converges to its true price. Even for a large time step ($h=1$), the option prices are accurate almost up to three decimal points. Such bias level is comparable to the advanced analytic methods. Overall speaking, we note that the bias in Table 3 (Case II) is lower than that in Table 2 (Case I). This is probably due to the observation that the geometric BM approximation for the conditional mean $\bar{F}_{0}^{T}$ in Eq. (12) holds better as $\beta=0.6$ (Case II) is closer to one than $\beta=0.3$ (Case I).
$K/F_{0}$ 0.2 0.4 0.8 1.0 1.2 1.6 2.0
Exact option price from FDM
FDM 0.84255 0.68906 0.40646 0.28502 0.18304 0.05343 0.01096
Bias of our simulation method  (×$(\times 10^{-3})$
$h=1$1.22 -1.49 -0.37 0.49 1.28 1.72 1.32
$h=1/4$0.46 -0.24 0.22 0.42 0.56 0.56 0.48
$h=1/16$0.34 -0.20 0.00 0.05 0.11 0.10 0.10
Standard deviation of our simulation method  (×$(\times 10^{-3})$
$h=1$ 1.97 1.83 1.50 1.31 1.08 0.63 0.38
$h=1/4$ 1.96 1.73 1.29 1.08 0.91 0.61 0.41
$h=1/16$ 1.89 1.75 1.44 1.28 1.06 0.53 0.22
Bias of various analytic approximation methods  (×$(\times 10^{-3})$
Hagan 22.35 23.64 17.94 13.81 9.38 2.54 0.82
ZC Map 0.37 0.51 1.39 2.29 3.20 4.02 2.66
Hyb ZC Map 4.64 5.81 4.07 2.29 0.43 -1.26 -0.30
Table 2: The European option prices from our simulation methods for Case I, compared with those from various analytic approximation methods (Hagan, ZC Map, and Hyb ZC Map) reported in Antonov and Spector (2012). The errors are measured against the benchmark option prices from the finite difference method (FDM). We used $N=100,000$ paths in each simulation run and repeated $m=50$ times. The average computation time is 0.8, 1.6, and 14.5 seconds for $h=1$, $1/4$, and $1/16$, respectively.
$K/F_{0}$ 0.2 0.4 0.8 1.0 1.2 1.6 2
Exact option price from FDM
FDM 0.82886 0.66959 0.39772 0.29118 0.20690 0.10018 0.05014
Bias of our simulation method   (×$(\times 10^{-3})$
$h=1$0.14 -0.30 -0.42 -0.43 -0.43 -0.40 -0.30
$h=1/4$ 0.45 0.37 0.27 0.20 0.10 -0.02 0.00
$h=1/16$ 0.01 -0.01 0.02 0.04 0.03 0.00 -0.03
Stdev of our simulation method   (×$(\times 10^{-3})$
$h=1$ 2.23 2.09 1.78 1.65 1.51 1.20 0.93
$h=1/4$ 2.21 2.10 1.85 1.70 1.51 1.14 0.88
$h=1/16$ 2.46 2.32 2.01 1.79 1.58 1.22 0.97
Bias of various analytic approximation methods   (×$(\times 10^{-3})$
Hagan 11.65 15.94 16.69 14.67 12.18 8.56 6.88
ZC Map -1.56 -1.57 0.56 2.37 4.02 5.77 5.33
Hyb ZC Map 3.07 4.53 3.86 2.37 1.00 -0.10 0.52
Table 3: The European option prices from our simulation methods for Case II, compared with those from various analytic approximation methods (Hagan, ZC Map, and Hyb ZC Map) reported in Antonov and Spector (2012). The errors are measured against the benchmark option prices from the finite difference method (FDM). We used $N=100,000$ paths in each simulation run and repeated $m=50$ times.
5.2 Accuracy and speed trade-off
In this subsection, we demonstrate the computational efficiency of our method by comparing the accuracy-speed trade-off with the earlier simulation methods, such as the Euler, low-bias (Chen et al., 2012), and piecewise semi-exact (PSE) (Cai et al., 2017) schemes. For the purpose, we use Cases III and IV. These cases have been tested by Cai et al. (2017), and we take advantage of the extensive records of the CPU time reported therein.
$K/F_{0}$ 0.4 0.8 1 1.2 1.6 2 Time (s)
Exact option price from FDM
FDM 0.04559 0.04141 0.03942 0.03750 0.03390 0.03061
Error (×$(\times 10^{-3})$ and CPU time of our simulation method
$h=1$ 0.00 0.00 0.00 0.00 -0.01 -0.01 0.03
Bias (×$(\times 10^{-3})$ and CPU time of the Euler scheme reported in Cai et al. (2017)
$h=1/400$ 1.6 1.5 1.5 1.4 1.3 1.2 49.4
$h=1/800$ 0.7 0.6 0.5 0.5 0.4 0.3 99.1
$h=1/1600$0.3 -0.3 -0.3 -0.3 -0.3 -0.3 194.0
Bias (×$(\times 10^{-3})$ and CPU time of the low-bias scheme reported in Cai et al. (2017)
$h=1/4$ 0.5 0.5 0.5 0.4 0.4 0.4 78.4
$h=1/8$ 0.4 0.4 0.4 0.3 0.3 0.2 175.8
Bias (×$(\times 10^{-3})$ and CPU time of the PSE scheme reported in Cai et al. (2017)
$h=1$ 0.1 0.2 0.2 0.2 0.2 0.2 98.3
Table 4: The European option prices from our simulation methods for Case III, compared with other MC methods: Euler, low-bias (Chen et al., 2012), and PSE (Cai et al., 2017) schemes. We used the price and computation time reported in Cai et al. (2017, Table 7). The biases are measured against the benchmark option prices from the finite difference method (FDM). All simulation methods used $N=100,000$ paths in each simulation run and repeated $m=50$ times.
The results for Case III in Table 4 show that our method is at least several hundred times faster than the competitors while the bias is much lower. Note that the correlation is zero in this case. When $\rho=0$, both our CEV approximation and Islah’s approximation become exact, and the source of bias is the accuracy of sampling $I_{t}^{h}$. The near-zero error of our method validates our SLN approximation for $I_{t}^{h}$. For the Euler and low-bias schemes, it is necessary to decrease the time step $h$ for more accurate sampling of $I_{t}^{h}$, resulting with slower computation. Although the PSE scheme exhibits good accuracy for single step ($h=1$), it requires prohibitive computation for numerical inverse transform for sampling $I_{t}^{h}$.
$N$ $h$ RMS error (×$\times 10^{-3}$) Time (s)
160,000 1 3.27 0.53
320,000 1/2 1.94 2.27
640,000 1/4 1.21 9.66
1,280,000 1/8 0.86 41.35
2,560,000 1/16 0.59 279.53
Table 5: The root mean square (RMS) error and CPU time from Case IV when the number of paths $N$ is doubled and time step $h$ is halved.
Next, we perform CPU time versus root-mean-square (RMS) error analysis on Case IV. The RMS error is defined by
RMS $\text{RMS error}=\sqrt{\text{Stdev}^{2}+\text{Bias}^{2}}.$ (21)
Table 5 presents the RMS error and CPU time for decreasing time step $h$ and increasing number of paths $N$. Similar to Cai et al. (2017, Table 9), we set $h\propto 1/N$. In each successive row, we double $N$ and halve $h$ to generate data points for the plots of RMS versus simulation time.
Figure 2: RMS error as a function of CPU time for simulation methods. The data points for our simulation scheme are from Table 5 while those for other methods are from Cai et al. (2017, Table 9 and Figure 6).
Based on the data points from Table 5, Figure 2 shows the plots of the trade-off between CPU time and RMS error. The figure explicitly demonstrates that, while the decay is slightly slower than the competing schemes, our simulation method achieves the same degree of RMS error using much less computation time, only about one hundredth of the Euler scheme. Thanks to the high efficiency for small time step $h$, our algorithm is also effective for evaluating path-dependent options with frequent discrete monitoring.
5.3 Comparison with Islah’s approximation
In this last numerical experiment using Case V, we demonstrate superiority of our CEV approximation of conditional $F_{T}$ in Section 4.1 to Islah’s approximation which has been commonly adopted in other schemes. The parameter set is also taken from Cai et al. (2017, Figure 4). We price the at-the-money European option for maturity, $T=1,2,\ldots,10$ with time step $h=1/2$ or 1. On one hand, our simulation scheme consists of (i) SLN sampling of $I_{t}^{h}$ and (ii) CEV model sampling of conditional $F_{T}$. On the other hand, we replace the second step with Islah’s approximation (Proposition 8).
Under Islah’s approximation, a power transformation of $F_{T}$ observes a CEV distribution; see Eq. (22). Therefore, Kang (2014)’s algorithm also provides a better way to simulate Islah’s approach. Accordingly, we use Algorithm 1 with $\beta$, $F_{0}$, and $\sigma_{0}^{2}T$ replaced by $\beta^{\prime}$, $\bar{F}_{T}^{\prime}$, and $\rho_{\ast}^{2}\sigma_{0}^{2}TI_{0}^{T}$ in Eq. (22), respectively, to sample $\left((\beta_{\ast}^{\prime}/\beta_{\ast})F_{T}^{\beta_{\ast}}\right)^{1/\beta_{\ast}^{\prime}}$. Then, $F_{T}$ is finally obtained.
(a) Average terminal price $F_{T}$
(b) European call option price
Figure 3: Comparison of Islah’s approximation and our conditional distribution of $F_{T}$
Figure 3 compares the results from the two approaches: our conditional distribution of $F_{T}$ and Islah’s approximation. Figure 3(a) shows the error of the terminal price $E(F_{T})$, which should be equal to $F_{0}=1.1$ in theory. As expected, the results show that our approach preserves the martingale property fairly well for all time-to-maturity $T$. That is, $E(F_{T})=F_{0}$ holds with very high accuracy regardless of $h$. In Islah’s approach, however, $E(F_{T})$ deviates from $F_{0}$. In particular, the deviation accumulates as the time-to-maturity $T$ becomes longer. The time step $h$ has to be smaller to reduce the deviation.
Figure 3(b) shows the error of the ATM option price from its true value computed with the FDM. Our approach is more accurate than Islah’s approach. Note that the higher option price error in Islah’s approach is due to failure of the martingale property. As the delta of ATM option is about 0.5, the option price error is approximately half of the forward price error. Conversely, the martingale preservation of our CEV approximation indeed improves European call option pricing. The option price error in our approach is mostly attributed to the error in $I_{t}^{h}$ distribution. The underpricing of option seems to be related to that the ex-kurtosis of $I_{t}^{h}$ sampled from the SLN approach, which is slightly smaller than that of true distribution (see Figure 1). The error is reduced when a smaller $h$ is used.
From Figure 3(b), we observe that under the same approach to sample $I_{0}^{T}$, and same size of time step $h$, the results using our CEV distribution of $F_{T}$ are very close to the benchmark FDM results. However, the results of Islah’s approximation have larger numerical bias as $T$ gets longer. These results illustrate that our CEV approximation of the conditional $F_{T}$ is superior to Islah’s approximation.
6 Conclusion
The stochastic-alpha-beta-rho (SABR) model proposed by Hagan et al. (2002) has been widely adopted for pricing derivatives under the stochastic volatility framework. The corresponding simulation methods have been studied extensively. The naive time-discretization schemes, such as the Euler or Milstein schemes, suffer from non-negligible bias caused by the truncation of negative forward values even when a small time step is used. Although researchers proposed various simulation schemes that overcome the limitation of the time-discretization schemes, the schemes have their own limitations. In sampling conditional average variance, the schemes are riddled with either heavy computation (Cai et al., 2017; Leitao et al., 2017a) or inaccurate approximation over a large simulation step (Chen et al., 2012). In sampling conditional forward price, almost all methods adopt Islah (2009)’s NCX2 approximation that neither preserves martingale property nor provides efficient sampling algorithm.
This paper present a series of innovations on the SABR simulation. Firstly, we derive the first four moments of the conditional average variance analytically, and sample the conditional average variance via an SLN random variable fitted to the first three moments. Our sampling procedure is accurate for reasonably large time-to-maturity and fast in computation in comparison to the numerical inverse transform methods. Secondly, we propose a martingale-preserving CEV approximation for the conditional forward price that preserves the martingale property. Lastly, we sample the conditional forward price efficiently with the exact CEV sampling algorithm of Makarov and Glew (2010) and Kang (2014). Numerical results show that our methods are highly efficient and accurate even under challenging parameter values, like large time-to-maturity $T$, when compared with other simulation methods.
Funding
The works of Lilian Hu and Yue Kuen Kwok are supported by the Guangzhou-HKUST (GZ) Joint Funding Program (Number 2024A03J0630).
Appendix A Conditional moments of $I_{t}^{h}$
The conditional average variance $I_{t}^{h}(\hat{Z})$ is closely related to the exponential functional of BM, defined by
$A_{t}:=\int_{0}^{t}e^{2Z_{s}}\,\mathrm{d}s,$
where $Z_{s}$ is a standard Brownian motion. The derivation of Proposition 2 is based on the analytic conditional moments of the exponential functional of BM (Matsumoto and Yor, 2005, (5.4)):
$E\left(\left(A_{t}\right)^{n}|Z_{t}=x\right)=\frac{e^{nx}}{n!t}\int_{x}^{\infty}be^{(x^{2}-b^{2})/2t}\left[\cosh(b)-\cosh(x)\right]^{n}\mathrm{d}b.$
The conditional average variance $I_{t}^{h}(\hat{Z})$ is expressed by $A_{t}$ conditional on the terminal location of the BM, $Z_{t}$:
$I_{t}^{h}(\hat{Z})=\frac{1}{\hat{\nu}^{2}}A_{\hat{\nu}^{2}}\,\Big{|}_{Z_{\hat{\nu}^{2}}=\hat{\nu}\hat{Z}}\,\quad\left(\hat{\nu}=\nu\sqrt{h}\right).$
Therefore, the conditional moments admit the following integral form:
$\mu^{\prime}_{k}$ $=E\left(\left(I_{t}^{h}\right)^{k}\right)=\frac{1}{\hat{\nu}^{2k}}E\left[\left(A_{\hat{\nu}^{2}}\right)^{k}\,\big{|}_{Z_{\hat{\nu}^{2}}=\hat{\nu}\hat{Z}}\right]$
$=\frac{e^{k\hat{\nu}\hat{Z}}}{k!\,\hat{\nu}^{2k}}\int_{\hat{Z}}^{\infty}se^{(\hat{Z}^{2}-s^{2})/2}\left[\cosh(\hat{\nu}s)-\cosh(\hat{\nu}\hat{Z})\right]^{k}\mathrm{d}s$
$=\frac{e^{k\hat{\nu}\hat{Z}}}{(k-1)!\,\hat{\nu}^{2k-1}}\int_{\hat{Z}}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\sinh(\hat{\nu}s)\left[\cosh(\hat{\nu}s)-\cosh(\hat{\nu}\hat{Z})\right]^{k-1}\mathrm{d}s,$
where we apply the change of variable, $s=\hat{\nu}b$, to obtain the second line, and integration by part to obtain the third line.
Using the following analytic integral
$\int_{z}^{\infty}e^{-s^{2}/2}\sinh(as)\,\mathrm{d}s=\frac{N(-z+a)-N(-z-a)}{2n(a)}=\frac{N(z+a)-N(z-a)}{2n(a)},$
we obtain for $k\neq 0$,
$\int_{\hat{Z}}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\sinh(k\hat{\nu}s)\,\mathrm{d}s=k\hat{\nu}\,m_{k}(\hat{Z})\quad\text{where}\quad m_{k}(\hat{Z}):=\frac{N(\hat{Z}+k\hat{\nu})-N(\hat{Z}-k\hat{\nu})}{2k\hat{\nu}\;n\left(\sqrt{\hat{Z}^{2}+(k\hat{\nu})^{2}}\right)},$
which will be useful in the derivation below. Here, we define $m_{k}(\hat{Z})$ in such a way that $m_{k}(\hat{Z})\to 1$ for any non-zero $k$ as $\hat{\nu}\to 0$.
Next, we evaluate $E((I_{t}^{h})^{k})$. For $k=1$, we obtain
$\mu=\frac{e^{\hat{\nu}\hat{Z}}}{\hat{\nu}}\int_{\hat{Z}}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\sinh(\hat{\nu}s)\,\mathrm{d}s=e^{\hat{\nu}\hat{Z}}\,m_{1}(\hat{Z})=\left(\frac{\sigma_{t+h}}{\sigma_{t}}\right)\,m_{1}.$
For $k=2$, using $\sinh(x)\cosh(x)=\frac{1}{2}\sinh(2x)$, we have
$\mu^{\prime}_{2}$ $=\frac{e^{2\hat{\nu}\hat{Z}}}{\hat{\nu}^{3}}\int_{z}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\sinh(\hat{\nu}s)\left[\cosh(\hat{\nu}s)-\cosh(\hat{\nu}z)\right]\,\mathrm{d}s$
$=\frac{e^{2\hat{\nu}\hat{Z}}}{\hat{\nu}^{3}}\int_{z}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\left[\frac{\sinh(2\hat{\nu}s)}{2}-\sinh(\hat{\nu}s)\cosh(\hat{\nu}z)\right]\,\mathrm{d}s$
$=\left(\frac{\sigma_{t+h}}{\sigma_{t}}\right)^{2}\frac{1}{\hat{\nu}^{2}}\left[m_{2}-cm_{1}\right].$
For $k=3$, using $\sinh(x)\cosh^{2}(x)=\frac{1}{4}\left[\sinh(3x)+\sinh(x)\right]$, we obtain
$\mu^{\prime}_{3}$ $=\frac{e^{3\hat{\nu}\hat{Z}}}{2\,\hat{\nu}^{5}}\int_{z}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\sinh(\hat{\nu}s)\left[\cosh(\hat{\nu}s)-\cosh(\hat{\nu}z)\right]^{2}\,\mathrm{d}s$
$=\frac{e^{3\hat{\nu}\hat{Z}}}{2\,\hat{\nu}^{5}}\int_{z}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\left[\frac{\sinh(3\hat{\nu}s)+\sinh(\hat{\nu}s)}{4}-\sinh(2\hat{\nu}s)\cosh(\hat{\nu}z)+\sinh(\hat{\nu}s)\cosh^{2}(\hat{\nu}z)\right]\,\mathrm{d}s$
$=\left(\frac{\sigma_{t+h}}{\sigma_{t}}\right)^{3}\frac{1}{8\hat{\nu}^{4}}\left[3m_{3}-8cm_{2}+\left(4c^{2}+1\right)m_{1}\right].$
For $k=4$, using $\sinh(x)\cosh^{3}(x)=\frac{1}{8}\left[\sinh(4x)+2\sinh(2x)\right]$, we have
$\mu^{\prime}_{4}$ $=\frac{e^{4\hat{\nu}\hat{Z}}}{6\,\hat{\nu}^{7}}\int_{z}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\sinh(\hat{\nu}s)\left[\cosh(\hat{\nu}s)-\cosh(\hat{\nu}z)\right]^{3}\,\mathrm{d}s$
$=\frac{e^{4\hat{\nu}\hat{Z}}}{6\,\hat{\nu}^{7}}\int_{z}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\left\{\frac{\sinh(4\hat{\nu}s)+2\sinh(2\hat{\nu}s)}{8}-\frac{3}{4}\left[\sinh(3\hat{\nu}s)+\sinh(\hat{\nu}s)\right]\cosh(\hat{\nu}z)\right.$
$\hskip 120.00018pt\left.+\,\frac{3}{2}\sinh(2\hat{\nu}s)\cosh^{2}(\hat{\nu}z)-\sinh(\hat{\nu}s)\cosh^{3}(\hat{\nu}z)\right\}\,\mathrm{d}s$
$=\left(\frac{\sigma_{t+h}}{\sigma_{t}}\right)^{4}\frac{1}{24\,\hat{\nu}^{6}}\left[2m_{4}-9cm_{3}+\left(12c^{2}+2\right)m_{2}-c\left(4c^{2}+3\right)m_{1}\right].$
In addition, the expansion of the mean, coefficient of variance, skewness, and ex-kurtosis (see Remark 5) around $\hat{\nu}=0$ are obtained as
$v=\frac{\hat{\nu}}{\sqrt{3}}+O(\hat{\nu}^{3}),\quad s=\frac{6\sqrt{3}}{5}\hat{\nu}+O(\hat{\nu}^{3}),\quad\text{and}\quad\kappa=\frac{276}{35}\hat{\nu}^{2}+O(\hat{\nu}^{4}).$
Appendix B Sampling Islah’s approximation via a CEV distribution
Though Islah’s approximation is less than desirable, we show that Islah’s distribution of $F_{t+h}$ can be expressed as a power of a CEV random variable. Therefore, the CEV simulation algorithm in Section 4.2 can be applied to expedite the sampling.
Let us define
$\beta^{\prime}:=\frac{\beta}{1-\beta_{\ast}\rho^{2}}\;(\geq\beta),\quad\beta_{\ast}^{\prime}:=1-\beta^{\prime}=\frac{\beta_{\ast}\rho_{\ast}^{2}}{1-\beta_{\ast}\rho^{2}}\;(\leq\beta_{\ast}),$
$y^{\prime}:=\left(\frac{\beta_{\ast}^{\prime}}{\beta_{\ast}}y^{\beta_{\ast}}\right)^{1/\beta_{\ast}^{\prime}},\quad\text{and}\quad\bar{F^{\prime}}_{t}^{h}:=\left|\frac{\beta_{\ast}^{\prime}}{\beta_{\ast}}\left(F_{t}^{\beta_{\ast}}+\frac{\beta_{\ast}\rho}{\nu}(\sigma_{t+h}-\sigma_{t})\right)\right|^{1/\beta_{\ast}^{\prime}}$
so that
$\frac{(y^{\prime})^{2\beta_{\ast}^{\prime}}}{(\beta_{\ast}^{\prime})^{2}}=\frac{y^{2\beta_{\ast}}}{\beta_{\ast}^{2}}\quad\text{and}\quad\frac{\left(\bar{F^{\prime}}_{t}^{h}\right)^{2\beta_{\ast}^{\prime}}}{\beta_{\ast}^{\prime 2}}=\frac{1}{\beta_{\ast}^{2}}\left[F_{t}^{\beta_{\ast}}+\frac{\beta_{\ast}\rho}{\nu}(\sigma_{t+h}-\sigma_{t})\right]^{2}.$
As a result, the complementary CDF of Islah (2009)’s approximation is equivalently expressed by
$\mathrm{Prob}(F_{t+h}\geq y)=\mathrm{Prob}\left(\left(\frac{\beta_{\ast}^{\prime}}{\beta_{\ast}}F_{t+h}^{\beta_{\ast}}\right)^{1/\beta_{\ast}^{\prime}}\geq y^{\prime}\right)=P_{\chi^{2}}\left(\frac{\left(\bar{F^{\prime}}_{t}^{h}\right)^{2\beta_{\ast}^{\prime}}}{\beta_{\ast}^{\prime 2}\rho_{\ast}^{2}\sigma_{0}^{2}TI_{0}^{T}};\frac{1}{\beta_{\ast}^{\prime}},\frac{\left(y^{\prime}\right)^{2\beta_{\ast}^{\prime}}}{\beta_{\ast}^{\prime 2}\rho_{\ast}^{2}\sigma_{0}^{2}TI_{0}^{T}}\right),$
and this implies that
$\left(\frac{\beta_{\ast}^{\prime}}{\beta_{\ast}}F_{t+h}^{\beta_{\ast}}\right)^{1/\beta_{\ast}^{\prime}}\sim\mathcal{CEV}_{\beta^{\prime}}\left(\bar{F^{\prime}}_{t}^{h},\rho_{\ast}^{2}\sigma_{t}^{2}hI_{t}^{h}\right).$ (22)
In other words, a power of $F_{t+h}$ instead of $F_{t+h}$ itself, follows a CEV distribution under Islah’s approximation. Therefore, $F_{t+h}$ can be sampled as
$F_{t+h}\sim\left(\frac{\beta_{\ast}}{\beta_{\ast}^{\prime}}\right)^{1/\beta_{\ast}}Y^{\beta_{\ast}^{\prime}/\beta_{\ast}}\quad\text{where}\quad Y\sim\mathcal{CEV}_{\beta^{\prime}}\left(\bar{F^{\prime}}_{t}^{h},\rho_{\ast}^{2}\sigma_{t}^{2}hI_{t}^{h}\right).$
Lastly, we obtain
$E\left(F_{t+h}^{\beta_{\ast}/\beta_{\ast}^{\prime}}\,|\,\sigma_{t+h},I_{t}^{h}\right)=\left(\frac{\beta_{\ast}}{\beta_{\ast}^{\prime}}\right)^{1/\beta_{\ast}^{\prime}}E(Y)=\left(\frac{\beta_{\ast}}{\beta_{\ast}^{\prime}}\right)^{1/\beta_{\ast}^{\prime}}\bar{F^{\prime}}_{t}^{h}=\left|F_{t}^{\beta_{\ast}}+\frac{\beta_{\ast}\rho}{\nu}(\sigma_{t+h}-\sigma_{t})\right|^{1/\beta_{\ast}^{\prime}}.$

### v2

@@
-Efficient simulation of the SABR model
+Efficient and accurate simulation of the stochastic-alpha-beta-rho model
 Abstract
-We propose an efficient and reliable simulation scheme for the stochastic-alpha-beta-rho (SABR) model.
-The two challenges of the SABR simulation lie in sampling (i) the integrated variance conditional on terminal volatility and (ii) the terminal price conditional on terminal volatility and integrated variance.
-For the first sampling procedure, we analytically derive the first four moments of the conditional average variance, and sample it from the moment-matched shifted lognormal approximation.
-For the second sampling procedure, we approximate the conditional terminal price as a constant-elasticity-of-variance (CEV) distribution.
+We propose an efficient, accurate and reliable simulation scheme for the stochastic-alpha-beta-rho (SABR) model.
+The two challenges of the SABR simulation lie in sampling (i) integrated variance conditional on terminal volatility and (ii) terminal forward price conditional on terminal volatility and integrated variance.
+For the first sampling procedure, we sample the conditional integrated variance using the moment-matched shifted lognormal approximation.
+For the second sampling procedure, we approximate the conditional terminal forward price as a constant-elasticity-of-variance (CEV) distribution.
 Our CEV approximation preserves the martingale condition and precludes arbitrage, which is a key advantage over Islah’s approximation used in most SABR simulation schemes in the literature.
-Then, we adopt the exact sampling method of the CEV distribution based on the shifted-Poisson-mixture Gamma random variable.
+We then adopt the exact sampling method of the CEV distribution based on the shifted-Poisson mixture Gamma random variable.
 Our enhanced procedures avoid the tedious Laplace inversion algorithm for sampling integrated variance and non-efficient inverse transform sampling of the forward price in some of the earlier simulation schemes.
@@
 keywords:
-SABR model, Monte Carlo simulation, constant elasticity of variance process, shifted-Poisson-mixture gamma distribution
-journal: arXiv.org
+SABR model, constant elasticity of variance process, simulation, shifted-Poisson-mixture Gamma distribution
+journal: European Journal of Operational Research
 1 Introduction
 The stochastic-alpha-beta-rho (SABR) model proposed by Hagan et al.
-(2002) has been widely adopted in option pricing due to its ability to capture volatility smile or skew using a few parameters, and exhibiting consistency in revealing the dynamic behavior between price and smile.
-On one hand, it is common in equity options that the implied volatility of out-of-money options are generally higher than that of the in-the-money counterparts, known as the volatility skew.
+(2002) has been widely adopted in option pricing due to its ability to capture volatility smile or skew using a few parameters, and exhibiting consistency in revealing the dynamic behavior between asset price and volatility smile.
+On one hand, it is common in equity options that the implied volatility of out-of-money options is generally higher than that of the in-the-money counterparts, known as the volatility skew.
 On the other hand, the relatively symmetric volatility smile is common in the foreign exchange options.
@@
 Notably, Antonov and Spector (2012) obtained a more accurate approximation by mapping the implied volatilities under the correlated cases to those of the uncorrelated cases, under which more analytic properties of the SABR model become available.
-Another stream of research, to which this paper belongs, is developing efficient simulation algorithms of the SABR model.
+This paper follows the other stream of research, which is the development of efficient simulation algorithms for pricing options under the SABR model.
 With growing popularity of the SABR model, there has been demand for pricing path-dependent derivatives using the model.
-Since analytic expansion approaches are limited to pricing the European options, the Monte Carlo method becomes a natural choice.
-In particular, the research on the SABR simulation has been stimulated by less than perfect performance of the Euler and Milstein time-discretization schemes.
-Since the SABR model imposes an absorbing boundary condition at the origin, the time discretization scheme should be more carefully implemented; naive truncation at the origin leads to large deviation from the original SABR model.
-Typically, it is necessary to use a small time discretization step to decrease the discretization errors due to its low order of convergence, thus suffering from heavy computation load.
-Several simulation algorithms (Chen et al., 2012; Cai et al., 2017; Leitao et al., 2017a, b; Cui et al., 2018; Grzelak et al., 2019; Kyriakou et al., 2023) have been proposed for efficient simulation of the SABR model.
-Simulating the SABR process over a relatively large time step requires sequentially sampling terminal volatility, average variance, and terminal price at the next simulation time point.
-Since the last two sampling procedures are challenging, existing methods are differentiated with reference to different choices of the algorithms for these two steps.
-These earlier simulation algorithms have achieved some success in that they overcome the issue of the time-discretization method.
-However, there are still many gaps to be filled in order to make SABR simulation more efficient.
-Regarding the average variance conditional on the terminal volatility, Chen et al.
-(2012) sampled the quantity through the mean-and-variance matched lognormal (LN) variable.
+Since analytic approximation approaches are limited to pricing European options, the Monte Carlo method becomes a better choice.
+In particular, the research on the SABR simulation started with the less than perfect performance of the Euler and Milstein time-discretization schemes.
+Since the SABR model imposes an absorbing boundary condition at the origin, the numerical boundary condition in the time discretization schemes have to be more carefully implemented.
+It has been well known that naive truncation at the origin may lead to large deviation from the original SABR model.
+Typically, it is necessary to use a small time discretization step to decrease the discretization errors due to its low order of convergence, thus causing heavy computation load.
+Several simulation algorithms (Chen et al., 2012; Cai et al., 2017; Leitao et al., 2017a, b; Cui et al., 2018; Grzelak et al., 2019; Kyriakou et al., 2023) have been proposed for the simulation of the SABR model.
+Simulation of the SABR model over one time step requires sequential sampling procedures of the terminal volatility, average variance, and terminal forward price at the next simulation time point.
+Since the last two sampling procedures are challenging, existing methods distinguish among themselves by adopting different choices of the algorithms for these two steps.
+These earlier simulation algorithms have achieved some successes to overcome the difficulty of imposing the absorbing boundary condition in the time-discretization methods.
+This paper aims to provide more efficient, accurate and reliable simulation scheme that competes favorably with the existing simulation schemes in the literature.
+Regarding the sampling of the average variance conditional on the terminal volatility, Chen et al.
+(2012) sampled this quantity through the mean-and-variance matched lognormal (LN) variable.
 While the sampling procedures are fast, the time step cannot be large since the derived mean and variance are valid in small-time limit only.
-For more accurate sampling, researchers used various version of the transform of the average variance.
+For more accurate sampling, later works employed various versions of the transform of the average variance.
 Leitao et al.
 (2017a, b) derived the approximate Fourier transform of the unconditional average variance recursively, while Cai et al.
-(2017) adopted the analytically available Laplace transform of the reciprocal of the conditional average variance (Matsumoto and Yor, 2005).
-However, these approaches are cumbersome and time-consuming due to the tedious numerical inverse transform for constructing the cumulative distribution function (CDF) and iterative root-finding for inverse transform sampling.
-Cui et al.
+(2017) used the analytic formula of the Laplace transform of the reciprocal of the conditional average variance (Matsumoto and Yor, 2005).
+However, these approaches are cumbersome and time-consuming since the tedious numerical inverse transform procedures are required to construct the cumulative distribution function (CDF) and implement the iterative root-finding for the inverse transform sampling.
+As an alternative approach, Cui et al.
 (2021) approximated the volatility dynamics with the continuous-time Markov chain (CTMC) over volatility grids.
-Their approach is also not immune from heavy computation and complicated implementation.
-Kyriakou et al.
-(2023) used moment-matched Pearson family of distributions for efficient sampling.
-For such purpose, they rely on numerical evaluation of the four moments from the conditional Laplace transform.
-Sampling the terminal price conditional on the terminal variance and average variance is perhaps more challenging.
-Unlike the other stochastic volatility models that are based on the geometric Brownian motion (BM), the constant-elasticity-of-variance (CEV) feature of the SABR model does not admit an analytically tractable form of the conditional price distribution.
-As a result, the distribution has to be approximated by an analytic form first even before the corresponding sampling algorithm is considered.
-For the approximation of the conditional price distribution, almost all existing algorithms (Chen et al., 2012; Cai et al., 2017; Leitao et al., 2017a, b; Cui et al., 2018; Grzelak et al., 2019; Kyriakou et al., 2023) adopted Islah (2009)’s noncentral chi-squared (NCX2) distribution approximation unquestionably.
-However, the failure of the martingale property in this approximation has not been critically examined among these papers.
+Unfortunately, their approach does not immune from heavy computation and cumbersome implementation.
+In another approach, Kyriakou et al.
+(2023) used the moment-matched Pearson family of distributions for efficient sampling.
+For such purpose, they rely on numerical evaluation of the four order moments from the Laplace transform of conditional average variance.
+Sampling the terminal forward price conditional on the terminal variance and average variance is much more challenging.
+Unlike the other stochastic volatility models that are based on the geometric Brownian motion (BM), the constant-elasticity-of-variance (CEV) feature of the SABR model does not admit an analytically tractable form of the conditional terminal forward price distribution.
+As a result, the price distribution has to be approximated in some analytic form first before the corresponding sampling algorithm is considered.
+For the approximation of the conditional terminal forward price distribution, most existing algorithms (Chen et al., 2012; Cai et al., 2017; Leitao et al., 2017a, b; Cui et al., 2018; Grzelak et al., 2019; Kyriakou et al., 2023) invariably adopted Islah (2009)’s noncentral chi-squared (NCX2) distribution approximation.
+However, the failure of the martingale property of the conditional forward price in this approximation has not been critically examined among these earlier schemes.
 Leitao et al.
 (2017a, b) applied an ad-hoc correction on the simulation prices to enforce the martingale property.
-Therefore, the possibility of better approximation of the conditional price distribution that preserves the martingale should be explored.
-The lack of fast sampling algorithm for the approximated distribution is the last hurdle in constructing an efficient SABR simulation algorithm.
-The inverse transform sampling with numerical root-finding is accurate but tediously slow.
+Our paper proposes the CEV approximation of the conditional terminal forward price distribution, which is considered to be better since it preserves the martingale property and provides more efficient simulation procedures.
+Indeed, the lack of fast sampling algorithm for the approximated conditional forward price distribution is the last challenge in constructing an efficient SABR simulation algorithm.
+The inverse transform sampling with numerical root-finding procedure is accurate but tediously slow.
 The mean-and-variance-match quadratic Gaussian sampling has been adopted in Chen et al.
@@
 The contributions of this paper are three-fold.
-Firstly, we derive the first four conditional moments of the conditional average variance analytically.
-We then use the shifted lognormal (SLN) distribution matching the moments for fast sampling of the average variance.
-As we use the moments that are higher in order and exact for any arbitrary time step, the validity of our algorithm can be extended to larger time step.
-Secondly, we present an alternative approximation of the conditional price distribution based on the CEV process.
-Our approximation is superior to the widely adopted Islah (2009)’s approximation because it preserves the martingale property by construction.
-Lastly, we employ the algorithms of Makarov and Glew (2010) and Kang (2014) that are capable of sampling the CEV price process for an arbitrary time step.
-Their algorithms are fast since the CEV price can be expressed by a compound random variable composed of three elementary (i.e., gamma-Poison-gamma) random variables.
+Firstly, we derive the first four moments of the conditional average variance analytically.
+We then use the shifted lognormal (SLN) distribution obtained from matching the moments for fast sampling of the average variance.
+Since we match the moments that are in higher order and exact for any arbitrary time step, good accuracy of our SLN distribution approximation can be maintained at larger time step.
+Secondly, we present an alternative approximation of the conditional forward price distribution based on the CEV process.
+Our approximation shows better accuracy when compared with the widely adopted Islah’s approximation (Islah, 2009) since it preserves the martingale property of the conditional forward price by its construction.
+Lastly, we employ the algorithms of Makarov and Glew (2010) and Kang (2014) that are capable of sampling the CEV process exactly.
+Their algorithms are fast since the CEV distribution can be expressed by a compound random variable composed of three elementary (namely, gamma-Poisson-gamma) random variables.
 This paper is organized as follows.
 Section 2 presents the formulation of the SABR model and reviews some of its analytic properties.
-Section 3 describes our SLN approximation for sampling the average variance.
+Section 3 describes our SLN approximation for sampling the average variance and presents its performance in numerical accuracy.
 In Section 4, we present the CEV distribution approximation of the conditional forward price, and an exact sampling algorithm for the CEV distribution via the shifted Poisson-mixture Gamma distribution.
-In Sections 3 and 4, our simulation algorithm is compared to those in the existing literature.
-In particular, we argue that the CEV distribution is a better alternative to the widely adopted Islah (2009)’s approximation as it preserves the martingale condition.
-Section 5 presents our comprehensive numerical experiments that serve to assess accuracy, efficiency and reliability of different simulation schemes and analytic approximation methods, focusing on the more challenging choices of parameters, such as large time-to-maturity.
+In particular, we argue that our CEV distribution approximation of the conditional forward price is better than the widely adopted Islah’s approximation as it preserves the martingale condition.
+Section 5 presents our comprehensive numerical experiments that serve to assess accuracy, efficiency and reliability of different simulation schemes and analytic approximation methods, displaying numerical comparisons under the more challenging choices of parameters, such as large time-to-maturity.
+Comparison of accuracy performance of our CEV approximation of the forward price with that of the Islah’s approximation is carefully examined.
 Conclusive remarks are summarized in Section 6.
-2 Formulation of the SABR model and its special cases
-In this section, we introduce the SABR model and review some of its analytic properties.
+2 Formulation of the SABR model and its two-step simulation
+In this section, we introduce the SABR model and review some of its analytic properties under various degenerate cases.
 The two steps of the SABR simulation are outlined.
@@
 2.1 SABR model
-The governing stochastic differential equations (SDE) for the SABR volatility model (Hagan et al., 2002) are given by
-$\frac{\mathrm{d}F_{t}}{F_{t}^{\beta}}=\sigma_{t}\,\mathrm{d}W_{t}\quad\text{and}\quad\frac{\mathrm{d}\sigma_{t}}{\sigma_{t}}=\nu\,\mathrm{d}Z_{t},\quad(0\leq\beta\leq 1),$ (1)
-where $F_{t}$ and $\sigma_{t}$ are the stochastic processes for the forward price and volatility, respectively, $\nu$ is the vol-of-vol, $\beta$ is the elasticity of variance parameter, and $W_{t}$ and $Z_{t}$ are the standard BMs that are correlated with correlation coefficient $\rho$.
-We also define two variables for later use:
+The governing stochastic differential equations (SDEs) for the SABR volatility model (Hagan et al., 2002) are given by
+$\frac{\mathrm{d}F_{t}}{F_{t}^{\beta}}=\sigma_{t}\,\mathrm{d}W_{t}\quad\text{and}\quad\frac{\mathrm{d}\sigma_{t}}{\sigma_{t}}=\nu\,\mathrm{d}Z_{t},\quad 0\leq\beta\leq 1,$ (1)
+where $F_{t}$ and $\sigma_{t}$ are the stochastic processes for the forward price and volatility, respectively, $\nu$ is the vol-of-vol, $\beta$ is the elasticity of variance parameter, and $W_{t}$ and $Z_{t}$ are the standard Brownian motions (BMs) that are correlated with correlation coefficient $\rho$.
+We also define two parameters for notational simplicity:
 $\rho_{\ast}:=\sqrt{1-\rho^{2}}\quad\text{and}\quad\beta_{\ast}:=1-\beta.$
-2.2 Average variance, conditional expectation, and two-step simulation procedures
-Regarding the simulation step, without loss of generality, we sample the volatility and forward price at time $t+h$ (i.e., $\sigma_{t+h}$ and $F_{t+h}$) for some time step $h$ given time $t$.
-Accordingly, the filtration up to time $t$ (e.g., $\sigma_{t}$ and $F_{t}$) is implicitly assumed.
-Since the volatility process in Eq.
-(1) follows a geometric BM, $\sigma_{t}=\sigma_{0}\exp\left(\nu Z_{t}-\nu^{2}t/2\right)$, simulating $\sigma_{t}$ from time $t$ to $t+h$ can be easily done by observing
+2.2 Average variance and two-step simulation procedures
+Regarding the simulation over one time step with size $h$, we sample the volatility and forward price at time $t+h$ ($\sigma_{t+h}$ and $F_{t+h}$) for discrete time step $h$ at given time $t$.
+Accordingly, the filtration up to time $t$ (say, $\sigma_{t}$ and $F_{t}$) is implicitly assumed.
+Since the volatility process $\sigma_{t}$ in Eq.
+(1) follows a geometric BM, we have $\sigma_{t}=\sigma_{0}\exp\left(\nu Z_{t}-\nu^{2}t/2\right).$ Simulating $\sigma_{t}$ from time $t$ to $t+h$ can be easily performed via
 $\sigma_{t+h}=\sigma_{t}\exp\left(\nu Z_{h}-\frac{\nu^{2}h}{2}\right)=\sigma_{t}\exp\left(\hat{\nu}\hat{Z}\right)\quad\quad\text{where}\quad\hat{Z}\sim\mathcal{N}\left(-\frac{\hat{\nu}}{2},1\right).$ (2)
-Here, $\hat{\nu}:=\nu\sqrt{h}$ is the standard deviation of $\ln\sigma_{t}$ over the time step $h$. This quantity is used as a measure of small-time limit.
-Like the simulation schemes of other SV models, integrated variance conditional on initial and terminal volatility values plays an important role in the SABR model simulation.
-Given the initial and terminal volatility (i.e., $\sigma_{t}$ and $\sigma_{t+h}$), we define the conditional average variance between $t$ and $t+h$ as
+Here, $\hat{\nu}:=\nu\sqrt{h}$ is the standard deviation of $\ln\sigma_{t}$ over the time step $h$. The quantity $\hat{\nu}$ is used as a measure of small-time limit.
+Like the simulation schemes in other stochastic volatility models, the average variance conditional on initial and terminal volatility values plays an important role in the SABR model simulation.
+Given the initial and terminal volatility (namely, $\sigma_{t}$ and $\sigma_{t+h}$), we define the dimensionless conditional average variance between $t$ and $t+h$ as
 $I_{t}^{h}(\sigma_{t+h}):=\frac{1}{\sigma_{t}^{2}h}\int_{t}^{t+h}\sigma_{s}^{2}\,\mathrm{d}s\;\Big|_{\sigma_{t+h}}.$ (3)
-Here, we normalize the integral by $\sigma_{t}^{2}h$ in order that $I_{t}^{h}$ becomes a dimensionless quantity in the order of one.
+Here, we normalize the integrated variance by $\sigma_{t}^{2}h$ in order that $I_{t}^{h}$ becomes a dimensionless quantity of order one.
 Specifically, $I_{t}^{h}$ converges to one in the limit of $\nu\downarrow 0$.
@@
 $I_{t}^{h}(\hat{Z})\;\sim\;\int_{0}^{1}e^{2\hat{\nu}Z_{s}}\,\mathrm{d}s\;\Big|_{Z_{1}=\hat{Z}}\;$.
-For notational simplicity, we omit the dependence on $\sigma_{t+h}$ or $\hat{Z}$ and simply write $I_{t}^{h}$.
-Almost all SABR simulation algorithms are required to perform the following two steps, given the filtration up to $t$, to sufficient accuracy:
+For notational simplicity, we simply write $I_{t}^{h}$ by omitting the dependence on $\sigma_{t+h}$ or $\hat{Z}$.
+Almost all SABR simulation algorithms strive to perform the following two steps of simulation to achieve efficiency and high accuracy, given the filtration up to $t$:
 Step 1
@@
 : simulation of the forward price $F_{t+h}$ conditional on $\sigma_{t+h}$ and $I_{t}^{h}$.
-As both simulation steps pose challenges, the previous studies on the topic have proposed numerous algorithms for the two steps.
-This paper also aims to innovate the two steps, details are presented in Sections 3 and 4, respectively.
-Regarding Step 2, let us also define the conditional expectation of $F_{t+h}$:
+This paper aims to innovate these two steps of simulation over existing algorithms, details are presented in Sections 3 and 4, respectively.
+Regarding Step 2, we define the conditional expectation of $F_{t+h}$:
 $\bar{F}_{t}^{h}(\sigma_{t+h},I_{t}^{h}):=E\left(F_{t+h}\,|\,\sigma_{t+h},I_{t}^{h}\right),$ (4)
-which is essential for the discussion of the simulation algorithms of Step 2. We also omit the dependence on $\sigma_{t+h}$ and $I_{t}^{h}$, and simply write $\bar{F}_{t}^{h}$.
+which is essential for the discussion of the simulation algorithms of Step 2. For notational simplicity, we omit the dependence on $\sigma_{t+h}$ and $I_{t}^{h}$, and simply write $\bar{F}_{t}^{h}$.
 2.3 Special cases
-Before introducing our simulation algorithm, we review several special cases of the SABR model.
-In these cases, Step 2 is relatively easier while Step 1 remains challenging.
-Reviewing these cases also helps building intuitions for our new algorithm.
+Before introducing our simulation algorithm, we review several special cases of the SABR model, which helps building intuition for our new algorithm.
 Normal ($\beta=0$) SABR model: When $\beta=0$, conditional on $\sigma_{T}$ and $I_{0}^{T}$, integrating $F_{t}$ from $t=0$ to $T$ reveals that the terminal forward price $F_{T}$ follows the normal distribution with mean $\bar{F}_{0}^{T}$ and variance $\rho_{\ast}^{2}\sigma_{0}^{2}TI_{0}^{T}$:
-$F_{T}\,|\,\sigma_{T},I_{0}^{T}\;\sim\;\mathcal{N}\left(\bar{F}_{0}^{T},\rho_{\ast}^{2}\sigma_{0}^{2}TI_{0}^{T}\right)$
+$F_{T}\,|\,\sigma_{T},I_{0}^{T}\;\sim\;\mathcal{N}\left(\bar{F}_{0}^{T},\rho_{\ast}^{2}\sigma_{0}^{2}TI_{0}^{T}\right),$
 where the conditional expectation $\bar{F}_{0}^{T}$ is given by
@@
 Choi et al.
-(2019) have also found an exact closed-form expression for sampling $F_{T}$ without separating Steps 1 and 2.
-Moreover, note that $F_{T}$, as well as $\bar{F}_{0}^{T}$ can be negative in this special case as the origin is not a boundary.
-The normal SABR model is different from the $\beta\downarrow 0$ limit of the SABR model, where the origin is an absorbing boundary.
-In the view of the above, the SABR model under $\beta=0$ is not a concern of our study.
-Lognormal ($\beta=1$) SABR model: When $\beta=1$, conditional on $\sigma_{T}$ and $I_{0}^{T}$, integrating $\ln F_{t}$ from $t=0$ to $T$ yields that the terminal forward price $F_{T}$ follows the LN distribution with mean $\bar{F}_{0}^{T}$ and log variance $\rho_{\ast}^{2}\sigma_{0}^{2}TI_{0}^{T}$:
+(2019) managed to find an exact closed-form expression for sampling $F_{T}$ without separating Steps 1 and 2.
+Moreover, note that $F_{T}$, as well as $\bar{F}_{0}^{T}$ can be negative under normal distribution as the origin is not a boundary.
+Accordingly, the normal SABR model is different from the $\beta\downarrow 0$ limit of the SABR model.
+In the later case, the origin is an absorbing boundary.
+In view of these subtleties, we discard further discussion of the SABR model under $\beta=0$.
+Lognormal ($\beta=1$) SABR model: When $\beta=1$, conditional on $\sigma_{T}$ and $I_{0}^{T}$, integrating $\ln F_{t}$ from $t=0$ to $T$ reveals that the terminal forward price $F_{T}$ follows the LN distribution with mean $\bar{F}_{0}^{T}$ and log variance $\rho_{\ast}^{2}\sigma_{0}^{2}TI_{0}^{T}$:
 $F_{T}\,|\,\sigma_{T},I_{0}^{T}\;\sim\;\bar{F}_{0}^{T}\,\exp\left(\,\rho_{\ast}\sigma_{0}\sqrt{TI_{0}^{T}}\,X-\frac{\rho_{\ast}^{2}\sigma_{0}^{2}TI_{0}^{T}}{2}\right),$ (6)
-where $X$ is a standard normal variate independent of $\sigma_{T}$ and $I_{0}^{T}$. The conditional expectation $\bar{F}_{0}^{T}$ is given by
+where $X$ is a standard normal variate independent of $\sigma_{T}$ and $I_{0}^{T}$. The conditional expectation $\bar{F}_{0}^{T}$ is explicitly given by
 $\bar{F}_{0}^{T}=F_{0}\exp\left(\frac{\rho}{\nu}(\sigma_{T}-\sigma_{0})-\frac{\rho^{2}\sigma_{0}^{2}TI_{0}^{T}}{2}\right).$ (7)
 Therefore, Step 2 becomes trivial as well when $\beta=1$.
-This special case provides an important observation, which derives an intuition for our CEV approximation when $10<\beta<$ 1 (to be discussed Section 4).
+This special case of $\beta=1$ provides the important intuition for our CEV approximation of the forward price under $0<\beta<1$ (to be discussed in Section 4).
 Remark 1.
-The joint distribution of $\sigma_{T}$ and $I_{0}^{T}$ satisfies
+From the law of total expectation, it follows that
+$E(\bar{F}_{0}^{T})=E\left(E\left(F_{T}|\sigma_{T},I_{0}^{T}\right)\right)=E(F_{T})=F_{0},$
+and the joint distribution of $\sigma_{T}$ and $I_{0}^{T}$ satisfies
 $E\left[\exp\left(\frac{\rho}{\nu}(\sigma_{T}-\sigma_{0})-\frac{\rho^{2}\sigma_{0}^{2}TI_{0}^{T}}{2}\right)\right]=1,$
 for any initial volatility $\sigma_{0}>0$.
-Proof.
-It follows from the law of total expectation:
-$E(\bar{F}_{0}^{T})=E\left(E\left(F_{T}|\sigma_{T},I_{0}^{T}\right)\right)=E(F_{T})=F_{0}.\qed$
-Zero vol-of-vol ($\nu=0$) SABR model: When $10<\beta<$ 1, the SABR model is visualized as the stochastic volatility version of the CEV model:
-$\frac{\mathrm{d}F_{t}}{F_{t}^{\beta}}=\sigma_{0}\,\mathrm{d}W_{t},$ (8)
-with the absorbing boundary condition at the origin.
+Zero vol-of-vol ($\nu=0$) SABR model: For $0<\beta<1$, the SABR model is now reduced to the CEV model:
+$\frac{\mathrm{d}F_{t}}{F_{t}^{\beta}}=\sigma_{0}\,\mathrm{d}W_{t},$ (8a)
+with the absorbing boundary condition at the origin. Here, $\sigma_{0}$ is constant. Indeed, the SABR model can be visualized as the stochastic volatility version of the CEV model.
 Definition 1.
-We define the distribution of $F_{T}$ resulting from the CEV dynamics in Eq. (8) as the CEV distribution,
-$F_{T}\sim\mathcal{CEV}_{\beta}(F_{0},\sigma_{0}^{2}T),$
-parameterized by the elasticity-of-variance parameter $\beta$, mean $F_{0}$ [^1] If $F_{T}\sim\mathcal{CEV}_{\beta}(F_{0},\sigma_{0}^{2}T)$, then $E(F_{T})=F_{0}$ regardless of the values of $\beta$ and $\sigma_{0}^{2}T$., and variance $\sigma_{0}^{2}T$.
-The probability distribution function (PDF) and CDF of the CEV distribution will be shown later in Proposition 5.
-Therefore, in the zero vol-of-vol ($\nu\downarrow 0$) limit, $F_{T}$ from the SABR model (unconditionally) follows $\mathcal{CEV}_{\beta}(F_{0},\sigma_{0}^{2}T)$.
+The CEV distribution is defined as the distribution of $F_{T}$ resulting from the CEV dynamics as depicted in Eq. (8a), which is denoted by
+$F_{T}\sim\mathcal{CEV}_{\beta}(F_{0},\sigma_{0}^{2}T).$ (8b)
+It is parameterized by the elasticity-of-variance parameter $\beta$, mean $F_{0}$, and variance $\sigma_{0}^{2}T$.
+The cumulative distribution function (CDF) and probability density function (PDF) of the CEV distribution are presented respectively in Eqs.
+(22a) and  (22b) later.
+If $F_{T}\sim\mathcal{CEV}_{\beta}(F_{0},\sigma_{0}^{2}T)$, then $E(F_{T})=F_{0}$ regardless of the values of $\beta$ and $\sigma_{0}^{2}T$.
+In the zero vol-of-vol ($\nu\downarrow 0$) limit, $F_{T}$ from the SABR model follows $\mathcal{CEV}_{\beta}(F_{0},\sigma_{0}^{2}T)$ unconditionally.
 Step 2 is reduced to sampling $F_{T}$ from the CEV distribution.
-However, this is not a trivial task as we discuss later in Section 4.
 Uncorrelated ($\rho=0$) SABR model: The terminal forward price $F_{T}$ also exhibits a CEV distribution under non-zero vol-of-vol ($\nu>0$) and zero correlation ($\rho=0$) case, albeit conditionally.
@@
 Basically, $TI_{0}^{T}$ plays the role of stochastic time clock of the CEV model.
-As in the $\nu=0$ case, Step 2 is reduced to sampling $F_{T}$ from the CEV distribution in this case.
-Unlike the $\beta=1$ case, $F_{T}$ is conditioned by $I_{0}^{T}$ only since conditional expectation is always $F_{0}$ without depending on $\sigma_{T}$.
+Like the $\nu=0$ case, Step 2 is reduced to sampling $F_{T}$ from the CEV distribution under $\rho=0$.
+Unlike the $\beta=1$ case, $F_{T}$ is only conditioned by $I_{0}^{T}$ since conditional expectation is always $F_{0}$ without depending on $\sigma_{T}$.
 Thus, $I_{0}^{T}$ should be understood as the unconditional average variance in this context.
-Therefore, the European option price can be simplified to one dimensional integral of the CEV option price over the (unconditional) distribution of $I_{0}^{T}$.
-Choi and Wu (2021b) have approximated the integral using the Gaussian quadrature together with the LN distribution fitted to the (unconditional) mean and variance of $I_{0}^{T}$.
-Completing the discussion on the four special cases, we are left with the most general case with $10<\beta<$ 1 and $\rho\neq 0$, under which the conditional distribution of $F_{T}$ is not analytically tractable.
-One has to resort to an approximation.
-The approximation should be reduced to the special cases discussed above under various continuous limits.
+Therefore, the European option price can be expressed as one dimensional integral of the CEV option price over the (unconditional) distribution of $I_{0}^{T}$.
+As a remark, Choi and Wu (2021b) approximated the integral using the Gaussian quadrature together with the LN distribution fitted to the (unconditional) mean and variance of $I_{0}^{T}$.
+Besides the above four special cases, one has to face with the general case with $0<\beta<1$ and $\rho\neq 0$, under which the conditional distribution of $F_{T}$ is not analytically tractable.
+The success of a simulation algorithm for the SABR model amounts to an ingenious choice chosen for the approximation of the conditional distribution of $F_{T}$.
 3 First simulation step: Sampling average variance $I_{t}^{h}$
 This section presents our simulation method for sampling $I_{t}^{h}$. We also compare with other most popular simulation methods, highlighting the computational advantages of our simulation scheme.
-3.1 Sampling $I_{t}^{h}$ from the moment-matched SLN approximation
+3.1 Sampling $I_{t}^{h}$ from the moment-matched shifted lognormal (SLN) approximation
 We sample $I_{t}^{h}$ from the SLN distribution matched to the first three moments of $I_{t}^{h}$.
-Definition 2 (SLN variable).
+Definition 2.
 The SLN random variable, $Y\sim\mathcal{SLN}(\mu,\sigma^{2},\lambda)$, is given by
@@
 where $X$ is a standard normal variate.
-Here, $\mu$ is the mean of $Y$, $\lambda$ and $\sigma^{2}$ are the weight and variance of the LN component, respectively.
+Here, $\mu$ is the mean of $Y$, $\lambda$ and $\sigma^{2}$ are the shifted parameter and variance of the LN component, respectively.
 When $\lambda=1$, $Y$ is reduced to the LN variable with mean $\mu$ and log variance $\sigma^{2}$.
-The mean, coefficient of variation, skewness, and ex-kurtosis of $Y$ are given by
-$E(Y)=\mu,\quad\text{Cv}(Y)=\lambda\sqrt{w},\quad\text{Sk}(Y)=\sqrt{w}(w+3),\quad\text{and}\quad\text{Ek}(Y)=w(w^{3}+6w^{2}+15w+16),$
-where $w=\exp\left(\sigma^{2}\right)-1\;(w>0)$.
+The coefficient of variation, skewness, and ex-kurtosis of $Y$ are given by
+$v(Y)=\lambda\sqrt{w},\quad\text{s}(Y)=\sqrt{w}(w+3),\quad\text{and}\quad\text{k}(Y)=w(w^{3}+6w^{2}+15w+16),$
+respectively, where $w=e^{\sigma^{2}}-1\;(w>0)$.
 Remark 2.
-While the skewness and ex-kurtosis of distribution $Y$ follow the conventional definitions, the coefficient of variation is defined as the ratio of standard deviation to mean:
-$\mathrm{Cv}(Y)=\frac{\sqrt{\mathrm{Var}(Y)}}{E(Y)}.$
-The coefficient of variation is an important dimensionless quantity that characterizes non-negative distributions such as the LN or SLN distributions.
+While the skewness and ex-kurtosis of distribution $Y$ follow the conventional definitions, the coefficient of variation is defined as the ratio of the standard deviation to mean:
+$v(Y)=\frac{\sqrt{\mathrm{Var}(Y)}}{E(Y)}.$
+The coefficient of variation is an important dimensionless quantity that characterizes non-negative distributions such as the LN and SLN distributions.
 Proposition 1 (SLN distribution matched to the first three moments).
-Given mean $\mu>0$, coefficient of variance $v>0$, and skewness $s>0$, the parameters of the SLN distribution matching $\mu$, $v$, and $s$ are analytically given by
-$\sigma=\sqrt{\ln(1+w)}\quad\text{for}\quad w=4\sinh^{2}\left(\frac{1}{6}\operatorname{arcosh}\left(1+\frac{s^{2}}{2}\right)\right)$
-$\text{and}\quad\lambda=\frac{v}{2\sinh\left(\frac{1}{6}\operatorname{arcosh}\left(1+\frac{s^{2}}{2}\right)\right)}.$
+Given mean $\mu>0$, coefficient of variance $v>0$, and skewness $s>0$, the parameters of the SLN distribution matching $\mu$, $v$, and $s$ admit the following analytic formulas:
+$\sigma=\sqrt{\ln(1+w)}\quad\text{for}\quad w=4\sinh^{2}\left(\frac{1}{6}\operatorname{arcosh}\left(1+\frac{s^{2}}{2}\right)\right),\quad\text{and}\quad\lambda=\frac{v}{2\sinh\left(\frac{1}{6}\operatorname{arcosh}\left(1+\frac{s^{2}}{2}\right)\right)}.$
 Proof.
 We first fit $\sigma$ to the given skewness $s$:
-$s^{2}=w(w+3)^{2}\quad\text{for}\quad w=\exp\left(\sigma^{2}\right)-1.$
+$s^{2}=w(w+3)^{2}\quad\text{for}\quad w=e^{\sigma^{2}}-1.$
 Substituting $x=w+2$ transforms the equation to a special case of the cubic equation:
@@
 $x=2\cosh\left(\frac{1}{3}\operatorname{arcosh}\left(1+\frac{s^{2}}{2}\right)\right).$
-The expression for $w$ follows from $\cosh\theta-1=2\sinh^{2}(\theta/2)$. Next, we fit $\lambda$ to the given coefficient of variance $v$:
-$\lambda=\frac{v}{\sqrt{w}}=\frac{v}{2\sinh\left(\frac{1}{6}\operatorname{arcosh}\left(1+\frac{s^{2}}{2}\right)\right)}.\qed$
+The expression for $w$ follows from $\cosh\theta-1=2\sinh^{2}\frac{\theta}{2}$. Next, we fit $\lambda$ to the given coefficient of variance $v$:
+$\lambda=\frac{v}{\sqrt{w}}=\frac{v}{2\sinh\left(\frac{1}{6}\operatorname{arcosh}\left(1+\frac{s^{2}}{2}\right)\right)}.$
 Remark 3.
@@
 Chen et al. (2012) has sampled $I_{t}^{h}$ through the LN random variable fitted to the mean and variance of $I_{t}^{h}$ in the small-time limit.
-Next, we derive the first four conditional moments of $I_{t}^{h}$, whose analytic formulas are presented in the next proposition.
+Next, we derive the first four conditional moments of $I_{t}^{h}$, whose analytic formulas are presented in Proposition 2.
 Proposition 2 (Conditional moments of average variance).
-Given $\sigma_{t}$ and $\sigma_{t+h}$, the first four raw moments of $I_{t}^{h}$ (i.e., $\mu=E(I_{t}^{h})$ and $\mu^{\prime}_{k}=E\left((I_{t}^{h})^{k}\right)$) are obtained as below.
+Given $\sigma_{t}$ and $\sigma_{t+h}$, the first four raw moments of $I_{t}^{h}$ (i.e., $\mu=E(I_{t}^{h})$ and $\mu^{\prime}_{k}=E\left((I_{t}^{h})^{k}\right)$) are given by
 $\mu$ $=\left(\frac{\sigma_{t+h}}{\sigma_{t}}\right)m_{1},$
-$\mu^{\prime}_{2}$ $=\left(\frac{\sigma_{t+h}}{\sigma_{t}}\right)^{2}\frac{1}{\hat{\nu}^{2}}\left[m_{2}-cm_{1}\right],$
+$\mu^{\prime}_{2}$ $=\left(\frac{\sigma_{t+h}}{\sigma_{t}}\right)^{2}\frac{1}{\hat{\nu}^{2}}\left(m_{2}-cm_{1}\right),$
 $\mu^{\prime}_{3}$ $=\left(\frac{\sigma_{t+h}}{\sigma_{t}}\right)^{3}\frac{1}{8\hat{\nu}^{4}}\left[3m_{3}-8cm_{2}+\left(4c^{2}+1\right)m_{1}\right],$
 $\mu^{\prime}_{4}$ $=\left(\frac{\sigma_{t+h}}{\sigma_{t}}\right)^{4}\frac{1}{24\,\hat{\nu}^{6}}\left[2m_{4}-9cm_{3}+\left(12c^{2}+2\right)m_{2}-c\left(4c^{2}+3\right)m_{1}\right],$
-where the conditioning variable, $\sigma_{t+h}$ and $\hat{Z}$, are exchangeable by $\hat{Z}=\frac{1}{\hat{\nu}}\ln(\sigma_{t+h}/\sigma_{t})$, and the coefficients, $c$ and $m_{k}\,(k=1,2,\dots,4)$, are the functions of them:
-$c(\hat{Z}):=\cosh(\hat{\nu}\hat{Z})=\frac{1}{2}\left(\frac{\sigma_{t+h}}{\sigma_{t}}+\frac{\sigma_{t}}{\sigma_{t+h}}\right)\quad\text{and}\quad m_{k}(\hat{Z})=\frac{N(\hat{Z}+k\hat{\nu})-N(\hat{Z}-k\hat{\nu})}{2k\hat{\nu}\;n\left(\sqrt{\hat{Z}^{2}+(k\hat{\nu})^{2}}\right)},$
+where the conditioning variable, $\sigma_{t+h}$ and $\hat{Z}$, are exchangeable by $\hat{Z}=\frac{1}{\hat{\nu}}\ln(\sigma_{t+h}/\sigma_{t})$, and the coefficients, $c$ and $m_{k}\,(k=1,2,3,4)$, are functions of $\sigma_{t}$, $\sigma_{t+h}$ and $\hat{Z}$:
+$c(\hat{Z}):=\cosh\hat{\nu}\hat{Z}=\frac{1}{2}\left(\frac{\sigma_{t+h}}{\sigma_{t}}+\frac{\sigma_{t}}{\sigma_{t+h}}\right)\quad\text{and}\quad m_{k}(\hat{Z})=\frac{N(\hat{Z}+k\hat{\nu})-N(\hat{Z}-k\hat{\nu})}{2k\hat{\nu}\;n\left(\sqrt{\hat{Z}^{2}+(k\hat{\nu})^{2}}\right)},$
 and $n(z)$ and $N(z)$ are the PDF and CDF of the standard normal distribution, respectively.
 Proof.
-The analytic derivation is based on the conditional moments of the exponential functional of BM (Matsumoto and Yor, 2005, (5.4)). See A for the derivation. ∎
+The analytic derivation is based on the conditional moments of the exponential functional of Brownian motion (Matsumoto and Yor, 2005, (5.4)). See A for the derivation. ∎
 Remark 4.
@@
 From the first four raw moments, the coefficient of variation $v$, skewness $s$, and ex-kurtosis $\kappa$ of $I_{t}^{h}$ are obtained as
-$v=\frac{\sqrt{\mu_{2}}}{\mu},\quad s=\frac{\mu^{\prime}_{3}-3\mu\mu^{\prime}_{2}+2\mu^{3}}{\mu_{2}\sqrt{\mu_{2}}},\quad\text{and}\quad\kappa=\frac{\mu^{\prime}_{4}-4\mu\mu^{\prime}_{3}+6\mu^{2}\mu^{\prime}_{2}-3\mu^{4}}{\mu_{2}^{2}}-3,$
+$v(I_{t}^{h})=\frac{\sqrt{\mu_{2}}}{\mu},\quad s(I_{t}^{h})=\frac{\mu^{\prime}_{3}-3\mu\mu^{\prime}_{2}+2\mu^{3}}{\mu_{2}\sqrt{\mu_{2}}},\quad\text{and}\quad\kappa(I_{t}^{h})=\frac{\mu^{\prime}_{4}-4\mu\mu^{\prime}_{3}+6\mu^{2}\mu^{\prime}_{2}-3\mu^{4}}{\mu_{2}^{2}}-3,$
 where $\mu_{2}=\mu^{\prime}_{2}-\mu^{2}$ is the variance.
 From Propositions 1 and 2, it is possible to sample $I_{t}^{h}$ from the SLN variable whose parameters are fitted to the first three moments of $I_{t}^{h}$ exactly.
-However, we take a simpler computational approach that uses a fixed $\lambda$ obtained from the small-time limit.
-It is shown to be effective as well.
-Proposition 3 (SLN in the small-time limit).
+However, we take a much simpler computational approach that uses a fixed $\lambda$ obtained from the small-time limit.
+It is shown to be effective as well (see Algorithm 1).
+In A, the coefficient of variation and skewness of $I_{t}^{h}$ are expanded around $\hat{\nu}=0$ as
+$v=\frac{\sqrt{\mu_{2}}}{\mu}=\frac{\hat{\nu}}{\sqrt{3}}+O(\hat{\nu}^{3})\quad\text{and}\quad s=\frac{\mu_{3}}{\mu_{2}\sqrt{\mu_{2}}}=\frac{6\sqrt{3}}{5}\hat{\nu}+O(\hat{\nu}^{3}),\quad\text{where }\hat{\nu}=\nu\sqrt{h}$
+Using Taylor’s expansion, we obtain
+$2\sinh\left(\frac{1}{6}\operatorname{arcosh}\left(1+\frac{s^{2}}{2}\right)\right)=\frac{s}{3}-\frac{s^{3}}{81}+O(s^{5}),$
+so that the shift parameter $\lambda$ converges to
+$\lambda=\frac{v}{2\sinh\left(\frac{1}{6}\operatorname{arcosh}\left(1+\frac{s^{2}}{2}\right)\right)}\to\frac{5}{6}\quad\text{as}\quad\hat{\nu}\downarrow 0.$
+Once $\lambda$ is determined, we fit $\sigma$ to match the coefficient of variation $v$:
+$\sigma=\sqrt{\ln\left(1+\frac{v^{2}}{\lambda^{2}}\right)}=\sqrt{\ln\left(1+\frac{36}{25}v^{2}\right)}.$
+As we demonstrate in the numerical results shown in Figure 1, this approximation is highly reliable not only for small $\hat{\nu}$ but even for $O(\hat{\nu})=1$.
+We summarize the above results into Algorithm 1 for approximating conditional $I_{t}^{h}$ as follows:
 In the small-time limit ($\hat{\nu}\downarrow 0$), $I_{t}^{h}$ is approximately sampled via an SLN variable with $\lambda=5/6$:
@@
 where $\mu$ and $v$ are the mean and coefficient of variance of $I_{t}^{h}$ given by Proposition 2 and Remark 5.
-Proof.
-In A, the coefficient of variation and skewness of $I_{t}^{h}$ are expanded around $\hat{\nu}=0$ as
-$v=\frac{\sqrt{\mu_{2}}}{\mu}=\frac{\hat{\nu}}{\sqrt{3}}+O(\hat{\nu}^{3})\quad\text{and}\quad s=\frac{\mu_{3}}{\mu_{2}\sqrt{\mu_{2}}}=\frac{6\sqrt{3}}{5}\hat{\nu}+O(\hat{\nu}^{3}).$
-Using Taylor’s expansion, we have
-$2\sinh\left(\frac{1}{6}\operatorname{arcosh}\left(1+\frac{s^{2}}{2}\right)\right)=\frac{s}{3}-\frac{s^{3}}{81}+O(s^{5}),$
-so that the shift parameter $\lambda$ converges to
-$\lambda=\frac{v}{2\sinh\left(\frac{1}{6}\operatorname{arcosh}\left(1+\frac{s^{2}}{2}\right)\right)}\to\frac{5}{6}\quad\text{as}\quad\hat{\nu}\downarrow 0.$
-Once $\lambda$ is determined, we fit $\sigma$ to match the coefficient of variation:
-$\sigma=\sqrt{\ln\left(1+\frac{v^{2}}{\lambda^{2}}\right)}=\sqrt{\ln\left(1+\frac{36}{25}v^{2}\right)}.$
-As we demonstrate in the numerical results shown in Figure 1, this approximation is highly reliable not only for small $\hat{\nu}$ but also for $O(\hat{\nu})=1$. ∎
-Figure 1: Variance (left), skewness (middle), and Ex-kurtosis (right) of $I_{t}^{h}$ as functions of $\hat{Z}$ for $\hat{\nu}=0.4$.
-We compare our SLN approximation with $\lambda=5/6$ (yellow), the SLN approximation fitted up to skewness (green), and the LN approximation fitted up to variance (blue) to the true values (dashed).
+Algorithm 1 Approximation of the conditional $I_{t}^{h}$
+Figure 1: Comparison of accuracy of calculating the skewness (left) and ex-kurtosis (right) of $I_{t}^{h}$ as functions of $\hat{Z}$ for $\hat{\nu}=0.4$.
+We compare our SLN approximation with $\lambda=5/6$ (orange), the SLN approximation fitted up to skewness (green), and the LN approximation fitted up to variance (blue) to the true values (dashed).
 3.2 Comparison to other simulation methods in literature
-Our sampling method for $I_{t}^{h}$ is much faster and easier to implement compared to other methods in the literature, such as Chen et al.
+Our sampling method for the approximation of conditional $I_{t}^{h}$ is much faster and easier to implement compared to other methods in the literature, such as Chen et al.
 (2012), Cai et al.
@@
 Our method is considered as an extension of Chen et al.
-(2012) as they use an LN approximation matched to mean and variance of $I_{t}^{h}$.
-However, they derive the mean and variance only in the small-time limit, so the time step should be small accordingly.
+(2012).
+Their LN approximation is matched up to mean and variance of $I_{t}^{h}$, while our approximation is matched up to the first three moments.
+Furthermore, they derive the mean and variance only in the small-time limit, so the time step has to be chosen to be small accordingly.
 We enhance the method by deriving the moments analytically and matching to higher moments with the SLN approximation.
-Figure 1 displays the accuracy of variance, skewness, and ex-kurtosis of $I_{t}^{h}$ obtained from our SLN approximation.
-The left figure shows that Chen et al.
-(2012)’s small-time approximation of the variance is not accurate across the dependency $Z$ while our approximation is exact.
-The middle figure shows that the SLN approximation gives the correct skewness value as the parameters are fitted up to skewness.
-Moreover, the skewness under the simple SLN scheme with $\lambda=5/6$ also agrees well for $\nu=0.2$.
-Even though the two SLN schemes do not use the ex-kurtosis in the parameter calibration, the ex-kurtosis is close to the true value, justifying our choice of the SLN random variable.
-Cai et al.
-(2017) computed the analytic form of the Laplace transform of the CDF of the reciprocal of the average variance 1/$I_{t}^{h}$, then used the Euler inversion formula to obtain the CDF of 1/$I_{t}^{h}$.
-After obtaining the CDF, they used the inverse transform method, which requires some tedious root-finding procedures such as Newton’s method or bisection method to sample $I_{t}^{h}$.
+Figure 1 displays accuracy of skewness and ex-kurtosis of $I_{t}^{h}$ obtained from our SLN approximation.
+The left figure shows that the SLN approximation (green curve) gives the correct skewness value across the dependency of $\hat{Z}$ since the parameters are fitted up to skewness.
+Moreover, the skewness under the simple SLN scheme with $\lambda=5/6$ (orange curve) also agrees well.
+Even though the two SLN schemes do not use ex-kurtosis in the parameter calibration, the SLN approximation of the ex-kurtosis of $I_{t}^{h}$ is close to the true value.
+This justifies that our chosen SLN approximation (either exact $\lambda$ or $\lambda=\frac{5}{6}$) is a good choice of approximation of $I_{t}^{h}$ (see right figure).
+As evidenced from the plots in Figure 1, the SLN approximation (either exact $\lambda$ or $\lambda=\frac{5}{6}$) outperforms the LN approximation.
+[…]
 Leitao et al.
 (2017a, b) used recursion schemes to find the characteristic function of the logarithm of $I_{t}^{h}$ and then employed the inverse Fourier technique COS method (Fang and Oosterlee, 2008) to recover its PDF.
-They then computed the Pearson correlation coefficient of $\ln I_{t}^{h}$ and $\ln\sigma_{t+h}$, then the bivariate copula distribution based on the CDF of $\ln\sigma_{T}$ and approximate CDF of $\ln I_{t}^{h}$.
+Next they computed the Pearson correlation coefficient of $\ln I_{t}^{h}$ and $\ln\sigma_{t+h}$, then the bivariate copula distribution based on the CDF of $\ln\sigma_{T}$ and approximate CDF of $\ln I_{t}^{h}$.
 Finally, they adopted the direct inverse transform method based on linear interpolation to sample $I_{t}^{h}$.
 Cui et al.
-(2021) used the continuous-time Markov chain (CTMC) to approximate the space of the volatility process with finite number of states.
+(2021) used the continuous-time Markov chain (CTMC) in the space of the volatility process with finite number of states.
 In their procedure, they first fixed the volatility grids, computed the transition probability density at these grids based on the exponential of the generator matrix and used the inverse transform method to sample the volatility.
 After obtaining the volatility values, they computed the conditional characteristic function of the integrated variance, again based on the exponential of the generator matrix.
-They then adopted the Fourier sampler to get the CDF of the integrated variance, which first computed the PDF of the integrated variance via the orthogonal projection.
+They then adopted the Fourier sampler to get the CDF of the integrated variance, by first computing the PDF of the integrated variance via the orthogonal projection.
 Finally, they simulated the integrated variance using the closed-form inverse function.
-Putting aside the issue of accuracy of the simulation methods of $I_{t}^{h}$, these simulation methods invariably include some tedious and time-consuming procedures such as the adoption of the inverse Laplace/Fourier transform calculation, root-finding steps in the inverse transform methods or computation of the exponential of matrices.
+Besides the issue of accuracy of the simulation of $I_{t}^{h}$, these simulation methods invariably include tedious and time-consuming procedures such as the adoption of the inverse Laplace/Fourier transform calculation, root-finding steps in the inverse transform methods or computation of the exponential of matrices.
 However, our SLN simulation methods for $I_{t}^{h}$ exhibits easy and fast implementation.
-As shown in the numerical experiments reported in Section 5, our simulation methods for $I_{t}^{h}$ demonstrate significant savings in computational time and high level of accuracy.
+As shown in the numerical experiments reported in Section 5, our simulation methods for approximating $I_{t}^{h}$ demonstrate significant savings in computational time and high level of accuracy.
 4 Second simulation step: Sampling forward price $F_{t+h}$
-This section presents our CEV approximation for $F_{t+h}$ and its sampling method.
-4.1 CEV approximation of the conditional price
-Proposition 4 proposes our approximation of $F_{t+h}$ for the general case.
-Our choice of the approximation is a very natural and intuitive extension of the special cases discussed in Section.
-2.3.
-The only technical component is the choice of the conditional terminal forward price that we adopt from Remark 1.
-Proposition 4 (CEV approximation of the conditional terminal forward price).
-For $10<\beta<$ 1 and $-1\leq\rho\leq 1$, conditional on $\sigma_{t+h}$ and $I_{t}^{h}$ (and the filtration up to time $t$), we approximate $F_{t+h}$ as a CEV distribution
-$F_{t+h}\,|\,\sigma_{t+h},I_{t}^{h}\;\sim\;\mathcal{CEV}_{\beta}(\bar{F}_{t}^{h},\rho_{\ast}^{2}\sigma_{t}^{2}hI_{t}^{h}),$ (10)
+This section presents our CEV approximation for the forward price $F_{t+h}$ and the sampling of the CEV distribution via the exact simulation scheme of the shifted Poisson mixture gamma variate.
+4.1 CEV approximation of the conditional forward price
+In the uncorrelated SABR model ($\rho=0$), we observe that the conditional forward price $F_{T}$ follows the CEV distribution.
+To construct our SABR simulation scheme, we propose the CEV approximation of the conditional $F_{T}$ under the general case of $\rho\neq 0$ and $0<\beta<1$.
+We employ the operator splitting technique to derive the CEV approximation.
+Recall that a similar operator splitting approach has been used in the design of the simulation schemes for the rough Heston model in Bayer and Breneis (2024).
+We rewrite the SDEs for the SABR volatility model [See Eq. (1)] into the following form:
+$\frac{\text{d}F_{t}}{F_{t}^{\beta}}=\sigma_{t}(\rho\,\text{d}Z_{t}+\sqrt{1-\rho^{2}}\,\text{d}Z_{t}^{\perp})\quad\text{and}\quad\frac{\text{d}\sigma_{t}}{\sigma_{t}}=\nu\,\text{d}Z_{t},$ (10)
+where $Z_{t}$ and $Z_{t}^{\perp}$ are uncorrelated BMs. Similar to the operator splitting approximation method in Lileika and Mackevičius (2021), we split the above SDEs into two parts:
+$\frac{\text{d}F_{t}^{(1)}}{[F_{t}^{(1)}]^{\beta}}=\rho\sigma_{t}\,\text{d}Z_{t}\quad\text{and}\quad\frac{\text{d}\sigma_{t}}{\sigma_{t}}=\nu\,\text{d}Z_{t}$ (11a)
+and
+$\frac{\text{d}F_{t}^{(2)}}{[F_{t}^{(2)}]^{\beta}}=\sqrt{1-\rho^{2}}\sigma_{t}\,\text{d}Z_{t}^{\perp}\quad\text{and}\quad\frac{\text{d}\sigma_{t}}{\sigma_{t}}=\nu\,\text{d}Z_{t}.$ (11b)
+The operator splitting approximation method dictates that an approximation to the solution of $F_{t}$ in Eq.
+(10) with initial condition $F_{0}$ is given by $F_{t}\sim F_{t}^{(2)}\left(F_{t}^{(1)}(F_{0})\right)$, where $F_{t}^{(1)}(F_{0})$ is the solution of $F_{t}^{(1)}$ in Eq.
+(11a) with the initial condition $F_{0}^{(1)}=F_{0}$ and $F_{t}^{(2)}\left(F_{t}^{(1)}(F_{0})\right)$ is the solution of $F_{t}^{(2)}$ in Eq.
+(11b) with the initial condition $F_{0}^{(2)}=F_{t}^{(1)}(F_{0})$.
+Next, we solve the simplified set of SDEs for $F_{t}^{(1)}$ and $F_{t}^{(2)}$.
+The SDEs in Eq.
+(11b) resembles the SABR model with $\rho=0$ so that $F_{t}^{(2)}|\,\sigma_{t},I_{0}^{t}$ follows the CEV distribution, where
+$F_{t}^{(2)}|\,\sigma_{t},I_{0}^{t}\;\sim\;\mathcal{CEV}_{\beta}(F_{0}^{(2)},\rho_{\ast}^{2}\sigma_{0}^{2}hI_{0}^{t}),\quad\rho_{\ast}^{2}=1-\rho^{2}.$ (12)
+However, there is no simple analytic solution to Eq.
+(11a) for $\rho\neq 0$ and $0<\beta<1$.
+We propose to employ the frozen coefficient approximation to approximate $F_{t}^{(1)}]^{\beta_{\ast}}$ by $F_{0}^{\beta_{\ast}}$ such that Eq.
+(11a) is approximated by
+$\frac{\text{d}F_{t}^{(1)}}{F_{t}^{(1)}}=\frac{\rho\sigma_{t}}{[F_{t}^{(1)}]^{\beta_{\ast}}}\,\text{d}Z_{t}\approx\frac{\rho\sigma_{t}}{F_{0}^{\beta_{\ast}}}\,\text{d}Z_{t}.$ (13)
+Let $F_{t}^{(a)}$ denote the solution of the above approximating SDE, which is considered as an approximation of $F_{t}^{(1)}$. By rewriting the associated SDE for $F_{t}^{(a)}$ into the form:
+$\text{d}\ln F_{t}^{(a)}=\frac{\rho}{\nu F_{0}^{\beta_{\ast}}}\,\text{d}\sigma_{t}-\frac{\rho^{2}\sigma_{t}^{2}}{2F_{0}^{2\beta_{\ast}}}\,\text{d}t,$
+the solution of $F_{t}^{(a)}$ can be easily found by direct integration, which gives
+$F_{t}^{(a)}(F_{0})=F_{0}\exp\left(\frac{\rho}{\nu F_{0}^{\beta_{\ast}}}(\sigma_{t}-\sigma_{0})-\frac{\rho^{2}\sigma_{0}^{2}tI_{0}^{t}}{2F_{0}^{2\beta_{\ast}}}\right).$ (14)
+By combining the above operator splitting procedure coupled with frozen coefficient approximation, the approximation solution of $F_{t}$ is obtained as $F_{t}^{(2)}\left(F_{t}^{(a)}(F_{0})\right)$.
+Conditional on $\sigma_{t}$ and $I_{0}^{t}$, we obtain the CEV approximation of the conditional terminal forward price $F_{t}$ in the SABR model as follows:
+$F_{t}\,|\,\sigma_{t},I_{0}^{t}\;\sim\;\mathcal{CEV}_{\beta}(\bar{F}^{t}_{0},\rho_{\ast}^{2}\sigma_{0}^{2}tI_{0}^{t}),$ (15a)
+where
+$\bar{F}^{t}_{0}=F_{0}\exp\left(\frac{\rho}{\nu F_{0}^{\beta_{\ast}}}(\sigma_{t}-\sigma_{0})-\frac{\rho^{2}\sigma_{0}^{2}tI_{0}^{t}}{2F_{0}^{2\beta_{\ast}}}\right).$ (15b)
+The above discussion considers the approximation of solution of $F_{t}$ with initial time zero and end time $t$.
+In our context, we solve for $F_{t+h}$ with initial time $t$ and end time $t+h$.
+Accordingly, we change the initial time zero to time $t$ (comparing Eqs.
+(15a, 15b) with Eqs.
+(16a, 16b)).
+The CEV approximation of the conditional forward price $F_{t+h}\,|\,\sigma_{t+h},I_{t}^{h}$ is summarized in Algorithm 2:
+For $0<\beta<1$ and $1\leq\rho\leq 1$, conditional on $\sigma_{t+h}$ and $I_{t}^{h}$ (filtration up to time $t$), we approximate $F_{t+h}$ as a CEV distribution as follows:
+$F_{t+h}\,|\,\sigma_{t+h},I_{t}^{h}\;\sim\;\mathcal{CEV}_{\beta}(\bar{F}_{t}^{h},\rho_{\ast}^{2}\sigma_{t}^{2}hI_{t}^{h}),$ (16a)
 where the conditional expectation $\bar{F}_{t}^{h}$ is approximated by
-$\bar{F}_{t}^{h}\approx F_{t}\exp\left(\frac{\rho(\sigma_{t+h}-\sigma_{t})}{\nu F_{t}^{\beta_{\ast}}}-\frac{\rho^{2}\sigma_{t}^{2}hI_{t}^{h}}{2F_{t}^{2\beta_{\ast}}}\right).$ (11)
-Proof.
-We provide the rationales for each component of the CEV approximation.
-First, with regard to distribution, the choice of $\mathcal{CEV}_{\beta}$ is a natural extension of the $\nu=0$ and $\rho=0$ cases, where the forward price is distributed by $\mathcal{CEV}_{\beta}$ unconditionally and conditionally, respectively.
-Given that we choose the CEV distribution as the approximation, we show how to select the mean and variance.
-The variance, $\rho_{\ast}^{2}\sigma_{t}^{2}hI_{t}^{h}$, is also a natural choice, since it is consistent with all special cases in Section 2.3.
-The variance in the $\rho=0$ case is merely the special cases of $\rho_{\ast}^{2}\sigma_{t}^{2}hI_{t}^{h}$ when $\rho_{\ast}=1$.
-Lastly, we choose the conditional expectation $\bar{F}_{t}^{h}$ as the mean of the CEV distribution.
-Note that it is analytically intractable to obtain the exact conditional expectation under the original SABR dynamics when $\rho\neq 0$.
-Instead, we approximate the SABR dynamics in the interval $[\,t,t+h]$ by the lognormal SABR model:
-$\frac{\mathrm{d}F_{s}}{F_{s}}=\frac{\sigma_{s}}{F_{s}^{\beta_{\ast}}}\,\mathrm{d}W_{s}\approx\frac{\sigma_{s}}{F_{t}^{\beta_{\ast}}}\,\mathrm{d}W_{s}\quad\text{for}\quad t\leq s\leq t+h,$ (12)
-where the stochastic volatility $\sigma_{s}$ is replaced by $\sigma_{s}/F_{t}^{\beta_{\ast}}$.
-Under this approximation, the conditional expectation in Eq.
-(11) follows from $\beta=1$ case in Eq.
-(7).
-Note that the lognormal SABR approximation is too loose to be used for sampling $F_{t+h}$.
-We only use it for approximating $\bar{F}_{t}^{h}$.
-Most importantly, our choice of $\bar{F}_{t}^{h}$ ensures the martingale property of $F_{t}$.
-Based on Remark 1, we observe
+$\bar{F}_{t}^{h}\approx F_{t}\exp\left(\frac{\rho(\sigma_{t+h}-\sigma_{t})}{\nu F_{t}^{\beta_{\ast}}}-\frac{\rho^{2}\sigma_{t}^{2}hI_{t}^{h}}{2F_{t}^{2\beta_{\ast}}}\right).$ (16b)
+Algorithm 2 CEV approximation of the conditional terminal forward price
+Note that our choice of $\bar{F}_{t}^{h}$ ensures the martingale property of conditional terminal forward price $F_{t+h}$. Based on Remark 1, we observe
 $E(F_{t+h})=E\left(E(F_{t+h}|\sigma_{t+h},I_{t}^{h})\right)=E(\bar{F}_{t}^{h})=F_{t},$ (17)
-when $F_{t+h}$ is sampled according to this proposition. The martingale condition significantly improves the accuracy of our simulation algorithm (see numerical results illustrated in Figure 3). ∎
-4.2 Sampling conditional price from a CEV distribution
-The availability of analytic distribution function does not guarantee an efficient sampling algorithm of the random variable.
-The slow root-finding step is often used to invert the distribution function, and it has been the case with some existing SABR simulation methods (e.g., Cai et al.
+when $F_{t+h}$ is sampled according to Algorithm 2.
+The satisfaction of the martingale condition significantly improves the accuracy of our simulation algorithm (see numerical results illustrated in Figure 3).
+4.2 Sampling conditional terminal forward price from a CEV distribution
+With availability of analytic distribution function, the slow root-finding step is often used to invert the underlying distribution function.
+This approach has been adopted by some existing SABR simulation methods (e.g., Cai et al.
 (2017)).
-Our SABR simulation algorithm would be completed when our CEV approximation in Proposition 4 is paired with an efficient sampling algorithm of the CEV distribution.
-For the purpose, we adopt the algorithms of Makarov and Glew (2010) and Kang (2014) for sampling the CEV distribution exactly.
-Despite its simplicity and efficiency, the algorithm is not widely known in the literature.
-Before further discussion, let us present probability distributions that are related to the CEV distribution, and review several related analytic properties.
+Our SABR simulation algorithm would be completed when our CEV approximation of the forward price in Algorithm 2 is paired with an efficient sampling algorithm of the CEV distribution.
+It is fortunate that we can use the algorithms of Makarov and Glew (2010) and Kang (2014) for the exact simulation of the CEV distribution.
+Despite its simplicity and efficiency, this exact simulation scheme of the CEV distribution is not widely known in the literature.
+First, we present the relevant probability distributions that are related to the CEV distribution and several related analytic properties.
 Definition 3 (Noncentral chi-squared random variable).
-Let ${\chi^{2}}(\delta,r)$ be the NCX2 random variable with degree of freedom $\delta$ and noncentrality $r$. The PDF and CDF of ${\chi^{2}}(\delta,r)$ are respectively expressed by
+Let ${\chi^{2}}(\delta,r)$ be the NCX2 random variable with degree of freedom $\delta$ and noncentrality $r$. The PDF and CDF of ${\chi^{2}}(\delta,r)$ are respectively given by
 $f_{\chi^{2}}(x\,;\,\delta,r)=\frac{1}{2}\left(\frac{x}{r}\right)^{\alpha/2}I_{\alpha}(\sqrt{x\,r})\;e^{-(x+r)/2}\quad\text{and}\quad P_{\chi^{2}}(x;\delta,r)=\int_{0}^{x}f_{\chi^{2}}(t\,;\,\delta,r)\,\mathrm{d}t,$ (18)
@@
 Remark 6 (Relation to the gamma distribution).
-The central chi-squared (i.e., $r=0$) distribution is a gamma distribution.
+The central chi-squared (corresponding to $r=0$) distribution is a gamma distribution.
 Let $\mathcal{G}(\alpha)$ be a gamma random variable with shape parameter $\alpha$ and unit scale parameter, the two random variables are related by
 ${\chi^{2}}(\delta,0)\sim 2\,\mathcal{G}(\delta/2).$
-The PDF and CDF of $\mathcal{G}(\alpha)$ are respectively expressed by  [^3] The CDF, $P_{\mathcal{G}}(x;\alpha)$, is the scaled lower incomplete gamma function $\gamma(\alpha,x)$: $P_{\mathcal{G}}(x;\alpha)=\frac{\gamma(\alpha,x)}{\Gamma(\alpha)}\quad\text{where}\quad\gamma(\alpha,x)=\int_{0}^{x}t^{\alpha-1}e^{-t}\,\mathrm{d}t.$
+The PDF and CDF of $\mathcal{G}(\alpha)$ are respectively given by  [^2] The CDF of $\mathcal{G}(\alpha)$, $P_{\mathcal{G}}(x;\alpha)$, is the scaled lower incomplete gamma function $\gamma(\alpha,x)$: $P_{\mathcal{G}}(x;\alpha)=\frac{\gamma(\alpha,x)}{\Gamma(\alpha)}\quad\text{where}\quad\gamma(\alpha,x)=\int_{0}^{x}t^{\alpha-1}e^{-t}\,\mathrm{d}t.$
 $f_{\mathcal{G}}(x;\alpha)=\frac{1}{\Gamma(\alpha)}x^{\alpha-1}e^{-x}\quad\text{and}\quad P_{\mathcal{G}}(x;\alpha)=\frac{1}{\Gamma(\alpha)}\int_{0}^{x}t^{\alpha-1}e^{-t}\,\mathrm{d}t.$ (19)
-Definition 4 (Shifted Poisson random variable).
-The shifted Poisson (SP) variate with intensity $\lambda$ and shift $\alpha$, $N\sim\mathcal{SP}(\lambda,\alpha)$, takes non-negative integer values with mass probability function:
+Definition 4 (Shifted Poisson variate).
+The shifted Poisson (SP) variate with intensity $\lambda$ and shift parameter $\alpha$, $N\sim\mathcal{SP}(\lambda,\alpha)$, takes non-negative integer values with mass probability function:
 $P_{\mathcal{SP}}(n;\lambda,\alpha)=\mathrm{Prob}(N=n)=\frac{1}{P_{\mathcal{G}}(\lambda;\alpha)}\frac{\lambda^{\alpha+n}\,e^{-\lambda}}{\Gamma(n+\alpha+1)}.$ (20)
@@
 It is well known that the transition density of $F_{T}$ is closely related to the NCX2 distribution (Schroder, 1989; Cox, 1996).
-Proposition 5.
-Define $z_{t}$ be a transformation of the CEV process, $F_{t}$, in Eq. (8) by the function $z(\cdot)$:
-$z_{t}:=z(F_{t})\quad(0\leq t\leq T)\quad\text{and}\quad z(y):=\frac{y^{2\beta_{\ast}}}{\beta_{\ast}^{2}\sigma_{0}^{2}T}.$ (17)
+We define $z_{T}$ be a transformation of the CEV process $F_{T}$ (see Eq.
+(8b)) by the function $z(\cdot)$:
+$z_{T}:=z(F_{T})\quad\text{and}\quad z(y):=\frac{y^{2\beta_{\ast}}}{\beta_{\ast}^{2}\sigma_{0}^{2}T}.$ (21)
 The PDF of $z_{T}>0$ is expressed in terms of the NCX2 distribution:
@@
 Since the location variable $z$ plays the role of the noncentrality parameter, the distribution of $z_{T}$ is not a NCX2 distribution.
-The complementary CDF of $F_{T}>0$ and the probability of absorption (i.e., mass at zero) are respectively given by
-$\mathrm{Prob}(F_{T}>y)$ $=\mathrm{Prob}(z_{T}>z(y))=P_{\chi^{2}}\left(z_{0};\frac{1}{\beta_{\ast}},z(y)\right),$ (18)
-$\mathrm{Prob}(F_{T}=0)$ $=\mathrm{Prob}(z_{T}=0)=1-P_{\chi^{2}}\left(z_{0};\,\frac{1}{\beta_{\ast}},0\right)=1-P_{\mathcal{G}}\left(\frac{z_{0}}{2};\,\frac{1}{2\beta_{\ast}}\right).$ (19)
+The complementary CDF of $F_{T}>0$ and the probability of absorption (probability mass at zero) are respectively given by
+$\mathrm{Prob}(F_{T}>y)=\mathrm{Prob}(z_{T}>z(y))=P_{\chi^{2}}\left(z_{0};\frac{1}{\beta_{\ast}},z(y)\right),$ (22a)
+$\mathrm{Prob}(F_{T}=0)=\mathrm{Prob}(z_{T}=0)=1-P_{\chi^{2}}\left(z_{0};\,\frac{1}{\beta_{\ast}},0\right)=1-P_{\mathcal{G}}\left(\frac{z_{0}}{2};\,\frac{1}{2\beta_{\ast}}\right).$ (22b)
 The last equality in the second line comes from Remark 6.
-Proposition 6 (Mixture gamma representation of the CEV distribution (Makarov and Glew, 2010, § 3.3)).
-Under the CEV model with absorbing boundary condition, given $z_{T}>0$ ($F_{t}$ is not absorbed at the origin until time $T$), the transition from $z_{0}$ to $z_{T}$ can be sampled by an SP-mixture gamma distribution:
-$z_{T}\sim 2\,\mathcal{G}(N+1)\quad\text{where}\quad N\sim\mathcal{SP}\left(\frac{z_{0}}{2},\frac{1}{2\beta_{\ast}}\right)$
+Proposition 3 (Mixture gamma representation of the CEV distribution).
+Under the CEV model with absorbing boundary condition, given $z_{T}>0$ ($F_{t}$ is not absorbed at the origin until time $T$), the transition from $z_{0}$ to $z_{T}$ can be sampled by an SP-mixture gamma distribution (Makarov and Glew, 2010, § 3.3):
+$z_{T}\sim 2\,\mathcal{G}(N+1)\quad\text{where}\quad N\sim\mathcal{SP}\left(\frac{z_{0}}{2},\frac{1}{2\beta_{\ast}}\right).$ (23)
 Proof.
@@
 $=\frac{1}{2}\sum_{k=0}^{\infty}\frac{(z_{0}/2)^{k+\alpha}e^{-z_{0}/2}}{\Gamma(k+\alpha+1)}\frac{(z_{T}/2)^{k}}{k!}\;e^{-z_{T}/2}$
-$=P_{\mathcal{G}}\left(\frac{z_{0}}{2};\,\alpha\right)\sum_{k=0}^{\infty}P_{\mathcal{SP}}\left(k;\frac{z_{0}}{2},\alpha\right)\cdot\frac{1}{2}f_{\mathcal{G}}\left(\frac{z_{T}}{2};k+1\right).$
+$=P_{\mathcal{G}}\left(\frac{z_{0}}{2};\,\alpha\right)\sum_{k=0}^{\infty}P_{\mathcal{SP}}\left(k;\frac{z_{0}}{2},\alpha\right)\cdot\frac{1}{2}f_{\mathcal{G}}\left(\frac{z_{T}}{2};k+1\right),$
 so that the transition density of $z_{T}$ conditional on $z_{T}>0$ is given by
@@
 ∎
-Proposition 7 (Sampling an SP variable (Kang, 2014, Algorithm 3)).
-The SP variable, $N\sim\mathcal{SP}(\lambda,\alpha)$, can be sampled by a gamma-mixture Poisson variable:
-$N\sim\mathcal{POIS}(\lambda-X)\quad\text{where}\quad X\sim\mathcal{G}(\alpha)\quad\text{conditional on}\quad X<\lambda.$
-Proof.
-For proof, see Kang (2014). ∎
-Combining Propositions 6 and 7 leads to an exact simulation algorithm for sampling the CEV distribution.
-Moreover, Proposition 7 can be further optimized in the context of the CEV simulation.
+As proposed by Kang (2014), the SP variable, $N\sim\mathcal{SP}(\lambda,\alpha)$, can be sampled by a gamma-mixture Poisson variable:
+$N\sim\mathcal{POIS}(\lambda-X)\quad\text{where}\quad X\sim\mathcal{G}(\alpha)\quad\text{conditional on}\quad X<\lambda.$ (24)
+Combining Eqs.
+(23) and (24) leads to an exact simulation algorithm for sampling the CEV distribution.
+Moreover, Scheme (24) can be further optimized in the context of the CEV simulation.
 This is the reason we select ‘Algorithm 3’ among the three algorithms in Kang (2014).
-As this advantage is not mentioned in Kang (2014), we state explicitly it as a remark:
+As this advantage is not mentioned explicitly in Kang (2014), we state this feature as a remark:
 Remark 9 (Kang (2014)’s ‘Algorithm 3’ for the CEV distribution).
-If simulating $N\sim\mathcal{SP}(\lambda,\alpha)$ is the final goal, $X<\lambda$ can be obtained by repeatedly sampling $X$ until $X<\lambda$.
-If $N$ is an intermediate variable for the CEV simulation in Proposition 6, however, the repeated sampling of $X$ is unnecessary since
+Suppose simulating $N\sim\mathcal{SP}(\lambda,\alpha)$ is the final goal, $X<\lambda$ can be obtained by repeatedly sampling $X$ until $X<\lambda$.
+If $N$ is an intermediate variable for the CEV simulation in Scheme (23), however, the repeated sampling of $X$ is unnecessary since
 $\mathrm{Prob}(X\geq\lambda)=1-P_{\mathcal{G}}\left(\frac{z_{0}}{2};\,\frac{1}{2\beta_{\ast}}\right)=\mathrm{Prob}(F_{T}=0).$
-Therefore, we put $F_{T}=0$ if $X\geq\lambda$.
+Therefore, we set $F_{T}=0$ when $X\geq\lambda$.
 In summary, an exact simulation of the CEV distribution consists of successive generation of elementary random variables, for which efficient numerical algorithms are available in standard numerical library.
@@
 end if
-Algorithm 1 Exact simulation of the CEV distribution $F_{T}$ in Eq. (8)
-The CEV sampling algorithm can be applied to the conditional distribution of $F_{t+h}$ in Proposition 4.
-We need to sample $z_{t+h}=z(F_{t+h})$ from $z_{t}=z(\bar{F}_{t}^{h})$, where the transformation $z(\cdot)$ is modified in the context of Proposition 4:
+Algorithm 3 Exact simulation of the CEV distribution of $F_{T}$ as depicted in Eq. (8b)
+The CEV sampling algorithm can be applied to the conditional distribution of $F_{t+h}$ in Eqs.
+(16a) and (16b).
+We need to sample $z_{t+h}=z(F_{t+h})$ from $z_{t}=z(\bar{F}_{t}^{h})$, where the transformation $z(\cdot)$ is modified to become
 $z(y)=\frac{y^{2\beta_{\ast}}}{\beta_{\ast}^{2}\rho_{\ast}^{2}\sigma_{t}^{2}hI_{t}^{h}}.$
@@
  $\sigma_{t+h}\leftarrow$ Eq. (2) ;
-       $I_{t}^{h}\leftarrow$ Proposition 3 ;
-       $\bar{F}_{t}^{h}\leftarrow$ Eq. (11) ;
+ $I_{t}^{h}\leftarrow$ Algorithm 1 ;
+ $\bar{F}_{t}^{h}\leftarrow$ Eq. (16b) ;
  $z_{t}\leftarrow\frac{\left(\bar{F}_{t}^{h}\right)^{2\beta_{\ast}}}{\beta_{\ast}^{2}\rho^{2}\sigma_{t}^{2}hI_{t}^{h}}$;
@@
  if $X\geq z_{t}/2$ then
-             $F_{T}\leftarrow 0$ ;
+    $F_{t+h}\leftarrow 0$ ;
     break ;
@@
 Algorithm 4 Simulation of $F_{T}$ under the SABR model
-4.3 Comparison to Islah (2009)’s approximation used in other simulation schemes
-While earlier SABR simulation methods show various approach for sampling $I_{t}^{h}$ (Step 1), almost all of them (Chen et al., 2012; Cai et al., 2017; Leitao et al., 2017a, b; Grzelak et al., 2019; Cui et al., 2021; Kyriakou et al., 2023) are based on Islah (2009)’s approximation for sampling $F_{t+h}$ (Step 2).
+4.3 Comparison to Islah’s approximation used in other simulation schemes
+While the earlier SABR simulation methods use various approaches for sampling $I_{t}^{h}$ (Simulation step 1), almost all of them (Chen et al., 2012; Cai et al., 2017; Leitao et al., 2017a, b; Grzelak et al., 2019; Cui et al., 2021; Kyriakou et al., 2023) are based on Islah’s approximation for sampling $F_{t+h}$ (Simulation step 2).
 We compare our approach to Islah (2009)’s in this section.
-Proposition 8 (Islah (2009)’s approximation).
-For $10<\beta<$ 1, conditional on $\sigma_{t+h}$ and $I_{t}^{h}$ (and the filtration up to time $t$), the distribution of $F_{t+h}$ is approximated by
+We state Islah’s approximation as follows.
+For $0<\beta<1$, conditional on $\sigma_{t+h}$ and $I_{t}^{h}$ (and the filtration up to time $t$), the distribution of $F_{t+h}$ is approximated by
 $\mathrm{Prob}(F_{t+h}\geq y)\approx P_{\chi^{2}}\left(z^{\prime}_{t};\frac{1-\beta_{\ast}\rho^{2}}{\beta_{\ast}\rho_{\ast}^{2}},z(y)\right),$ (25)
 where
-$z^{\prime}_{t}=\frac{1}{\beta_{\ast}^{2}\rho_{\ast}^{2}\sigma_{t}^{2}hI_{t}^{h}}\left[F_{t}+\frac{\beta_{\ast}\rho}{\nu}(\sigma_{t+h}-\sigma_{t})\right]^{2}\quad\text{and}\quad z(y)=\frac{y^{2\beta_{\ast}}}{\beta_{\ast}^{2}\rho_{\ast}^{2}\sigma_{t}^{2}hI_{t}^{h}}.$
-Proof.
-For the ideas behind this approximation, see Islah (2009, § 2.2), Chen et al. (2012, Result 2.4), or Cui et al. (2021, Appendix B.1)
-Islah’s approximation is compared with the probability under our CEV approximation in Proposition 4:
+$z^{\prime}_{t}=\frac{1}{\beta_{\ast}^{2}\rho_{\ast}^{2}\sigma_{t}^{2}hI_{t}^{h}}\left[F_{t}^{\beta_{\ast}}+\frac{\beta_{\ast}\rho}{\nu}(\sigma_{t+h}-\sigma_{t})\right]^{2}\quad\text{and}\quad z(y)=\frac{y^{2\beta_{\ast}}}{\beta_{\ast}^{2}\rho_{\ast}^{2}\sigma_{t}^{2}hI_{t}^{h}}.$
+For the rationales behind this approximation, see Islah (2009, § 2.2), Chen et al. (2012, Result 2.4), or Cui et al. (2021, Appendix B.1).
+Islah’s approximation is compared with the probability under our CEV approximation in Eqs. (16a) and (16b):
 $\mathrm{Prob}(F_{t+h}\geq y)\approx P_{\chi^{2}}\left(z_{t};\frac{1}{\beta_{\ast}},z(y)\right)\quad\text{with}\quad z_{t}=z\left(\bar{F}_{t}^{h}\right).$
-The key difference between the two is that Islah’s approximation uses the modified degree of freedom, $(1-\beta_{\ast}\rho^{2})/(\beta_{\ast}\rho_{\ast}^{2})$, in Eq.
+The key difference between these two approximations is that Islah’s approximation uses the modified degree of freedom, $(1-\beta_{\ast}\rho^{2})/(\beta_{\ast}\rho_{\ast}^{2})$, in Eq.
 (25), and it is not equal to $1/\beta_{\ast}$ (except $\rho=0$).
 However, the transformation $z(\cdot)$ uses $\beta_{\ast}$ inconsistently.
-On the other hands, our CEV approximation uses $1/\beta_{\ast}$, which is the same degree of freedom for the CEV distribution.
+On the other hand, our CEV approximation uses $1/\beta_{\ast}$, which is the same degree of freedom for the CEV distribution.
 This means that $F_{t+h}$ under Islah’s approximation is not a CEV distribution, which leads to two drawbacks discussed in the next two remarks.
@@
 (2017a, b) artificially add an additive adjustment, commonly called the martingale correction, to the simulated samples of $F_{t+h}$.
-This ad-hoc correction is inferior to our approach where the martingale condition is naturally observed.
-Apart from the failure of martingale property in Islah’s approximation, a fast and exact algorithm for sampling $F_{t+h}$ according to Eq.
+This ad-hoc martingale correction is inferior when compared with our approach where the martingale condition is naturally observed.
+Apart from the failure of the martingale property in Islah’s approximation, a fast and exact algorithm for sampling $F_{t+h}$ according to Eq.
 (25) has not been available in the literature.
@@
 (25), which is proven to be effective in the Heston simulation (Andersen, 2008).
-While fast, it becomes inaccurate for a larger time step as it only uses the first two moments.
+While being fast, it becomes inaccurate for a larger time step as it only uses the first two moments.
 This approach belongs to the latter type.
-It is interesting to note that the exact CEV sampling algorithm in Section 4.2 can also be modified to sample $F_{t+h}$ under Islah’s approximation though it is not recommended due to the inherent drawbacks of Islah’s approximation.
-We outline the procedure in B.
+It is interesting to note that the exact CEV sampling algorithm in Section 4.2 can also be modified to sample $F_{t+h}$ under Islah’s approximation, though it is not recommended due to the inherent drawbacks of Islah’s approximation.
+We outline the simulation algorithm in B.
 With the use of the exact CEV sampling procedure, we numerically illustrate the supermartingale property of $F_{t}$ in Section 5.
 5 Numerical Results
-We demonstrate the accuracy and effectiveness of our simulation algorithm with an extensive set of numerical examples.
-Table 1 shows the five parameter sets used in our numerical tests.
-For easier comparison to other existing methods, we select the sets from previous literature: the first two from Antonov and Spector (2012) and the rest from Cai et al.
+We demonstrate numerical accuracy and effectiveness of our simulation algorithm with an extensive set of numerical examples.
+In Subsection 5.1, we examine separately the errors caused by the two steps of simulation approximation: (i) SLN approximation of the average variance, (ii) CEV approximation of the forward price distribution.
+We then compare accuracy and speed tradeoff with analytic approximation formula pf European option prices in Subsection 5.2 and other simulation schemes in Subsection 5.3.
+In the last subsection, we show the competitive edge of accuracy performance of our CEV approximation of the forward price distribution with that of the Islah’s approximation.
+For easier comparison to other existing methods, we select the sets from the previous literature: the first two from Antonov and Spector (2012) and the rest from Cai et al.
 (2017).
@@
 Pricing European options also benefits from the availability of highly accurate benchmark option prices obtained from using the finite difference method (FDM) (Cai et al., 2017).
-Cases $F_{0}$ $\sigma_{0}$ $\nu$ $\rho$ $\beta$ $T$ $K$
-Case I 1 0.25 0.3 $-0.8$ 0.3 10 [0.2, 2.0]
-Case II 1 0.25 0.3 $-0.5$ 0.6 10 [0.2, 2.0]
-Case III 0.05 0.4 0.6 0.0 0.3 1 [0.02, 0.10]
-Case IV 1.1 0.4 0.8 $-0.3$ 0.3 4 1.1
-Case V 1.1 0.3 0.5 $-0.8$ 0.4 [1, 10] 1.1
-Table 1: Parameter sets used in our numerical experiments.
 In our numerical tests, unless otherwise specified, we used $N=10^{5}$ paths for each simulation, and repeated $m=50$ times.
@@
 (2017) used Matlab 7 on an Intel Core2 Q9400 2.66GHZ processor.
-One may argue that comparing CPU time in this manner may not be fair as we used a newer system.
-However, the two CPUs are not significantly different to overturn the conclusion.
-As shown, our algorithm is faster than the existing algorithms by one or two orders of magnitude.
-5.1 Accuracy and comparison to analytic approximations
+We obtained the computer codes of Cai et al.
+(2017) and reran the codes again using our computer.
+The CPU times of their algorithm have been updated in Table 6 and Figure 2, where we examine the errors versus CPU time tradeoff.
+As shown in the comparison, our algorithm is seen to remain faster than the existing algorithms by at least one order of magnitude.
+5.1 Examination of numerical errors caused by separate simulation steps
+Recall that there are two separate simulation steps that may contribute to the overall errors in the full simulation scheme.
+The first step involves the SLN approximation of the average variance.
+The second step involves the CEV approximation of the forward price.
+The second step gives rise to errors from the operator splitting procedure and frozen coefficient approximation.
+The simulation of the CEV distribution via the SP mixture Gamma variate is exact, so this does not lead to simulation errors.
+We manage to isolate the errors that are separated into contributions from the SLN approximation and frozen coefficient approximation by considering accuracy performance of several degenerate cases of the SABR model.
+The accuracy analysis of the overall errors in the full simulation scheme in the general SABR model is discussed in Subsections 5.2 and 5.3.
+First, we consider the lognormal SABR model, where $\beta=1$.
+The corresponding conditional forward price $F_{T}$ follows the lognormal distribution (see Eq.
+(6)), so one does not need to introduce operator splitting and frozen coefficient procedures to obtain an approximation of conditional $F_{T}$.
+The lognormal SABR model serves to isolate the error contributed by the SLN approximation only.
+We performed accuracy analysis of European option prices computed using our simulation scheme with the SLN approximation and the benchmark numerical results obtained from the finite difference method (FDM).
+We compute the relative error using the formula:
+$\text{relative error}=\frac{\text{simulation calculation}-\text{FDM}}{\text{FDM}}.$
+In our calculations, we use the following parameter values: $\beta=1$, $T=1$, $\nu=0.2$, $\rho=-0.75$ and $h=1$ for the base case, and perturb the parameter values to $\rho=-0.5$, $\rho=-0.25$, $\nu=0.4$, $\nu=0.8$, $h=0.5$ and $h=0.25$ as well.
+The other parameter values include $\sigma_{0}=0.2$ and $F_{0}=K=1$.
+The relative errors of the simulation calculations are listed in Table 1.
+$\rho$ -0.75 -0.75 -0.75 -0.5 -0.25 -0.75 -0.75
+$\nu$ 0.2 0.2 0.2 0.2 0.2 0.4 0.6
+$h$ 1 0.5 0.25 1 1 1 1
+Rel. Err. 0.00353% 0.00489% 0.0110% 0.00700% 0.00275% 0.00808% 0.0198%
+Table 1: The relative errors contributed by the SLN approximation in the calculations of European call option prices for the lognormal SABR model ($\beta=1$).
+The FDM value for $\rho=-0.75$, $\nu=0.2$ is 0.07910, $\rho=-0.5$, $\nu=0.2$ is 0.07942, $\rho=-0.25$, $\nu=0.2$ is 0.07969, $\rho=-0.75$, $\nu=0.4$ is 0.07860 and $\rho=-0.75$, $\nu=0.6$ is 0.07811.
+We observe that the errors contributed by the SLN approximation for the calculations of European option prices for the lognormal SABR model ($\beta=1$) are exceedingly small, only a very small fraction of one percentage point.
+In most cases, the agreement of the simulation result and FDM value is up to 3 significant figures.
+The errors increase with higher value of $\nu$, but they remain to be very small.
+The decrease of time step $h$ is irrelevant in decreasing the errors.
+Indeed, errors increase due to roundoffs with more calculation due to an increase of the number of time steps.
+In conclusion, the SLN approximation of the average variance is highly accurate and contributes almost insignificant errors.
+Next, we consider the SABR models with $\rho=1$, $\rho=0.75$ and $\rho=0$ while $0<\beta<1$.
+The full correlation case ($\rho=1$) exhibits errors in calculating the European option prices contributed by the SLN approximation (should be exceedingly small from the above discussion) and frozen coefficient approximation, but no errors contributed from the operator splitting procedures.
+The conditional forward price of the zero correlation ($\rho=0$) SABR model follows the CEV distribution exactly, so there are no errors contributed by the operator splitting and frozen coefficient procedures.
+The errors arise only from the SLN approximation and simulation of the CEV distribution via the SP mixture Gamma variate.
+For $\rho=0.75$, it should have all three error sources: SLN approximation, operator splitting and frozen coefficient approximation.
+Table 2 lists the relative errors of the simulation calculations, which are benchmarked with the FDM results.
+Parameter values for $\sigma_{0}$, $F_{0}$, $K$ and $T$ are the same as those used in Table 1.
+$\rho=1$ $\nu=0.2$ $\nu=0.2$ $\nu=0.2$ $\nu=0.4$ $\nu=0.8$
+$\beta=0.4$ $\beta=0.6$ $\beta=0.8$ $\beta=0.8$ $\beta=0.8$
+$h=1$ 0.518% 0.348% 0.164% 0.404 0.746%
+$h=0.5$ 0.418% 0.225% 0.0683% 0.199 0.350%
+$h=0.25$ 0.244% 0.119% 0.0299% 0.0947 0.224%
+$\rho=0.75$ $\nu=0.2$ $\nu=0.2$ $\nu=0.2$ $\nu=0.4$ $\nu=0.8$
+$\beta=0.4$ $\beta=0.6$ $\beta=0.8$ $\beta=0.8$ $\beta=0.8$
+$h=1$ 0.415% 0.306% 0.125% 0.333% 0.421%
+$h=0.5$ 0.237% 0.130% 0.0647% 0.0773% 0.301%
+$h=0.25$ 0.0399% 0.0287% 0.0242% 0.0597% 0.215%
+$\rho=0$ $\nu=0.2$ $\nu=0.2$ $\nu=0.2$ $\nu=0.4$ $\nu=0.8$
+$\beta=0.4$ $\beta=0.6$ $\beta=0.8$ $\beta=0.8$ $\beta=0.8$
+$h=1$ -0.0562% -0.00574% 0.0704% 0.0257% 0.123%
+$h=0.5$ -0.0138% 0.0151% -0.0537% 0.0803% 0.0292%
+$h=0.25$ -0.0526% 0.0123% -0.0162% 0.0616% 0.0758%
+Table 2: The relative errors in calculating the European call option prices for the full correlation ($\rho=1$), $\rho=0.75$ and zero correlation ($\rho=0$) SABR models.
+The FDM value for $\rho=1$: $\nu=0.2\,,\beta=0.4$ is 0.07989, $\nu=0.2,\,\beta=0.6$ is 0.08002, $\nu=0.2,\,\beta=0.8$ is 0.08017, $\nu=0.4,\,\beta=0.8$ is 0.08044 and $\nu=0.8,\,\beta=0.8$ is 0.08043.
+The FDM values for $\rho=0.75$ are 0.07998, 0.08008, 0.08018, 0.08083 and 0.08276, respectively.
+The FDM values for $\rho=0$ are 0.07996, 0.07994, 0.07992, 0.08068 and 0.08355, respectively.
+For the full correlation ($\rho=1$) SABR model, the main source of errors comes from the frozen coefficient approximation.
+In all cases of different values of $\nu$ and $\beta$, the errors show decrease in magnitude with respect to the time step $h$.
+The frozen coefficient approximation is sensitive to the size of the time step $h$, smaller approximation error with smaller time step.
+Also, the errors show increase in magnitude with respect to $\beta_{\ast}=1-\beta$ and $\nu$.
+Such trends hold similarly for $\rho=0.75$, whose errors are from the three main sources: SLN approximation, operator splitting and frozen coefficient approximation.
+It is seen that the errors from operator splitting and frozen coefficient approximation may counteract with each other, leading to a smaller error when compared with the full correlation ($\rho=1$).
+On the other hand, for zero correlation ($\rho=0$) SABR model, the sources of errors come from the SLN approximation and exact simulation of the CEV distribution, both are exceedingly small (well below one percentage point).
+There is no pattern of dependence on $\nu$, $\beta$ and $h$ since errors are mostly from roundoff.
+5.2 Comparison to analytic approximations
+Cases $F_{0}$ $\sigma_{0}$ $\nu$ $\rho$ $\beta$ $T$ $K$
+Case I 1 0.25 0.3 0.8-0.8 0.3 10 [0.2, 2.0]
+Case II 1 0.25 0.3 0.5-0.5 0.6 10 [0.2, 2.0]
+Case III 0.05 0.4 0.6 0.0 0.3 1 [0.02, 0.10]
+Case IV 1.1 0.4 0.8 0.3-0.3 0.3 4 1.1
+Case V 1.1 0.3 0.5 0.8-0.8 0.4 [1, 10] 1.1
+Table 3: Parameter sets used in our numerical experiments.
+Table 3 shows the five parameter sets used in our numerical tests.
 In this subsection, we test the accuracy of our algorithm (Algorithm 4) using Cases I and II in Table 3.
-Tables 2 and 3 show the numerical results for the two sets.
+Tables 4 and 5 show the numerical results for the two sets.
 The two examples are previously tested in Antonov and Spector (2012) to demonstrate the performance of the cutting-edge analytic approximations for European vanilla options, map to the zero correlation (ZC Map, and hybrid map to the zero correlation (Hyb ZC Map), against the widely used Hagan et al.
@@
 Therefore, we also display those option prices for comparison.
-Tables 2 and 3 demonstrate that our simulation scheme is highly accurate.
+Tables 4 and 5 demonstrate that our simulation scheme is highly accurate.
 As we decrease the time step $h$ from 1 to $1/16$, the option price converges to its true price.
@@
 This is probably due to the observation that the geometric BM approximation for the conditional mean $\bar{F}_{0}^{T}$ in Eq.
-(12) holds better as $\beta=0.6$ (Case II) is closer to one than $\beta=0.3$ (Case I).
+(13) holds better as $\beta=0.6$ (Case II) is closer to $\beta=1$ (lognormal SABR model) than $\beta=0.3$ (Case I).
 $K/F_{0}$ 0.2 0.4 0.8 1.0 1.2 1.6 2.0
@@
 FDM 0.84255 0.68906 0.40646 0.28502 0.18304 0.05343 0.01096
-Bias of our simulation method  (×$(\times 10^{-3})$
-$h=1$1.22 -1.49 -0.37 0.49 1.28 1.72 1.32
-$h=1/4$0.46 -0.24 0.22 0.42 0.56 0.56 0.48
-$h=1/16$0.34 -0.20 0.00 0.05 0.11 0.10 0.10
-Standard deviation of our simulation method  (×$(\times 10^{-3})$
+Bias of our simulation method  $(\times 10^{-3})$
+$h=1$ -1.22 -1.49 -0.37 0.49 1.28 1.72 1.32
+$h=1/4$ -0.46 -0.24 0.22 0.42 0.56 0.56 0.48
+$h=1/16$ -0.34 -0.20 0.00 0.05 0.11 0.10 0.10
+Standard deviation of our simulation method  $(\times 10^{-3})$
 $h=1$ 1.97 1.83 1.50 1.31 1.08 0.63 0.38
@@
 $h=1/16$ 1.89 1.75 1.44 1.28 1.06 0.53 0.22
-Bias of various analytic approximation methods  (×$(\times 10^{-3})$
+Bias of various analytic approximation methods  $(\times 10^{-3})$
 Hagan 22.35 23.64 17.94 13.81 9.38 2.54 0.82
@@
 FDM 0.82886 0.66959 0.39772 0.29118 0.20690 0.10018 0.05014
-Bias of our simulation method   (×$(\times 10^{-3})$
-$h=1$0.14 -0.30 -0.42 -0.43 -0.43 -0.40 -0.30
+Bias of our simulation method   $(\times 10^{-3})$
+$h=1$ -0.14 -0.30 -0.42 -0.43 -0.43 -0.40 -0.30
 $h=1/4$ 0.45 0.37 0.27 0.20 0.10 -0.02 0.00
 $h=1/16$ 0.01 -0.01 0.02 0.04 0.03 0.00 -0.03
-Stdev of our simulation method   (×$(\times 10^{-3})$
+Stdev of our simulation method   $(\times 10^{-3})$
 $h=1$ 2.23 2.09 1.78 1.65 1.51 1.20 0.93
@@
 $h=1/16$ 2.46 2.32 2.01 1.79 1.58 1.22 0.97
-Bias of various analytic approximation methods   (×$(\times 10^{-3})$
+Bias of various analytic approximation methods   $(\times 10^{-3})$
 Hagan 11.65 15.94 16.69 14.67 12.18 8.56 6.88
@@
 In this subsection, we demonstrate the computational efficiency of our method by comparing the accuracy-speed trade-off with the earlier simulation methods, such as the Euler, low-bias (Chen et al., 2012), and piecewise semi-exact (PSE) (Cai et al., 2017) schemes.
-For the purpose, we use Cases III and IV.
-These cases have been tested by Cai et al.
-(2017), and we take advantage of the extensive records of the CPU time reported therein.
+We use Cases III and IV to achieve better comparison since these cases have been tested by Cai et al.
+(2017).
+We take advantage of the extensive records of the CPU time reported therein.
 $K/F_{0}$ 0.4 0.8 1 1.2 1.6 2 Time (s)
@@
 FDM 0.04559 0.04141 0.03942 0.03750 0.03390 0.03061
-Error (×$(\times 10^{-3})$ and CPU time of our simulation method
+Error $(\times 10^{-3})$ and CPU time of our simulation method
 $h=1$ 0.00 0.00 0.00 0.00 -0.01 -0.01 0.03
-Bias (×$(\times 10^{-3})$ and CPU time of the Euler scheme reported in Cai et al. (2017)
-$h=1/400$ 1.6 1.5 1.5 1.4 1.3 1.2 49.4
-$h=1/800$ 0.7 0.6 0.5 0.5 0.4 0.3 99.1
-$h=1/1600$0.3 -0.3 -0.3 -0.3 -0.3 -0.3 194.0
-Bias (×$(\times 10^{-3})$ and CPU time of the low-bias scheme reported in Cai et al. (2017)
-$h=1/4$ 0.5 0.5 0.5 0.4 0.4 0.4 78.4
-$h=1/8$ 0.4 0.4 0.4 0.3 0.3 0.2 175.8
-Bias (×$(\times 10^{-3})$ and CPU time of the PSE scheme reported in Cai et al. (2017)
-$h=1$ 0.1 0.2 0.2 0.2 0.2 0.2 98.3
+Bias $(\times 10^{-3})$ and CPU time of the Euler scheme reported in Cai et al. (2017)
+$h=1/400$ 1.6 1.5 1.5 1.4 1.3 1.2 4.12
+$h=1/800$ 0.7 0.6 0.5 0.5 0.4 0.3 8.26
+$h=1/1600$ -0.3 -0.3 -0.3 -0.3 -0.3 -0.3 16.2
+Bias $(\times 10^{-3})$ and CPU time of the low-bias scheme reported in Cai et al. (2017)
+$h=1/4$ 0.5 0.5 0.5 0.4 0.4 0.4 6.54
+$h=1/8$ 0.4 0.4 0.4 0.3 0.3 0.2 14.7
+Bias $(\times 10^{-3})$ and CPU time of the PSE scheme reported in Cai et al. (2017)
+$h=1$ 0.1 0.2 0.2 0.2 0.2 0.2 8.19
 Table 6: The European option prices from our simulation methods for Case III, compared with other MC methods: Euler, low-bias (Chen et al., 2012), and PSE (Cai et al., 2017) schemes.
@@
 All simulation methods used $N=100,000$ paths in each simulation run and repeated $m=50$ times.
-The results for Case III in Table 4 show that our method is at least several hundred times faster than the competitors while the bias is much lower.
-Note that the correlation is zero in this case.
-When $\rho=0$, both our CEV approximation and Islah’s approximation become exact, and the source of bias is the accuracy of sampling $I_{t}^{h}$.
-The near-zero error of our method validates our SLN approximation for $I_{t}^{h}$.
+The results for Case III in Table 6 show that our method is at least one hundred times faster than the competitors while the bias is much lower.
+Under $\rho=0$ in this case, both our CEV approximation and Islah’s approximation become exact, and the source of bias is the accuracy of sampling $I_{t}^{h}$.
+The near-zero error of our method validates extremely high accuracy of our SLN approximation for $I_{t}^{h}$.
 For the Euler and low-bias schemes, it is necessary to decrease the time step $h$ for more accurate sampling of $I_{t}^{h}$, resulting with slower computation.
 Although the PSE scheme exhibits good accuracy for single step ($h=1$), it requires prohibitive computation for numerical inverse transform for sampling $I_{t}^{h}$.
-$N$ $h$ RMS error (×$\times 10^{-3}$) Time (s)
+NN $h$ RMS error ($\times 10^{-3}$) Time (s)
 160,000 1 3.27 0.53
@@
 2,560,000 1/16 0.59 279.53
-Table 5: The root mean square (RMS) error and CPU time from Case IV when the number of paths $N$ is doubled and time step $h$ is halved.
+Table 7: The root mean square (RMS) error and CPU time from Case IV when the number of paths NN is doubled and time step $h$ is halved.
 Next, we perform CPU time versus root-mean-square (RMS) error analysis on Case IV. The RMS error is defined by
-RMS $\text{RMS error}=\sqrt{\text{Stdev}^{2}+\text{Bias}^{2}}.$ (21)
+$\text{RMS error}=\sqrt{\text{Stdev}^{2}+\text{Bias}^{2}}.$ (26)
 Table 7 presents the RMS error and CPU time for decreasing time step $h$ and increasing number of paths $N$.
@@
 5.4 Comparison with Islah’s approximation
-In this last numerical experiment using Case V, we demonstrate superiority of our CEV approximation of conditional $F_{T}$ in Section 4.1 to Islah’s approximation which has been commonly adopted in other schemes.
+In this last numerical experiment using Case V, we demonstrate superiority of our CEV approximation of conditional $F_{T}$ in Section 4.1 to Islah’s approximation which has been commonly adopted in other simulation schemes.
 The parameter set is also taken from Cai et al.
@@
 On one hand, our simulation scheme consists of (i) SLN sampling of $I_{t}^{h}$ and (ii) CEV model sampling of conditional $F_{T}$.
-On the other hand, we replace the second step with Islah’s approximation (Proposition 8).
+On the other hand, we replace the second step with Islah’s approximation (Eq.
+(25)).
 Under Islah’s approximation, a power transformation of $F_{T}$ observes a CEV distribution; see Eq.
@@
 Figure 3(a) shows the error of the terminal price $E(F_{T})$, which should be equal to $F_{0}=1.1$ in theory.
-As expected, the results show that our approach preserves the martingale property fairly well for all time-to-maturity $T$.
+As expected, the results show that our approach preserves the martingale property of the forward price fairly well for all time-to-maturity $T$.
 That is, $E(F_{T})=F_{0}$ holds with very high accuracy regardless of $h$.
 In Islah’s approach, however, $E(F_{T})$ deviates from $F_{0}$.
-In particular, the deviation accumulates as the time-to-maturity $T$ becomes longer.
+In particular, the deviation from $F_{0}$ accumulates as the time-to-maturity $T$ becomes longer.
 The time step $h$ has to be smaller to reduce the deviation.
@@
 Although researchers proposed various simulation schemes that overcome the limitation of the time-discretization schemes, the schemes have their own limitations.
-In sampling conditional average variance, the schemes are riddled with either heavy computation (Cai et al., 2017; Leitao et al., 2017a) or inaccurate approximation over a large simulation step (Chen et al., 2012).
-In sampling conditional forward price, almost all methods adopt Islah (2009)’s NCX2 approximation that neither preserves martingale property nor provides efficient sampling algorithm.
+In sampling conditional average variance, the simulation schemes are riddled with either heavy computation (Cai et al., 2017; Leitao et al., 2017a) or inaccurate approximation over a large simulation step (Chen et al., 2012).
+In sampling conditional forward price, almost all simulation algorithms adopt Islah (2009)’s NCX2 approximation that neither preserves martingale property nor provides efficient sampling algorithm.
 This paper present a series of innovations on the SABR simulation.
@@
 Numerical results show that our methods are highly efficient and accurate even under challenging parameter values, like large time-to-maturity $T$, when compared with other simulation methods.
-Funding
-The works of Lilian Hu and Yue Kuen Kwok are supported by the Guangzhou-HKUST (GZ) Joint Funding Program (Number 2024A03J0630).
+Acknowledgment
+The works of Lilian Hu and Yue Kuen Kwok are supported by the Guangzhou-HKUST (GZ) Joint Funding Program (Number 2024A03J0630). The work of Yue Kuen Kwok is also supported by Project 2023CX10X1.
 Appendix A Conditional moments of $I_{t}^{h}$
-The conditional average variance $I_{t}^{h}(\hat{Z})$ is closely related to the exponential functional of BM, defined by
+[…]
 $A_{t}:=\int_{0}^{t}e^{2Z_{s}}\,\mathrm{d}s,$
-where $Z_{s}$ is a standard Brownian motion. The derivation of Proposition 2 is based on the analytic conditional moments of the exponential functional of BM (Matsumoto and Yor, 2005, (5.4)):
-$E\left(\left(A_{t}\right)^{n}|Z_{t}=x\right)=\frac{e^{nx}}{n!t}\int_{x}^{\infty}be^{(x^{2}-b^{2})/2t}\left[\cosh(b)-\cosh(x)\right]^{n}\mathrm{d}b.$
-The conditional average variance $I_{t}^{h}(\hat{Z})$ is expressed by $A_{t}$ conditional on the terminal location of the BM, $Z_{t}$:
+where $Z_{s}$ is a standard BM. The derivation of Proposition 2 is based on the analytic conditional moments of the exponential functional of BM (Matsumoto and Yor, 2005, (5.4)):
+$E\left(\left(A_{t}\right)^{n}|Z_{t}=x\right)=\frac{e^{nx}}{n!t}\int_{x}^{\infty}be^{(x^{2}-b^{2})/2t}\left(\cosh b-\cosh x\right)^{n}\mathrm{d}b.$
+[…]
 $I_{t}^{h}(\hat{Z})=\frac{1}{\hat{\nu}^{2}}A_{\hat{\nu}^{2}}\,\Big|_{Z_{\hat{\nu}^{2}}=\hat{\nu}\hat{Z}}\,\quad\left(\hat{\nu}=\nu\sqrt{h}\right).$
-Therefore, the conditional moments admit the following integral form:
+Accordingly, the conditional moments of $I_{t}^{h}$ admit the following integral form:
 $\mu^{\prime}_{k}$ $=E\left(\left(I_{t}^{h}\right)^{k}\right)=\frac{1}{\hat{\nu}^{2k}}E\left[\left(A_{\hat{\nu}^{2}}\right)^{k}\,\big|_{Z_{\hat{\nu}^{2}}=\hat{\nu}\hat{Z}}\right$]
-$=\frac{e^{k\hat{\nu}\hat{Z}}}{k!\,\hat{\nu}^{2k}}\int_{\hat{Z}}^{\infty}se^{(\hat{Z}^{2}-s^{2})/2}\left[\cosh(\hat{\nu}s)-\cosh(\hat{\nu}\hat{Z})\right]^{k}\mathrm{d}s$
-$=\frac{e^{k\hat{\nu}\hat{Z}}}{(k-1)!\,\hat{\nu}^{2k-1}}\int_{\hat{Z}}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\sinh(\hat{\nu}s)\left[\cosh(\hat{\nu}s)-\cosh(\hat{\nu}\hat{Z})\right]^{k-1}\mathrm{d}s,$
+$=\frac{e^{k\hat{\nu}\hat{Z}}}{k!\,\hat{\nu}^{2k}}\int_{\hat{Z}}^{\infty}se^{(\hat{Z}^{2}-s^{2})/2}\left(\cosh\hat{\nu}s-\cosh\hat{\nu}\hat{Z}\right)^{k}\mathrm{d}s$
+$=\frac{e^{k\hat{\nu}\hat{Z}}}{(k-1)!\,\hat{\nu}^{2k-1}}\int_{\hat{Z}}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\sinh\hat{\nu}s\left(\cosh\hat{\nu}s-\cosh\hat{\nu}\hat{Z}\right)^{k-1}\mathrm{d}s,$
 where we apply the change of variable, $s=\hat{\nu}b$, to obtain the second line, and integration by part to obtain the third line.
 Using the following analytic integral
-$\int_{z}^{\infty}e^{-s^{2}/2}\sinh(as)\,\mathrm{d}s=\frac{N(-z+a)-N(-z-a)}{2n(a)}=\frac{N(z+a)-N(z-a)}{2n(a)},$
-we obtain for $k\neq 0$,
-$\int_{\hat{Z}}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\sinh(k\hat{\nu}s)\,\mathrm{d}s=k\hat{\nu}\,m_{k}(\hat{Z})\quad\text{where}\quad m_{k}(\hat{Z}):=\frac{N(\hat{Z}+k\hat{\nu})-N(\hat{Z}-k\hat{\nu})}{2k\hat{\nu}\;n\left(\sqrt{\hat{Z}^{2}+(k\hat{\nu})^{2}}\right)},$
-which will be useful in the derivation below. Here, we define $m_{k}(\hat{Z})$ in such a way that $m_{k}(\hat{Z})\to 1$ for any non-zero $k$ as $\hat{\nu}\to 0$.
-Next, we evaluate $E((I_{t}^{h})^{k})$. For $k=1$, we obtain
-$\mu=\frac{e^{\hat{\nu}\hat{Z}}}{\hat{\nu}}\int_{\hat{Z}}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\sinh(\hat{\nu}s)\,\mathrm{d}s=e^{\hat{\nu}\hat{Z}}\,m_{1}(\hat{Z})=\left(\frac{\sigma_{t+h}}{\sigma_{t}}\right)\,m_{1}.$
-For $k=2$, using $\sinh(x)\cosh(x)=\frac{1}{2}\sinh(2x)$, we have
-$\mu^{\prime}_{2}$ $=\frac{e^{2\hat{\nu}\hat{Z}}}{\hat{\nu}^{3}}\int_{z}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\sinh(\hat{\nu}s)\left[\cosh(\hat{\nu}s)-\cosh(\hat{\nu}z)\right]\,\mathrm{d}s$
-$=\frac{e^{2\hat{\nu}\hat{Z}}}{\hat{\nu}^{3}}\int_{z}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\left[\frac{\sinh(2\hat{\nu}s)}{2}-\sinh(\hat{\nu}s)\cosh(\hat{\nu}z)\right]\,\mathrm{d}s$
-$=\left(\frac{\sigma_{t+h}}{\sigma_{t}}\right)^{2}\frac{1}{\hat{\nu}^{2}}\left[m_{2}-cm_{1}\right].$
-For $k=3$, using $\sinh(x)\cosh^{2}(x)=\frac{1}{4}\left[\sinh(3x)+\sinh(x)\right]$, we obtain
-$\mu^{\prime}_{3}$ $=\frac{e^{3\hat{\nu}\hat{Z}}}{2\,\hat{\nu}^{5}}\int_{z}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\sinh(\hat{\nu}s)\left[\cosh(\hat{\nu}s)-\cosh(\hat{\nu}z)\right]^{2}\,\mathrm{d}s$
-$=\frac{e^{3\hat{\nu}\hat{Z}}}{2\,\hat{\nu}^{5}}\int_{z}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\left[\frac{\sinh(3\hat{\nu}s)+\sinh(\hat{\nu}s)}{4}-\sinh(2\hat{\nu}s)\cosh(\hat{\nu}z)+\sinh(\hat{\nu}s)\cosh^{2}(\hat{\nu}z)\right]\,\mathrm{d}s$
+$\int_{z}^{\infty}e^{-s^{2}/2}\sinh as\,\mathrm{d}s=\frac{N(-z+a)-N(-z-a)}{2n(a)}=\frac{N(z+a)-N(z-a)}{2n(a)},$
+for $k\neq 0$, we obtain
+$\int_{\hat{Z}}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\sinh k\hat{\nu}s\,\mathrm{d}s=k\hat{\nu}\,m_{k}(\hat{Z})\quad\text{where}\quad m_{k}(\hat{Z}):=\frac{N(\hat{Z}+k\hat{\nu})-N(\hat{Z}-k\hat{\nu})}{2k\hat{\nu}\;n\left(\sqrt{\hat{Z}^{2}+(k\hat{\nu})^{2}}\right)},$
+[…]
+Next, we evaluate the moments $E((I_{t}^{h})^{k})$, $k=1,2,3,4$. For $k=1$, we obtain
+$\mu=\frac{e^{\hat{\nu}\hat{Z}}}{\hat{\nu}}\int_{\hat{Z}}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\sinh\hat{\nu}s\,\mathrm{d}s=e^{\hat{\nu}\hat{Z}}\,m_{1}(\hat{Z})=\left(\frac{\sigma_{t+h}}{\sigma_{t}}\right)\,m_{1}.$
+For $k=2$, using $\sinh x\cosh x=\frac{1}{2}\sinh 2x$, we have
+$\mu^{\prime}_{2}$ $=\frac{e^{2\hat{\nu}\hat{Z}}}{\hat{\nu}^{3}}\int_{z}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\sinh\hat{\nu}s\left(\cosh\hat{\nu}s-\cosh\hat{\nu}z\right)\,\mathrm{d}s$
+$=\frac{e^{2\hat{\nu}\hat{Z}}}{\hat{\nu}^{3}}\int_{z}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\left(\frac{\sinh 2\hat{\nu}s}{2}-\sinh\hat{\nu}s\cosh\hat{\nu}z\right)\,\mathrm{d}s$
+$=\left(\frac{\sigma_{t+h}}{\sigma_{t}}\right)^{2}\frac{1}{\hat{\nu}^{2}}\left(m_{2}-cm_{1}\right).$
+For $k=3$, using $\sinh x\cosh^{2}x=\frac{1}{4}\left(\sinh 3x+\sinh x\right)$, we obtain
+$\mu^{\prime}_{3}$ $=\frac{e^{3\hat{\nu}\hat{Z}}}{2\,\hat{\nu}^{5}}\int_{z}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\sinh\hat{\nu}s\left(\cosh\hat{\nu}s-\cosh\hat{\nu}z\right)^{2}\,\mathrm{d}s$
+$=\frac{e^{3\hat{\nu}\hat{Z}}}{2\,\hat{\nu}^{5}}\int_{z}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\left(\frac{\sinh 3\hat{\nu}s+\sinh\hat{\nu}s}{4}-\sinh 2\hat{\nu}s\cosh\hat{\nu}z+\sinh\hat{\nu}s\cosh^{2}\hat{\nu}z\right)\,\mathrm{d}s$
 $=\left(\frac{\sigma_{t+h}}{\sigma_{t}}\right)^{3}\frac{1}{8\hat{\nu}^{4}}\left[3m_{3}-8cm_{2}+\left(4c^{2}+1\right)m_{1}\right].$
-For $k=4$, using $\sinh(x)\cosh^{3}(x)=\frac{1}{8}\left[\sinh(4x)+2\sinh(2x)\right]$, we have
-$\mu^{\prime}_{4}$ $=\frac{e^{4\hat{\nu}\hat{Z}}}{6\,\hat{\nu}^{7}}\int_{z}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\sinh(\hat{\nu}s)\left[\cosh(\hat{\nu}s)-\cosh(\hat{\nu}z)\right]^{3}\,\mathrm{d}s$
-$=\frac{e^{4\hat{\nu}\hat{Z}}}{6\,\hat{\nu}^{7}}\int_{z}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\left\{\frac{\sinh(4\hat{\nu}s)+2\sinh(2\hat{\nu}s)}{8}-\frac{3}{4}\left[\sinh(3\hat{\nu}s)+\sinh(\hat{\nu}s)\right]\cosh(\hat{\nu}z)\right.$
-$\hskip 120.00018pt\left.+\,\frac{3}{2}\sinh(2\hat{\nu}s)\cosh^{2}(\hat{\nu}z)-\sinh(\hat{\nu}s)\cosh^{3}(\hat{\nu}z)\right\}\,\mathrm{d}s$
+For $k=4$, using $\sinh x\cosh^{3}x=\frac{1}{8}\left(\sinh 4x+2\sinh 2x\right)$, we have
+$\mu^{\prime}_{4}$ $=\frac{e^{4\hat{\nu}\hat{Z}}}{6\,\hat{\nu}^{7}}\int_{z}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\sinh\hat{\nu}s\left(\cosh\hat{\nu}s-\cosh\hat{\nu}z\right)^{3}\,\mathrm{d}s$
+$=\frac{e^{4\hat{\nu}\hat{Z}}}{6\,\hat{\nu}^{7}}\int_{z}^{\infty}e^{(\hat{Z}^{2}-s^{2})/2}\left[\frac{\sinh 4\hat{\nu}s+2\sinh 2\hat{\nu}s}{8}-\frac{3}{4}\left(\sinh 3\hat{\nu}s+\sinh\hat{\nu}s\right)\cosh\hat{\nu}z\right$.
+$\hskip 120.00018pt\left.+\,\frac{3}{2}\sinh 2\hat{\nu}s\cosh^{2}\hat{\nu}z-\sinh\hat{\nu}s\cosh^{3}\hat{\nu}z\right]\,\mathrm{d}s$
 $=\left(\frac{\sigma_{t+h}}{\sigma_{t}}\right)^{4}\frac{1}{24\,\hat{\nu}^{6}}\left[2m_{4}-9cm_{3}+\left(12c^{2}+2\right)m_{2}-c\left(4c^{2}+3\right)m_{1}\right].$
 In addition, the expansion of the mean, coefficient of variance, skewness, and ex-kurtosis (see Remark 5) around $\hat{\nu}=0$ are obtained as
-$v=\frac{\hat{\nu}}{\sqrt{3}}+O(\hat{\nu}^{3}),\quad s=\frac{6\sqrt{3}}{5}\hat{\nu}+O(\hat{\nu}^{3}),\quad\text{and}\quad\kappa=\frac{276}{35}\hat{\nu}^{2}+O(\hat{\nu}^{4}).$
+$Cv=\frac{\hat{\nu}}{\sqrt{3}}+O(\hat{\nu}^{3}),\quad s=\frac{6\sqrt{3}}{5}\hat{\nu}+O(\hat{\nu}^{3}),\quad\text{and}\quad\kappa=\frac{276}{35}\hat{\nu}^{2}+O(\hat{\nu}^{4}).$
 Appendix B Sampling Islah’s approximation via a CEV distribution
-Though Islah’s approximation is less than desirable, we show that Islah’s distribution of $F_{t+h}$ can be expressed as a power of a CEV random variable.
-Therefore, the CEV simulation algorithm in Section 4.2 can be applied to expedite the sampling.
-Let us define
+Though Islah’s approximation of $F_{t+h}$ is less than desirable since it fails the martingale property, nevertheless, we show that Islah’s approximation can be expressed as a power function of a CEV random variable.
+For theoretical interest, we show that the CEV simulation algorithm in Section 4.2 can be applied to expedite the sampling of $F_{t+h}$ based on Islah’a approximation.
+We define
 $\beta^{\prime}:=\frac{\beta}{1-\beta_{\ast}\rho^{2}}\;(\geq\beta),\quad\beta_{\ast}^{\prime}:=1-\beta^{\prime}=\frac{\beta_{\ast}\rho_{\ast}^{2}}{1-\beta_{\ast}\rho^{2}}\;(\leq\beta_{\ast}),$
@@
 $\frac{(y^{\prime})^{2\beta_{\ast}^{\prime}}}{(\beta_{\ast}^{\prime})^{2}}=\frac{y^{2\beta_{\ast}}}{\beta_{\ast}^{2}}\quad\text{and}\quad\frac{\left(\bar{F^{\prime}}_{t}^{h}\right)^{2\beta_{\ast}^{\prime}}}{\beta_{\ast}^{\prime 2}}=\frac{1}{\beta_{\ast}^{2}}\left[F_{t}^{\beta_{\ast}}+\frac{\beta_{\ast}\rho}{\nu}(\sigma_{t+h}-\sigma_{t})\right]^{2}.$
-As a result, the complementary CDF of Islah (2009)’s approximation is equivalently expressed by
-$\mathrm{Prob}(F_{t+h}\geq y)=\mathrm{Prob}\left(\left(\frac{\beta_{\ast}^{\prime}}{\beta_{\ast}}F_{t+h}^{\beta_{\ast}}\right)^{1/\beta_{\ast}^{\prime}}\geq y^{\prime}\right)=P_{\chi^{2}}\left(\frac{\left(\bar{F^{\prime}}_{t}^{h}\right)^{2\beta_{\ast}^{\prime}}}{\beta_{\ast}^{\prime 2}\rho_{\ast}^{2}\sigma_{0}^{2}TI_{0}^{T}};\frac{1}{\beta_{\ast}^{\prime}},\frac{\left(y^{\prime}\right)^{2\beta_{\ast}^{\prime}}}{\beta_{\ast}^{\prime 2}\rho_{\ast}^{2}\sigma_{0}^{2}TI_{0}^{T}}\right),$
+As a result, the complementary CDF of Islah’s approximation is equivalently expressed by
+$\mathrm{Prob}(F_{t+h}\geq y)=\mathrm{Prob}\left(\left(\frac{\beta_{\ast}^{\prime}}{\beta_{\ast}}F_{t+h}^{\beta_{\ast}}\right)^{1/\beta_{\ast}^{\prime}}\geq y^{\prime}\right)=P_{\chi^{2}}\left(\frac{\left(\bar{F^{\prime}}_{t}^{h}\right)^{2\beta_{\ast}^{\prime}}}{\beta_{\ast}^{\prime 2}\rho_{\ast}^{2}\sigma_{t}^{2}hI_{t}^{h}};\frac{1}{\beta_{\ast}^{\prime}},\frac{\left(y^{\prime}\right)^{2\beta_{\ast}^{\prime}}}{\beta_{\ast}^{\prime 2}\rho_{\ast}^{2}\sigma_{t}^{2}hI_{t}^{h}}\right),$
 and this implies that
 $\left(\frac{\beta_{\ast}^{\prime}}{\beta_{\ast}}F_{t+h}^{\beta_{\ast}}\right)^{1/\beta_{\ast}^{\prime}}\sim\mathcal{CEV}_{\beta^{\prime}}\left(\bar{F^{\prime}}_{t}^{h},\rho_{\ast}^{2}\sigma_{t}^{2}hI_{t}^{h}\right).$ (27)
-In other words, a power of $F_{t+h}$ instead of $F_{t+h}$ itself, follows a CEV distribution under Islah’s approximation. Therefore, $F_{t+h}$ can be sampled as
+In other words, instead of $F_{t+h}$ itself, a power function of $F_{t+h}$ follows a CEV distribution under Islah’s approximation. Therefore, $F_{t+h}$ can be sampled as
 $F_{t+h}\sim\left(\frac{\beta_{\ast}}{\beta_{\ast}^{\prime}}\right)^{1/\beta_{\ast}}Y^{\beta_{\ast}^{\prime}/\beta_{\ast}}\quad\text{where}\quad Y\sim\mathcal{CEV}_{\beta^{\prime}}\left(\bar{F^{\prime}}_{t}^{h},\rho_{\ast}^{2}\sigma_{t}^{2}hI_{t}^{h}\right).$